### v1

# Evaluating LLM-corrupted Crowdsourcing Data Without Ground Truth

Abstract
The recent success of generative AI highlights the crucial role of high-quality human feedback in building trustworthy AI systems. However, the increasing use of large language models (LLMs) by crowdsourcing workers poses a significant challenge: datasets intended to reflect human input may be compromised by LLM-generated responses. Existing LLM detection approaches often rely on high-dimension training data such as text, making them unsuitable for annotation tasks like multiple-choice labeling. In this work, we investigate the potential of peer prediction — a mechanism that evaluates the information within workers’ responses without using ground truth — to mitigate LLM-assisted cheating in crowdsourcing with a focus on annotation tasks. Our approach quantifies the correlations between worker answers while conditioning on (a subset of) LLM-generated labels available to the requester. Building on prior research, we propose a training-free scoring mechanism with theoretical guarantees under a crowdsourcing model that accounts for LLM collusion. We establish conditions under which our method is effective and empirically demonstrate its robustness in detecting low-effort cheating on real-world crowdsourcing datasets.
1 Introduction
High-quality human feedback plays a crucial role in shaping modern AI systems, such as ChatGPT, to be trustworthy, safe, and aligned with human preference (Ouyang et al., 2022). This feedback is usually collected through crowdsourcing, where a requester (principal) posts a set of tasks and recruits workers (agents) — who may be trained experts or crowdsourcing workers — to complete them. For example, researchers widely use platforms like Amazon Mechanical Turk and Prolific for tasks such as data labeling and survey administration.
However, the integrity of this data collection process is increasingly threatened by the very AI models it seeks to improve. Evidence suggests that many agents are using large language models (LLMs) to complete tasks such as abstract summarization (Veselovsky et al., 2023) and peer review (Liang et al., 2024), leading to datasets that no longer reflect genuine human input. While LLM-assisted annotations or responses may sometimes be more accurate or coherent than human’s (Gilardi et al., 2023; Törnberg, 2023; Kaikaus et al., 2023), their widespread use clashes with the goal of gathering diverse and authentic human opinions. This issue is particularly concerning in applications that depend on human subjectivity or expertise, such as preference labeling and toxicity detection, where human judgment remains the gold standard.
How to detect potential LLM cheaters and evaluate the quality of a human-contributed dataset? Existing black-box methods for detecting LLM-generated content primarily focus on distinguishing human-written text from AI-generated text by analyzing statistical differences in writing style (Gehrmann et al., 2019; Mao et al., 2024; Koike et al., 2024; Hu et al., 2023). While sometimes effective in open-ended text generation tasks, these training-based approaches are not well-suited for annotation tasks, where the discrete responses lack the rich textual features needed for classification.
In this paper, we develop a theoretically robust scoring metric to evaluate the information in human contributions that is not captured by LLM responses. Our method builds on peer prediction, a scoring mechanism designed to incentivize truthful reporting in the absence of ground truth (Miller et al., 2005; Shnayder et al., 2016). Traditional peer prediction methods evaluate a worker based on the correlation between her responses and those of her peers. Under the standard model, every worker independently observes only high-effort signals — denoted as $X_{i}$ for worker $i$ — after working on a set of i.i.d. questions. Peer prediction mechanisms are shown to be information monotone, meaning that workers maximize their expected score by truthfully reporting their high-effort signals.
However, these methods are ineffective when workers can coordinate on cheap signals, denoted as $Z_{i}$ for worker $i$, that are highly correlated and require minimal effort to produce. In our settings, LLM responses serve as such cheap signals. To address this limitation, we extend peer prediction by measuring correlations between workers’ responses while conditioning on a cheap signal $Z$ that is observable to the requester. For example, $Z_{i}$ and $Z$ could represent responses to the questions independently generated by the same LLM or different LLMs. Intuitively, by conditioning on $Z$, our method scores the information within workers’ responses beyond what can be obtained from LLMs alone, ensuring that human effort is properly rewarded while LLM-assisted cheating is penalized.
Theoretically, we identify the conditions under which our mechanism maintains information monotonicity under two regimes: (1) for any form of cheating strategy and (2) for a subset of practical cheating strategies, which we refer to as naive strategies. Under a naive strategy, rather than reporting honestly on all assigned tasks, the agent sometimes reports noise (random reporting) or copies the LLM output, thereby mixing high-effort signals with uninformative signals. To achieve information monotonicity in the general case, we require that for any two workers $i$ and $j$, $Z_{i}$ (the cheap signal of $i$) is approximately uncorrelated with both $X_{j}$ (the high-effort signal of $j$) and $Z_{j}$ (the cheap signal of $j$), once we condition on $Z$ (see Assumptions 4.1 and 4.2). For naive strategies, a weaker condition suffices: conditioning on $Z$, the correlation between high-effort signals $(X_{i},X_{j})$ is stronger than the correlation between either $(Z_{i},X_{j})$ or $(Z_{i},Z_{j})$ on average (see 4.5).
Empirically, we test our assumptions using two subjective labeling datasets and five types of commonly used LLMs. Our findings show that human responses consistently exhibit stronger correlations than LLM-generated responses when samples of $Z$ and $Z_{i}$ originate from the same model, supporting the validity of 4.5. However, the correlations between independent samples of LLMs responses are usually non-zero, suggesting that Assumptions 4.1 and 4.2 are too strong to hold in practice. This implies that it is feasible to use peer prediction to detect cheating with naive strategies, but detecting arbitrary cheating — where agents choose to cheat even after seeing the high-effort signal — may require substantial ground truth verifications. Finally, under scenarios where 4.5 holds, we simulate cheating agents and evaluate the effectiveness of our method in identifying them. Echoing our theoretical insights, we show that our approach has the most robust performance on mixed crowds compared with all the baselines.
2 Related Work
Our study is relevant to the usage of LLMs on crowdsourcing platforms. On one hand, if used properly, LLMs can assist better human-AI collaborations or even replace human labor on particular crowdsourcing tasks. For example, Cegin et al. (2023) find that ChatGPT can, to some extent, replace crowdsourcing on paraphrase generation for intent classification tasks. Kaikaus et al. (2023) focus on emotion, sentiment, and cognitive dissonance annotations, where they found that GPT labels are more consistent and reliable than human annotators. Ashktorab et al. (2021) demonstrate that “batch labeling”, a framework of using AI to assist human labeling by allowing a single labeling action to apply to multiple records, can increase the overall labeling accuracy and increase the speed. On the other hand, however, many studies have warned of the dangers of improperly using LLMs in various crowdsourcing tasks. del Rio-Chanona et al. (2023) show that over recent years, LLMs have decreased the activities on Stack Overflow, the leading online Q&A platform for computer programming, by 16%. Regarding the quality of LLM-generated content, evidences exist showing that LLMs tend to have higher bias than human annotators on qualitative analysis (Ashwin et al., 2023), and they tend to exhibit a high degree of unreliability in questions answered in the context of undergraduate computer science. However, LLMs evolve rapidly — a considerable amount of recent work has focused on getting LLMs to emulate human expertise via extensive crowdsourced data (Bach et al., 2022; Zhang et al., 2025; Chiang and Lee, 2023; Köpf et al., 2023). We anticipate that the gap between LLMs and human performance will continue to shrink in the near future.
What’s more, the usage of LLMs in crowdsourcing has uncontrollably increased. In a study, it is reported that as high as 40% of MTurk workers used LLMs to cheat in text summarization tasks (Veselovsky et al., 2023). Furthermore, using a pre-trained LLM, Liang et al. (2024) reported that the use of LLMs in paper writing has significantly increased in recent years — the largest increase in CS papers was as high as 17%. Even worse, an empirical study suggests that humans tend to over-reliant on AI suggestions, even when they are incorrect (Campero et al., 2022). Taking into account both the opportunities and challenges, our work aims to address LLM cheating and foster a healthy and reliable crowdsourcing environment.
Our work builds on advancements in information elicitation, particularly peer prediction. First introduced by Miller et al. (2005), peer prediction mechanisms aim to incentivize truthful reporting without requiring ground truth verification. The core idea is to reward agents based on how well their responses predict those of others. The original mechanism relied on the principal’s knowledge of the correlation between agents’ signals (the information they observe). Later, Dasgupta and Ghosh (2013) extended this idea, removing the need for such knowledge by assuming agents answer multiple i.i.d. binary questions. Subsequent works have expanded this framework to non-binary (but finite) signals (Shnayder et al., 2016; Kong and Schoenebeck, 2019b), infinite signals (Schoenebeck and Yu, 2020), truthful guarantees with fewer questions (Kong, 2020b), heterogeneous agents (Agarwal et al., 2017), adversarial agents (Schoenebeck et al., 2021), and broader cheating strategies (Zhang and Schoenebeck, 2023b).
As a generalization of the correlated agreement (CA) mechanism (Shnayder et al., 2016), our method can score the quality of human responses conditioned on an external information source. We chose the CA mechanism for its simplicity and robust empirical performance, as alternative mechanisms often have complex designs or exhibit high output variance (Burrell and Schoenebeck, 2021; Zhang and Schoenebeck, 2023a; Xu et al., 2024). Our approach is also inspired by Kong and Schoenebeck (2018), who proposed using conditioned mutual information to elicit hierarchical information. Unlike their framework, which assumes an information hierarchy where more effort yields strictly more informative signals, our method does not require hierarchical structures or sequential elicitation. Instead, we assume the principal has access to samples of low-effort information and focus on empirically verifying the conditions under which our method works (or not). Lastly, we propose a heuristic generalization of our method to handle textual and other forms of complex data (Appendix C). This approach is related to the discussion of applying peer prediction to elicit textual information using LLMs (Lu et al., 2024; Wu and Hartline, 2024).
The studies on LLM-content detection are also related to this work. Common approaches for LLM content detection include watermarking (Gu et al., 2022; Kirchenbauer et al., 2023; Lee et al., 2023; Zhao et al., 2023; Yang et al., 2023a), which embeds identifiable markers into data, models, or generated text; statistical detection, which leverage differences in output by analyzing logits statistics (Su et al., 2023; Vasilatos et al., 2023) or vocabulary statistics (Gehrmann et al., 2019; Mao et al., 2024; Yang et al., 2023b); and deep learning-based methods, which train models to distinguish AI-generated content from human-written text (Koike et al., 2024; Hu et al., 2023; Shah et al., 2023). For a more comprehensive review, we refer to the survey paper (Wu et al., 2024). However, these methods typically rely on abundant textual data for training and detection. In contrast, our approach is training-free and addresses cheating in scenarios where the critical distinction between human-written and AI-generated content lies in the information level rather than the vocabulary level. For instance, our method is designed for tasks like answering multiple-choice questions, where no direct textual response is available.
3 Model
Consider a truth-discovery setting where there are $m$ i.i.d. questions. The principal distributes these questions to $n$ agents such that each agent answers multiple questions and each question is answered by more than one agent. We first present the information structure when two agents, denoted as $i$ and $j$, answer an arbitrary question. Because questions are i.i.d., the same information structure will be applied to any question.
Figure 1: The causal relationship among key variables.
Signals  Suppose the underlying ground truth of the question is $Y\in\Sigma$. For example, if the answer to the question is either “yes” or “no”, then $|\Sigma|=2$. Suppose agent $i$ and $j$ both answer this question. After careful thinking, they each observe a signal $X_{i}$ and $X_{j}$ respectively. For simplicity, we assume the signal space is the same as the ground truth space, i.e. $X_{i},X_{j}\in\Sigma$. We can interpret $X_{i}$ and $X_{j}$ as the high-effort signals, which are agents’ best guess of $Y$ after exerting a substantial amount of effort.
In addition to their high-effort signals, agents can obtain cheap signals $Z_{i},Z_{j}\in\Sigma$ that may or may not depend on $Y$. The cheap signals are usually strongly correlated, sometimes even more correlated than agents’ high-effort signals. For example, these can be the responses of two (potentially different) LLMs.
We present the causal relationship of these variables in Fig. 1. Note that $X_{i}$ and $Z_{i}$ are independent conditioned on $Y$, meaning that the high-effort signal is only correlated with the cheap signal via the ground truth. However, $Z_{i}$ and $Z_{j}$ may have additional correlations conditioned on $Y$. This may happen when agents cheat with LLMs trained on similar data.
Strategies  The principal’s goal is to collect high-effort signals from agents. This means that we consider scenarios where human expertise remains the gold standard, or AI underperforms human agents. Such situations arise when the notion of ground truth is inherently subjective, shaped by human preferences — such as when labelers are asked to rank LLM responses according to personal preference, a common practice in reinforcement learning with human feedback.
However, agents can strategically report based on their high-effort signal $X_{i}$ and the cheap signal $Z_{i}$. Mathematically, the report $\hat{X}_{i}$ is a (random) function of $X_{i}$ and $Z_{i}$, denoted as $\hat{X}_{i}=\theta_{i}(X_{i},Z_{i})$ where $\theta_{i}\in\Theta$. Let $\tau$ be the truth-telling strategy, i.e. $\tau(X_{i},Z_{i})=X_{i}$, and let $\mu_{i}$ denote an uninformative strategy such that $\mu_{i}(X_{i},Z_{i})$ is independent of $X_{i}$. The definitions for $\hat{X}_{j}$ and $\theta_{j}$ are analogous.
We highlight a special type of strategy, called the naive strategy, which is more natural and common on crowdsourcing platforms.
Definition 3.1.
A naive strategy $\nu\in\Theta_{\nu}$ is a convex combination between truth-telling and an uninformative strategy $\mu$, i.e.
$\nu(X_{i},Z_{i})=\begin{cases}X_{i}&\text{with probability }p_{i}<1,\\ \mu(X_{i},Z_{i})&\text{otherwise.}\end{cases}$
Examples of naive strategies include: exerting effort and reporting $X_{i}$ for some questions while randomly selecting answers for the rest; or fully relying on an LLM and always reporting $Z_{i}$. In contrast, non-naive strategies are typically complex and impractical, often resembling malicious behavior rather than mere low-effort cheating. For example, an agent may first exert effort to obtain $X_{i}$ for each question and then mix it with $Z_{i}$ by reporting “yes” if either $X_{i}$ or $Z_{i}$ is “yes”, and “no” otherwise.
Principal Information  The principal can observe the reports $\hat{X}_{i}$ and $\hat{X}_{j}$, but not the underlying signals $X_{i},X_{j}$ and $Z_{i},Z_{j}$. Furthermore, the principal cannot directly evaluate human responses using the ground truth $Y$. Without knowing $Y$, the principal cannot distinguish the real information $X_{i}$ and the cheap signal $Z_{i}$ without further assumptions. Our method assumes that the principal can observe a signal $Z$ that is (strongly) correlated with the agents’ cheap signals. In the context of LLM cheating, for example, the principal can generate answers to a fraction of the questions using a popular LLM.
The effectiveness of our method necessarily depends on the quality of $Z$ — how well the principal’s information can block the correlation between agents’ cheap signals. We will introduce our method and the assumptions in Section 4 and empirically verify the assumptions in Section 5.3.
3.1 Objectives and Problem Statement
In practice, the principal may collect responses from multiple agents, resulting in an $n\times m$ response matrix $\hat{X}$ where each entry $\hat{X}_{i,q}\in\Sigma$ if agent $i$ answers question $q$ and $\hat{X}_{i,q}=\phi$ otherwise. Given $\hat{X}$ and the principal’s samples of the cheap signal $Z$, a scoring mechanism designs a score for each agent. Our goal is to design a scoring mechanism that satisfies information monotonicity. Let $S_{i}(\hat{X}_{i},\hat{X}_{-i}\mid Z)$ denote the expected score of agent $i$, where $\hat{X}_{-i}$ is the reports of all agents but $i$.
Definition 3.2.
A mechanism is information monotone for naive strategies if, for any agent $i$, $S_{i}(X_{i},X_{-i}\mid Z)>S_{i}(\hat{X}_{i},\hat{X}_{-i}\mid Z)$ for any naive strategy $\nu_{i}\in\Theta_{\nu}$ with $\hat{X}_{i}=\nu_{i}(X_{i},Z_{i})$.
A mechanism is $\epsilon$-information monotone if it is information monotone for naive strategies and, additionally, $S_{i}(X_{i},X_{-i}\mid Z)\geq S_{i}(\hat{X}_{i},\hat{X}_{-i}\mid Z)-\epsilon$ holds for any strategy $\theta_{i}\in\Theta$ where $\hat{X}_{i}=\theta_{i}(X_{i},Z_{i})$.
At the individual level, information monotone scores can be used to identify unreliable labelers or LLM cheaters. At the dataset level, the score distribution across agents can help interpret the quality of the crowdsourced dataset.
We note that the same concept is termed informed truthfulness or informed strategy-proofness in the peer prediction literature, where the objective is to incentivize agents to report truthful information (Shnayder et al., 2016).
4 Our Method
Classic peer prediction methods often struggle with cheap signals that are highly correlated with one another but only weakly correlated with the ground truth. In this section, we generalize a prior work — the correlated agreement (CA) mechanism (introduced in Appendix A) — and propose a mechanism that achieves information monotonicity when the principal can obtain some noisy samples of the cheap signal. In the context of LLM cheating, these samples can be LLM-generated labels for the crowdsourced questions. . We name our generalization as the conditioned CA mechanism, which scores an agent based on the correlation between her responses and her peers’ conditioned on the principal’s samples of $Z$. Our method has two main steps.
Learn the scoring function. A scoring function determines whether a pair of signals “agree” or not, which is learned based on data. We first estimate the empirical conditioned joint distribution $\tilde{P}(\hat{X}_{i},\hat{X}_{j}\mid Z)$, computed using the frequency of observing a pair of signals on the same question conditioned on $Z$.***We can use all agents’ reports to estimate the same distribution because we assume agents are homogeneous and questions are i.i.d.. This assumption is primarily made for practical concerns, as there is usually a shortage of data to estimate the distribution accurately for each pair of agents. Next, we compute the delta tensor
$\tilde{\Delta}_{h,l,k}=\tilde{P}\left(\hat{X}_{i}=h,\hat{X}_{j}=l\mid Z=k\right)-\tilde{P}\left(\hat{X}_{i}=h\mid Z=k\right)\tilde{P}\left(\hat{X}_{j}=l\mid Z=k\right),$
which is the difference between the joint distribution and the product of marginal distributions of the reports, conditioned on $Z$. This gives us the scoring function $T_{\tilde{\Delta}}=\text{Sign}(\tilde{\Delta})$, i.e. $T_{\tilde{\Delta}}(h,l,k)=1$ if $\tilde{\Delta}_{h,l,k}>0$ and 00 otherwise. This step becomes unnecessary if the true scoring function $T_{\Delta}=\text{Sign}(\Delta)$ is already known, where $\Delta$ is defined analogously to $\tilde{\Delta}$, but based on the true signals $X_{i}$ and $X_{j}$. For example, for binary questions with positively correlated signals, $\Delta_{\cdot,\cdot,k}$ is an identity matrix for any $k\in\Sigma$.
Compute the bonus/penalty score. We then repeatedly sample a bonus question $q$ and a penalty question $q^{\prime}$ while ensuring $Z_{q}=Z_{q^{\prime}}=k$. For each pair of samples, the score for agent $i$ is:
$T_{\tilde{\Delta}}(\hat{X}_{i,q},\hat{X}_{j,q},k)-T_{\tilde{\Delta}}(\hat{X}_{i,q},\hat{X}_{j,q^{\prime}},k),$
where $j$ is a randomly selected agent who answers question $q^{\prime}$. The final score for agent $i$ is obtained by averaging over all sampled pairs of $q$ and $q^{\prime}$. Algorithm 1 presents the procedures of our method.
Input: The crowdsourced dataset $\hat{X}$, the samples of cheap signal $Z$, the (estimated) scoring function $T_{\tilde{\Delta}}$
Estimate the empirical marginal distribution of $Z$, i.e. $\tilde{P}(Z=\sigma)$ for $\sigma\in\Sigma$.
for agent $i\in[n]$ do
       Let $M_{i}$ be the set of questions answered by agent $i$ Initialize $s_{i}=0$;
       for $k\in\Sigma$ do
             Initialize $x=0$;
             for question $q\in M_{i}\cap M_{Z}(k)$ do
                   Randomly select a peer $j$ who answers $q$;
                   $x\mathrel{+}=T_{\tilde{\Delta}}(\hat{X}_{i,q},\hat{X}_{j,q},k)$;
                   Randomly select a question $q^{\prime}\in M_{j}\cap M_{Z}(k)$;
                   $x\mathrel{-}=T_{\tilde{\Delta}}(\hat{X}_{i,q},\hat{X}_{j,q^{\prime}},k)$
            $s_{i}\mathrel{+}=x\cdot\tilde{P}(Z=k)/|M_{i}\cap M_{Z}(k)|$
return $s$
Algorithm 1 The conditioned CA mechanism
Intuitively, the conditioned CA mechanism encourages agents to agree on the same question and penalizes agents when they agree on two distinct questions, given that the principal’s sampled signal is identical on both questions. If agents rely purely on the cheap signal or respond independently of the true signal, the bonus score and the penalty score cancel out in expectation. This results in a score that is close to zero.
4.1 Theoretical Insights
We demonstrate the effectiveness of our method and clarify the conditions under which the mechanism performs reliably. Before we start, we first present an intuitive way to interpret the score computed by our mechanism. In a prior work, Kong and Schoenebeck (2019a) show that the score of the CA mechanism can be interpreted as a type of mutual information between agents’ reports called the total variation distance (TVD) mutual information. Inspired by this prior work, we show that when agents report the high-effort signal and the underlying scoring function $T_{\Delta}$ is applied, the expected score computed by the conditioned CA mechanism is essentially the TVD conditioned mutual information:
$I_{\text{TVD}}(X_{i};X_{j}\mid Z)=\sum_{z\in\Sigma}\Pr(Z=z)\sum_{\sigma,\sigma^{\prime}\in\Sigma}\left|\Pr(X_{i}=\sigma,X_{j}=\sigma^{\prime}\mid z)-\Pr(X_{i}=\sigma\mid z)\Pr(X_{j}=\sigma^{\prime}\mid z)\right|.$
4.1.1 Conditions for $\epsilon$-Information Monotonicity
We first present the following pair of assumptions which require that the principal’s information $Z$ can approximately block any correlation between an agent’s cheap signal and the ground truth (4.1) and the correlation between agents’ cheap signals (4.2).
Assumption 4.1.
We assume $Z_{i}$ and $Y$ are approximately independent conditioned on $Z$, i.e. $|\Pr(Z_{i},Y\mid Z)-\Pr(Z_{i}\mid Z)\Pr(Y\mid Z)|\leq\epsilon_{i}$ for any $i\in[n]$.
Assumption 4.2.
We assume $Z_{i}$ and $Z_{j}$ are approximately independent conditioned on $Y$ and $Z$, i.e. $|\Pr(Z_{i},Z_{j}\mid Y,Z)-\Pr(Z_{i}\mid Y,Z)\Pr(Z_{j}\mid Y,Z)|\leq\epsilon$ for any $i,j\in[n]$.
We first show that these two assumptions are sufficient to guarantee (approximate) information monotonicity.
Theorem 4.3.
If $T_{\Delta}$ is known and $I_{\text{TVD}}(X_{i};X_{j}\mid Z)>\hat{\epsilon}\coloneqq\left((\epsilon+\epsilon_{i}+\epsilon_{j})|\Sigma|^{2}+(\epsilon_{i}+\epsilon_{j})|\Sigma|^{3}\right)$, then under 4.1 and 4.2, the conditioned CA mechanism is $\hat{\epsilon}$-information monotone.
We defer the proof to Section B.1. When agents report the high-effort signals truthfully, the expected score is the TVD mutual information between $X_{i}$ and $X_{j}$. Theorem 4.3 suggests that if $Z$ approximately blocks the correlation between $Z_{i}$ with $X_{j}$ and $Z_{j}$, then any cheating strategy can only decrease the TVD mutual information between the reports $\hat{X}_{i}$ and $\hat{X}_{j}$ up to an error. In the context of LLM cheating, suppose the principal suspects that some agents are using LLMs to generate their labels. Our mechanism suggests that the principal can generate labels using a (potentially different) LLM and compute the CA score conditioned on these generated labels. According to Theorem 4.3, if the principal’s generated labels satisfy two conditions — 1) they block correlations among the labels generated by potential cheating agents (4.2), and 2) they block correlations between those labels and the ground truth (4.1) — then the conditioned CA mechanism’s score is information monotone.
We further show the necessity of the assumptions for information monotonicity. We construct an example to show that if $Z_{i}$ and $Z_{j}$ have some correlation that is not captured by $Z$, then there exists a signal structure such that the score computed by the conditioned CA mechanism is not maximized by reporting $X_{i}$ and $X_{j}$.
Proposition 4.4.
Suppose $T_{\Delta}$ is known. If there exist $(z_{i},z_{j},y,z)$ such that $|\Pr(Z_{i}=z_{i},Z_{j}=z_{j}\mid Y=y,Z=z)-\Pr(Z_{i}=z_{i}\mid Y=y,Z=z)\Pr(Z_{j}=z_{j}\mid Y=y,Z=z)|=d$ for some $0<d\leq\frac{1}{4}$, then there exists a joint distribution $\Pr(X_{i},X_{j},Z_{i},Z_{j},Y,Z)$ that satisfies the causal relationship described in Fig. 1 and a strategy profile $(\theta_{i},\theta_{j})\neq(\tau,\tau)$ such that $S_{i}(\hat{X}_{i},\hat{X}_{j})\geq S_{i}(X_{i},X_{j})+8d^{2}$ where $\hat{X}_{i}=\theta_{i}(X_{i},Z_{i})$ and $\hat{X}_{j}=\theta_{j}(X_{j},Z_{j})$.
Before we provide the full proof (deferred to Section B.2), we briefly introduce the high-level idea. Consider a signal structure with a binary signal space $\Sigma=\{0,1\}$. Suppose agent $i$’s high-effort signal $X_{i}$ is perfectly accurate at predicting $X_{j}$ when $X_{i}=0$ but is noisy when $X_{i}=1$, i.e. $\Pr(X_{j}=0|X_{i}=0,Z)=1$ and $\Pr(X_{j}=1|X_{i}=1,Z)<1$. On the other hand, the cheap signal $Z_{i}$ is perfectly accurate at predicting $Z_{j}$ when $Z_{i}=1$ but is noisy when $Z_{i}=0$, i.e. $\Pr(Z_{j}=1|Z_{i}=1,Z)=1$ and $\Pr(Z_{j}=0|Z_{i}=0,Z)<1$. In this case, it is intuitive that agent $i$ can be better off if both agents combine their signals — reporting 0 if and only if $X_{i}=Z_{i}=0$ and reporting $1$ otherwise. Note that although this example only indicates that 4.2 is necessary, the same recipe can be straightforwardly applied to show the necessity of 4.1.
4.1.2 Conditions for Information Monotonicity Under Naive Strategies
We know 4.1 and 4.2 are necessary for the conditioned CA mechanism to be $\epsilon$-information monotone. However, these assumptions are designed to guard against complex, unrealistic strategies, such as those where cheaters first obtain the high-effort signal and then blend it with the cheap signal. If we instead restrict our focus to naive strategies, these assumptions on $Z$ can be significantly relaxed.
Assumption 4.5.
For any pair of agents $i$ and $j$, we assume the TVD conditioned mutual information between $X_{i}$, $X_{j}$, $Z_{i}$, $Z_{j}$ satisfies that
$I_{\text{TVD}}(X_{i};X_{j}\mid Z)>\max\left(I_{\text{TVD}}(Z_{i};Z_{j}\mid Z),I_{\text{TVD}}(Z_{i};X_{j}\mid Z)\right).$
Proposition 4.6.
If $T_{\Delta}$ is known and 4.5 holds, the conditioned CA mechanism is information monotone for naive strategies.
We defer the proof to Section B.3. Intuitively, a naive strategy is essentially a mixture of truth-telling and uninformative strategies. We show that when both agents follow naive strategies, the score-maximizing equilibrium must fall into one of the following three cases: 1) both agents exert full effort, achieving an expected score of $I_{\text{TVD}}(X_{i};X_{j}\mid Z)$; 2) both agents report cheap signals, with expected score $I_{\text{TVD}}(Z_{i};Z_{j}\mid Z)$; 3) one agent exerts full effort while the other reports a cheap signal, yielding expected score $I_{\text{TVD}}(Z_{i};X_{j}\mid Z)$. Therefore, the inequality in 4.5 ensures that full effort strictly dominates naive strategies, establishing the proposition.
Note that 4.5 is significantly weaker than 4.1 and 4.2, which require both $I_{\text{TVD}}(Z_{i};Z_{j}\mid Z)$ and $I_{\text{TVD}}(Z_{i};X_{j}\mid Z)$ to be close to zero.
4.1.3 Learning the scoring function $T_{\Delta}$.
Thus far, we have assumed that the true scoring function $T_{\Delta}$ is known. For complex signal structures, especially when the signal space is large, we have to learn the scoring function from the data. This introduces a concern: if a significant fraction of agents are cheating, the learned scoring function $T_{\tilde{\Delta}}$ might be manipulated in a way that benefits the cheating reports. We mitigate this concern using the following proposition. Let $S_{i}^{T}(\theta_{i},\theta_{j})$ be the expected score computed under the scoring function $T$ when agents’ strategies are $\theta_{i}$ and $\theta_{j}$ respectively.
Proposition 4.7.
Suppose 4.1 and 4.2 hold with no error such that $\epsilon=\epsilon_{i}=0$ for any $i\in[n]$. Then, for any scoring function $T\in\{0,1\}^{|\Sigma|\times|\Sigma|}$ and any strategy profile $(\theta_{i},\theta_{j})$, $S_{i}^{T}(\theta_{i},\theta_{j})\leq S_{i}^{T_{\Delta}}(\tau,\tau)$, where $\tau$ denote the truth-telling strategy.
We defer the proof to Section B.4. Intuitively, the proposition holds because when $T_{\Delta}$ is known and agents are truth-telling, the expected score is $I_{\text{TVD}}(X_{i};X_{j}\mid Z)$. However, if $T_{\Delta}$ is mislearned, the expected score is upper bounded by $I_{\text{TVD}}(\hat{X}_{i};\hat{X}_{j}\mid Z)$. Since our assumptions ensure that $Z_{i}$ and $Z_{j}$ are independent given $Z$, the data processing inequality guarantees that $I_{\text{TVD}}(\hat{X}_{i};\hat{X}_{j}\mid Z)<I_{\text{TVD}}(X_{i};X_{j}\mid Z)$.
4.7 implies that the expected score of the conditioned CA mechanism is maximized if every agent exerts full effort and reports $X_{i}$, while any strategy that manipulates the true scoring function will only (weakly) decrease the expected score.***Here we assume that the principal can learn the correct $T_{\Delta}$ if everyone reports their high-effort signal.
5 Experiments
We evaluate the effectiveness of our proposed method using real-world crowdsourcing datasets. We focus on labeling tasks where human judgment remains as the gold standard, given that current LLMs still fall short of perfect performance. We first empirically examine the conditions under which the key assumptions underlying our method are likely to hold, as well as when they may fail. Next, we simulate a mixed, noisy crowd of labelers and assess whether our method can detect unreliable labelers better than several well-established baselines.
5.1 Datasets
Hatefulness/Offensiveness/Toxicity Labeling  We first consider a toxicity labeling dataset which includes $3459$ social media user comments posted in response to political news posts and videos on Twitter, YouTube, and Reddit in August 2021 (Schöpke-Gonzalez et al., 2025). MTurk workers are assigned these comments and are requested to annotate them as hateful, offensive, and/or toxic. The definitions of hatefulness, offensiveness, and toxicity are selected from highly cited previous literature. In this dataset, every question is answered by $5$ workers, and each worker answers $73$ questions on average.
Preference Alignment
We further consider an LLM alignment dataset where each question is a comparison between two answers to the same prompt generated by two different LLMs. Labelers rate their preference on a 0–4 scale (0: A clearly better, 4: B clearly better, 2: tie). The comparison is scored from several more detailed angles, such as the helpfulness and the truthfulness of the answer. Nonetheless, we only focus on the overall preference. In total, there are 10 4611046110\,46110 461 questions where each question is answered by two normal labelers and two experts. On average, each labeler answers around $180$ questions. All labelers are recruited from Prolific. For more details, refer to Miranda et al. (2024).
5.2 Large Language Model Usage
In this work, we select five well-known LLMs to generate labels or annotations for all used datasets, including GPT-3.5-turbo, GPT-4, Gemma-2-2b-it***https://huggingface.co/google/gemma-2-2b-it, Phi-3.5-mini-Instruct ***https://huggingface.co/microsoft/Phi-3.5-mini-instruct, and Mistral-7B-Instruct-v0.3***https://huggingface.co/mistralai/Mistral-7B-Instruct-v0.3. The corresponding prompt templates are provided in Appendix E. Note that we choose only five models because other commonly used LLMs, such as LLaMA and Gemini, refuse to answer toxicity-related questions. In Appendix C, we use GPT-4 to process each human review and LLM-generated review into a $19$-dimension vector to implement our method.
5.3 Assumption Verification
We empirically test the following hypotheses while varying the LLMs that generate $Z_{i}$, $Z_{j}$, and $Z$.
  1. 1.
    $Z_{i}$ and $X_{j}$ are approximately independent given $Z$, i.e. $I_{\text{TVD}}(Z_{i};X_{j}\mid Z)\approx 0$.
  2. 2.
    $Z_{i}$ and $Z_{j}$ are approximately independent given $Z$, i.e. $I_{\text{TVD}}(Z_{i};Z_{j}\mid Z)\approx 0$.
  3. 3.
    $X_{i}$ and $X_{j}$ are more correlated than $Z_{i}$ and $X_{j}$ given $Z$, i.e. $I_{\text{TVD}}(X_{i};X_{j}\mid Z)>I_{\text{TVD}}(Z_{i};X_{j}\mid Z)$.
  4. 4.
    $X_{i}$ and $X_{j}$ are more correlated than $Z_{i}$ and $Z_{j}$ given $Z$, i.e. $I_{\text{TVD}}(X_{i};X_{j}\mid Z)>I_{\text{TVD}}(Z_{i};Z_{j}\mid Z)$.
The first two hypotheses correspond to 4.1 and 4.2 respectively and the other two hypotheses correspond to 4.5. More broadly, our results in this section provide valuable insights into the correlations between different LLMs and their correlations compared with human labelers in the context of subjective labeling questions.
For each considered LLM, we prompt it to independently generate responses to each question in the dataset three times using the default temperature. These responses are treated as samples of the cheap signals $Z_{i}$, $Z_{j}$, and $Z$. Furthermore, we randomly select two human responses in the original dataset as samples of $X_{i}$ and $X_{j}$, respectively. Using these samples across all tasks, we estimate the conditioned joint distributions and then compute the conditioned mutual information. We report the results for $I_{\text{TVD}}(Z_{i};X_{j}\mid Z)$ on two datasets in Fig. 2, while results for other datasets and for $I_{\text{TVD}}(Z_{i};Z_{j}\mid Z)$ are deferred to Section D.1.
(a) Hatefulness
(b) Preference Alignment
Figure 2: The TVD mutual information between $Z_{i}$ and $X_{j}$ conditioned on $Z$, i.e. $I_{\text{TVD}}(Z_{i};X_{j}\mid Z)$. LLM names on the x-axis denote the models used to sample $Z$ while each bar in the same group represents the model used to sample $Z_{i}$. We use “Random Cheaters” as a reference, who randomly selects an answer to each question according to the prior.
Hypotheses 1 and 2 generally do not hold.  In Fig. 2, we observe that when $Z_{i}$ and $Z$ are sampled from different models, the mutual information between $Z_{i}$ and $X_{j}$ is often significantly higher than that expected from the random reporting (red bar). Even when $Z_{i}$ and $Z$ are independently sampled by the same model, conditioning on $Z$ reduces but does not eliminate this correlation. This indicates that independently generated responses from the same LLM can remain dependent after conditioning on another independent sample, contradicting Hypothesis 1. As presented in Section D.1, this pattern continues to hold for $I_{\text{TVD}}(Z_{i};Z_{j}\mid Z)$, rejecting hypothesis 2.
Hypothesis 3 and 4 generally hold when conditioning on the “right” model.  Encouragingly, the conditioned correlation between an LLM and a human labeler is consistently weaker than that between two human labelers when $Z_{i}$ and $Z$ are sampled from the same model. For example, in Fig. 2(b), the unconditioned mutual information between GPT-4-generated labels and human labels initially exceeds that of two expert labels. However, after conditioning on independent GPT-4 responses, $I_{\text{TVD}}(X_{i};X_{j}\mid Z)$ becomes significantly higher than $I_{\text{TVD}}(Z_{i};X_{j}\mid Z)$. This pattern is consistent across all tested datasets, providing strong support for hypothesis 3. Results for Hypothesis 4 are deferred to the appendix.
Our empirical results suggest that as of the time of this study and based on the subjective labeling tasks we tested, human labelers exhibit distinct correlation patterns that no LLM replicates.
5.4 Detecting Cheating Agents
In this subsection, we focus on identifying cheating agents with naive strategies and comparing various methods based on their area under the ROC curve (AUC).
5.4.1 Simulating Noisy Crowds
Real-world crowdsourcing datasets do not include labels identifying whether an agent is cheating or working diligently. To enable evaluation, we simulate cheating agents and treat the original dataset labels as representing “high-effort” responses. Specifically, we replace a randomly selected fraction of labelers in the original dataset with cheaters adopting naive strategies. To reflect the noisy nature of real-world crowdsourcing data, we simulate three types of cheaters, each representing a common naive strategy observed in practice (Yang et al., 2019; Becker et al., 2021):
  1. 1.
    An $\alpha_{llm}$ fraction of labelers are LLM cheaters, who report the LLM-generated labels on all assigned questions.
  2. 2.
    An $\alpha_{r}$ fraction of labelers are random cheaters, who generate labels by independently sampling from the marginal distribution over labels for each assigned question.
  3. 3.
    An $\alpha_{b}$ fraction of labelers are biased cheaters, who report the dataset’s majority label on 90% of questions and choose uniformly at random on the rest.
Figure 3: Distribution of the conditioned CA scores with $(\alpha_{llm},\alpha_{r},\alpha_{b})=(0.1,0.05,0.05)$, where GPT-4 generates samples for $Z_{i}$, $Z_{j}$, and $Z$ using the preference alignment dataset.
Figure 3 shows the distribution of conditioned CA scores, where both the LLM cheaters’ responses and the principal’s signals are independently sampled from separate GPT-4 outputs. As shown, human workers achieve higher scores than any type of cheater on average. However, some workers fall below the average score of cheaters. Several factors contribute to this issue. For example, a subset of workers may themselves be cheaters in the dataset, or they may have been assigned many tasks that are inherently difficult to classify. Consequently, no detection method can achieve perfect accuracy. Next, we will examine how well our method distinguishes workers from cheaters compared to other baselines.
5.4.2 Baseline Measurements
We introduce the key ideas of the following four baselines.
  1. 1.
    We consider the Output Agreement (OA) mechanism that directly counts the frequency of agreements between two agent’s responses. Specifically, while scoring labeler $i$, we iteratively pair her with a different labeler $j$ and let $M_{i,j}$ denote the set of questions answered by both labelers. Then, the score of labeler $i$ is the average of the empirical probability of agreement, i.e. $\frac{1}{n}\sum_{j\in[n]\backslash\{i\}}\sum_{q\in M_{i,j}}\frac{\mathbbm{1}\left[\hat{X}_{i,q}=\hat{X}_{j,q}\right]}{|M_{i,j}|}$, and let $\frac{0}{0}=0$.
  2. 2.
    We use the Correlated Agreement (CA) mechanism as a baseline, which is introduced in Appendix A. The expected CA score is the unconditioned mutual information $I_{\text{TVD}}(\hat{X}_{i};\hat{X}_{j})$.
  3. 3.
    Another common method to estimate the reliability of the noisy agents is the Expectation-Maximization (EM) algorithm under the DS model (Dawid and Skene, 1979). Under the DS model, each agent $i$ makes mistakes according to a confusion matrix $\Gamma^{i}$ where $\Gamma^{i}_{h,l}$ denotes the probability that labeler $i$ reports $h$ when the ground truth answer is $l$. The EM algorithm first initializes a prior of the ground truth for each question and a confusion matrix for each agent. Then, it iteratively goes through an E-step and an M-step. In the E-step, the algorithm computes the posterior of the ground truth of each question given the prior and the confusion matrix. In the M-step, the confusion matrix of each agent is updated using the posterior computed in the E-step. The above two steps are iterated until the likelihood of the data matrix $\hat{X}$ converges. Finally, we score the reliability of agent $i$ as $\sum_{h\in\Sigma}\Gamma^{i}_{h,h}\tilde{P}(h)$ where $\tilde{P}(h)$ is the empirical marginal distribution of observing $h$ in the dataset.
  4. 4.
    Lastly, we consider the Conditioned Agreement ($\text{OA}_{Z}$) which counts the frequency of agreements only when the agreement is different from the principal’s sample on that question. Similar to the agreement metric, the score of labeler $i$ is the average of the empirical probability of agreement conditioned on $Z$, i.e. $\frac{1}{n}\sum_{j\in[n]\backslash\{i\}}\sum_{q\in M_{i,j}}\frac{\mathbbm{1}\left[\hat{X}_{i,q}=\hat{X}_{j,q}\&\hat{X}_{i,q}\neq Z_{q}\right]}{|M_{i,j}|}$. Intuitively, if labeler $i$ uses an LLM whose generated labels are similar to $Z$, then $\hat{X}_{i,q}\neq Z_{q}$ will happen with a small probability, meaning that the conditioned agreement score will be small.
We note that there are many more heuristic methods to estimate the reliability of crowdsourcing agents (Yin et al., 2008; Embretson and Reise, 2013; Piech et al., 2013). Many of these methods are variants of the EM algorithm under a different context-dependent model. We nonetheless only use the EM algorithm with the DS model to represent this line of methods because they all ignore the principal’s information $Z$ and thus are unlikely to perform well with LLM-corrupted datasets.
5.4.3 Results
(a) AUC v.s. fraction of LLM cheaters.
(b) AUC v.s. sampled questions.
Figure 4: The area under the ROC curve (AUC) for various methods on the preference alignment dataset as we vary (a) the fraction of LLM cheaters and (b) the proportion of questions the principal can sample. In (a), GPT-4 generates samples for $Z_{i}$, $Z_{j}$, and $Z$; in (b), AUC curves for all five LLM-cheater types are shown (omitting error bars for clarity).
Given a dataset $\hat{X}$ and samples of the cheap signal $Z$, we use our method as well as the baselines to score each labeler. We use the AUC score to measure the ability of each discussed metric to distinguish the simulated cheaters from the original human labelers.
Figure 4(a) presents a comparison of different measures while varying the fraction of LLM cheaters in the crowds. For each value of $\alpha_{llm}$ ranging from 00 to $0.2$, we randomly sample $\alpha_{r}$ and $\alpha_{b}$ uniformly from $[0,0.2]$ and report the average and error bars over $50$ such trials. Our result suggests that classic methods, such as the EM algorithm and the CA mechanism, perform well when the number of LLM cheaters is low. However, their AUCs degrade significantly as the fraction of LLM cheaters increases. On the other hand, the performance of $\text{OA}_{Z}$, which scores agents based on the similarity between their reports and $Z$, tends to have large variance and performs poorly when there are few LLM cheaters. In contrast, our proposed conditioned CA mechanism has the most robust performance for the mixed crowds and has the lowest variance.
Table 1 summarizes the results from Figure 4(a), reporting the average AUC score and the bottom 10% quantile for each method on each dataset w.r.t. the randomness of cheater fractions. A high average AUC reflects strong overall performance across random cheater compositions, while a high 10% quantile indicates robustness to the worst-case scenarios. The results support our earlier claim that our method consistently demonstrates the highest robustness across all tested settings.
Table 1: Average AUC (and bottom 10% quantiles) across datasets and methods with GPT-4 cheaters.
Dataset OA CA OAZ EM CAZ (ours)
Hatefulness 0.66 (0.45) 0.84 (0.75) 0.75 (0.64) 0.81 (0.70) 0.85 (0.81)
Offensiveness 0.80 (0.68) 0.92 (0.88) 0.92 (0.85) 0.89 (0.82) 0.94 (0.91)
Toxicity 0.60 (0.40) 0.89 (0.83) 0.89 (0.82) 0.85 (0.79) 0.91 (0.87)
Preference Alignment 0.65 (0.40) 0.70 (0.43) 0.70 (0.45) 0.68 (0.41) 0.85 (0.77)
Our experiments do not claim that our method outperforms all baselines in every scenario. Traditional methods like EM often perform better when all agents are random or biased cheaters. However, our method demonstrates the most robust performance across all settings due to its theoretical guarantees. This is particularly useful as, in practice, the designer usually has little clue about the composition of the crowds. In contrast, the baselines exhibit cases where the AUC drops below 0.5, meaning the algorithm completely misclassifies good workers as cheating agents.
Detection Performance vs. Number of Samples
We further examine how many questions the principal needs to sample to effectively detect cheaters. In Fig. 4(b), the AUC of the conditioned CA mechanism is shown with the LLM-cheater fraction fixed at 0.15, while the fraction of random and biased cheaters is each independently sampled from $[0,0.2]$. Similar plots for other datasets are provided in Section D.2.2. As the sampling rate increases from 20% to 100%, the average AUC also increases. While the results vary across the datasets and the LLMs that cheaters use, we observe that, except for the hatefulness labeling dataset with GPT-3.5 cheaters, sampling 50% of the questions is sufficient to outperform the baseline CA mechanism. Furthermore, in cases where human and AI labels are well separated, sampling just 20% can suffice.
Additional results in the appendix.
First, Section D.2.1 presents additional figures and tables for other datasets and LLMs, further supporting the claims made in the main body. Furthermore, when the principal is uncertain about which LLM the cheaters are using, we propose a method that computes the CA score conditioned on each candidate LLM’s outputs and uses the minimum score across them as the final score (Section D.2.3). Finally, in Appendix C and Section D.3, we present a heuristic extension of our method to open-ended textual responses. This serves as a preliminary step toward generalizing our approach beyond labeling tasks, with further refinement and large-scale validation left for future work.
6 Limitations
We acknowledge several limitations of our method. First, our method is primarily designed for crowdsourcing datasets comprising subjective questions or tasks where human experts are considered the gold standard. If LLMs significantly surpass human performance, the correlations between human responses conditioned on $Z$ may no longer reflect the ground truth $Y$. In such cases, applying our method could mistakenly reward incorrect correlations.
Second, as discussed in our experiment section, our method does not outperform all baselines in every scenario. In particular, when LLM responses are diverse, e.g. every cheating agent applies a different model or there are only a few cheating agents, heuristic methods such as the EM algorithm work well. This is because diverse or sparse cheaters are less correlated with the majority of the crowd and can be easily flagged as outliers using the classic methods. Despite this, our method demonstrates the most robust performance across a wide range of settings. Future work could explore integrating the principal’s samples of potential cheating signals into heuristic approaches like EM to enhance their performance and optimize detection.
Third, our experiments evaluate only five language models, as some models, such as LLaMA and Gemini, refuse to respond to toxicity-related prompts. While this is sufficient to demonstrate the method’s effectiveness, additional evaluations across more diverse models and task types would provide a more comprehensive assessment.
7 Conclusion
This work addresses the challenge of LLM-assisted cheating in crowdsourcing by developing a theoretically grounded quality measure for annotation tasks. Leveraging an extension of peer prediction, our method evaluates human information beyond what LLMs can generate, ensuring reliable human contributions. We establish theoretical guarantees under different cheating strategies and validate our approach empirically on real-world datasets. Our results demonstrate that our method effectively detects low-effort cheating while maintaining robustness across various conditions. Future work can focus on refining our method to better differentiate between harmful cheating and productive human-AI collaboration. This includes optimizing scoring mechanisms to recognize when LLMs are used as assistive tools rather than substitutes for human effort and developing adaptive detection techniques that adjust based on task characteristics and worker behavior.
Acknowledgment
This work was partly carried out at the DIMACS Center at Rutgers University. Jinlong Pang and Yang Liu are partially supported by the National Science Foundation (NSF) under grant IIS-2143895.
Appendix A Preliminary
What do we know about the problem? Under the assumption that there are no cheap signals, the peer prediction literature suggests that estimating the mutual information between $X_{i}$ and $X_{j}$ is an information monotone metric [Kong and Schoenebeck, 2019a]. There are various ways to estimate the mutual information using agents’ responses to multiple questions [Kong, 2020a, Schoenebeck and Yu, 2020] and many information monotone metrics that can be explained using mutual information [Dasgupta and Ghosh, 2013, Shnayder et al., 2016]. In this section, we discuss a simple yet intuitive mechanism called the correlated agreement (CA) mechanism [Shnayder et al., 2016].
The CA mechanism rewards agents when they “agree” on the same question and penalizes them when they “agree” on distinct questions. Instead of directly checking agreement, the mechanism redefines agreement based on the correlation between signals. A pair of signals $\sigma,\sigma^{\prime}\in\Sigma$ is correlated agreed if $\Pr(X_{1}=\sigma,X_{2}=\sigma^{\prime})-\Pr(X_{1}=\sigma)\Pr(X_{2}=\sigma^{\prime})>0$, meaning that $\sigma$ and $\sigma^{\prime}$ are more likely to appear on the same question rather than on distinct questions. In particular, the mechanism utilizes the delta matrix $\Delta_{\sigma,\sigma^{\prime}}=\Pr(X_{1}=\sigma,X_{2}=\sigma^{\prime})-\Pr(X_{1}=\sigma)\Pr(X_{2}=\sigma^{\prime})$ to design the scoring function $T_{\Delta}=\text{Sign}(\Delta)$ where $\text{Sign}(x)=1$ if $x>0$ and 0 otherwise.
In practice, the mechanism can be implemented in two ways depending on the information available to the principal. First, if $T_{\Delta}$ is known, the mechanism can be directly implemented, as shown in Algorithm 2. This case happens if the information structure is simple, e.g. when questions have binary answers, $T_{\Delta}$ is likely to be a $2\times 2$ identical matrix. Second, if $T_{\Delta}$ is unknown, the mechanism can first estimate the joint distribution, denoted as $\widetilde{\Pr}(X_{1},X_{2})$ using the report matrix $\hat{X}$. Then, an estimate of $T_{\Delta}$ can be computed using the estimated Delta matrix $\tilde{\Delta}_{\sigma,\sigma^{\prime}}=\widetilde{\Pr}(X_{1}=\sigma,X_{2}=\sigma^{\prime})-\widetilde{\Pr}(X_{1}=\sigma)\Pr(X_{2}=\sigma^{\prime})$. Next, Algorithm 2 can be applied.
We emphasize that penalizing agreements on distinct questions is crucial for the CA mechanism to maintain information monotonicity. Without such penalization, over-reporting the majority signal could result in a high score. This form of manipulation is particularly problematic when the dataset is biased. For instance, in a task to label messages as toxic or not, if only 5% of the data points are toxic, consistently reporting “not toxic” could yield a higher score than truthfully reporting the actual signal.
Therefore, the score computed by the CA mechanism can be interpreted in the following way. Let the marginal distribution of agent $i$’s reports be $\Pr(\hat{X}_{i})$. Imagine an agent $k$ who reports randomly for each question according to $\Pr(\hat{X}_{i})$. Then, the CA score for agent $i$ can be understood as the average number of questions on which agent $i$ agrees with another randomly selected agent $j$ more than agent $k$ agrees with $j$.
Input: The crowdsourced dataset $\hat{X}$, the (estimated) scoring function $T_{\tilde{\Delta}}$
Output: A score vector $s$
for agent $i\in[n]$ do
       Let $M_{i}$ be the set of questions answered by agent $i$
       Initialize $x\leftarrow 0$
       for question $q\in M_{i}$ do
             Randomly select a peer $j$ who answered question $q$
             $x\leftarrow x+T_{\tilde{\Delta}}(\hat{X}_{i,q},\hat{X}_{j,q})$ Randomly select a question $q^{\prime}\neq q$ answered by agent $j$
             $x\leftarrow x-T_{\tilde{\Delta}}(\hat{X}_{i,q},\hat{X}_{j,q^{\prime}})$
      $s_{i}\leftarrow x/|M_{i}|$
return $s$
Algorithm 2 The Correlated Agreement Mechanism [Shnayder et al., 2016]
Appendix B Proofs
B.1 Sufficiency of the Assumptions
Proof of Theorem 4.3.
Intuitively, 4.1 and 4.2 imply that conditioned on $Z$, $Z_{i}$ is almost independent of $X_{i}$, $X_{j}$, and $Z_{j}$. Therefore, any manipulation that makes $\hat{X}_{i}$ depend on $Z_{i}$ will not increase the expected score. To prove this, we first write down the expected score when agent $i$ and $j$ play strategy $\theta_{i}$ and $\theta_{j}$ respectively.
$S_{i}(\theta_{i},\theta_{j})$
$=$ $\sum_{z\in\Sigma}\Pr(Z=z)\sum_{x_{i},x_{j},z_{i},z_{j}\in\Sigma}\underbrace{\Pr(X_{i}=x_{i},X_{j}=x_{j},Z_{i}=z_{i},Z_{j}=z_{j}\mid z)\;T_{\Delta}(\theta_{i}(x_{i},z_{i}),\theta_{j}(x_{j},z_{j}),z)}_{\text{the expected score of the bonus question}}$
$-\sum_{z\in\Sigma}\Pr(Z=z)\sum_{x_{i},x_{j},z_{i},z_{j}\in\Sigma}\underbrace{\Pr(X_{i}=x_{i},Z_{i}=z_{i}\mid z)\;\Pr(X_{j}=x_{j},Z_{j}=z_{j}\mid z)\;T_{\Delta}(\theta_{i}(x_{i},z_{i}),\theta_{j}(x_{j},z_{j}),z)}_{\text{the expected score of the penalty question}}.$
For comparison, we next write down the expected score when both agents report truthfully.
$S_{i}(\tau,\tau)=$ $\sum_{z\in\Sigma}\Pr(Z=z)\sum_{x_{i},x_{j}\in\Sigma}\underbrace{\Pr(X_{i}=x_{i},X_{j}=x_{j}\mid z)\;T_{\Delta}(x_{i},x_{j},z)}_{\text{the expected score of the bonus question}}$
$-\sum_{z\in\Sigma}\Pr(Z=z)\sum_{x_{i},x_{j}\in\Sigma}\underbrace{\Pr(X_{i}=x_{i}\mid z)\;\Pr(X_{j}=x_{j}\mid z)\;T_{\Delta}(x_{i},x_{j},z)}_{\text{the expected score of the penalty question}}$
$=$ $\sum_{z\in\Sigma}\Pr(Z=z)\sum_{x_{i},x_{j}\in\Sigma}(\Pr(x_{i},x_{j}\mid z)-\Pr(x_{i}\mid z)\;\Pr(x_{j}\mid z))\;T_{\Delta}(x_{i},x_{j},z)$
$=$ $\sum_{z\in\Sigma}\Pr(Z=z)\sum_{x_{i},x_{j}\in\Sigma}|\Pr(x_{i},x_{j}\mid z)-\Pr(x_{i}\mid z)|$
$=$ $I_{\text{TVD}}(X_{i};X_{j}\mid Z)$
We want to show that $S_{i}(\theta_{i},\theta_{j})\leq S_{i}(\tau,\tau)$ up to an error. As shown above, the expected score can be decomposed into the expected score of the bonus question and the expected score of the penalty question. Next, we provide an upper bound on the bonus score and a lower bound on the penalty score.
Lower bound the penalty score.  We first provide a lower bound for $\Pr(X_{i},Z_{i}\mid Z)$, which is summarized in the following lemma.
Lemma B.1.
Under 4.1, $\Pr(X_{i},Z_{i}\mid Z)\geq\Pr(Z_{i}\mid Z)\Pr(X_{i}\mid Z)-\epsilon_{i}\sum_{y\in\Sigma}\Pr(X_{i}\mid Y=y,Z)$.
Then, we provide a lower bound of the penalty score while fixing a $Z=z$.
$\sum_{x_{i},x_{j},z_{i},z_{j}\in\Sigma}\Pr(x_{i},z_{i}\mid z)\;\Pr(x_{j},z_{j}\mid z)\;T_{\Delta}(\theta_{i}(x_{i},z_{i}),\theta_{j}(x_{j},z_{j}),z)$
$\geq$ $\sum_{x_{i},x_{j},z_{i},z_{j}\in\Sigma}\left(\Pr(z_{i}\mid z)\Pr(x_{i}\mid z)-\epsilon_{i}\sum_{y\in\Sigma}\Pr(x_{i}\mid y,z)\right)\;\Pr(x_{j},z_{j}\mid z)\;T_{\Delta}(\theta_{i}(x_{i},z_{i}),\theta_{j}(x_{j},z_{j}),z).$ (By Lemma B.1)
By reordering the summations,
$=$ $\sum_{x_{i},x_{j},z_{i},z_{j}\in\Sigma}\Pr(z_{i}\mid z)\Pr(x_{i}\mid z)\;\Pr(x_{j},z_{j}\mid z)\;T_{\Delta}(\theta_{i}(x_{i},z_{i}),\theta_{j}(x_{j},z_{j}),z)$
$-\epsilon_{i}\sum_{y,z_{i}\in\Sigma}\sum_{x_{j},z_{j}\in\Sigma}\Pr(x_{j},z_{j}\mid z)\sum_{x_{i}\in\Sigma}\Pr(x_{i}\mid y,z)T_{\Delta}(\theta_{i}(x_{i},z_{i}),\theta_{j}(x_{j},z_{j}),z).$
We focus on the second term. First note that $T_{\Delta}\leq 1$. Therefore, the second term is lower bounded by $-\epsilon_{i}\sum_{y,z_{i}\in\Sigma}\sum_{x_{j},z_{j}\in\Sigma}\Pr(x_{j},z_{j}\mid z)\sum_{x_{i}\in\Sigma}\Pr(x_{i}\mid y,z)$. Marginalizing over $x_{j},z_{j}$ and $x_{i}$ and observing that $y,z_{i}\in\Sigma$, the quality is lower-bounded by $\epsilon_{i}|\Sigma|^{2}$. This means the original quantity is lower-bounded by
$\geq$ $\sum_{x_{i},x_{j},z_{i},z_{j}\in\Sigma}\Pr(z_{i}\mid z)\Pr(x_{i}\mid z)\;\Pr(x_{j},z_{j}\mid z)\;T_{\Delta}(\theta_{i}(x_{i},z_{i}),\theta_{j}(x_{j},z_{j}),z)-\epsilon_{i}|\Sigma|^{2}.$
Next, we apply the same derivation for $j$.
$=$ $\sum_{x_{i},x_{j},z_{i},z_{j}\in\Sigma}\Pr(z_{i}\mid z)\Pr(x_{i}\mid z)\;\left(\Pr(z_{j}\mid z)\Pr(x_{j}\mid z)-\epsilon_{j}\sum_{y\in\Sigma}\Pr(x_{j}\mid y,z)\right)$ (1)
$\qquad\qquad\cdot T_{\Delta}(\theta_{i}(x_{i},z_{i}),\theta_{j}(x_{j},z_{j}),z)-\epsilon_{i}|\Sigma|^{2}$ (By Lemma B.1)
$\geq$ $\sum_{x_{i},x_{j},z_{i},z_{j}\in\Sigma}\Pr(z_{i}\mid z)\Pr(x_{i}\mid z)\Pr(z_{j}\mid z)\Pr(x_{j}\mid z)\;T_{\Delta}(\theta_{i}(x_{i},z_{i}),\theta_{j}(x_{j},z_{j}),z)-(\epsilon_{i}+\epsilon_{j})|\Sigma|^{2}.$ (2)
Upper bound the bonus score. We first present a lemma that bounds the difference between $\Pr(Z_{i},Z_{j}\mid Y,Z)$ and $\Pr(Z_{i}\mid Z)\Pr(Z_{j}\mid Z)$.
Lemma B.2.
Under 4.1 and 4.2, $\left|\Pr(Z_{i},Z_{j}\mid Y,Z)-\Pr(Z_{i}\mid Z)\Pr(Z_{j}\mid Z)\right|\leq\epsilon+\frac{\epsilon_{i}+\epsilon_{j}}{\Pr(Y|Z)}$.
With this lemma, we further provide an upper bound of the expected score on the bonus question. Fixing $Z=z$,
$\sum_{x_{i},x_{j},z_{i},z_{j}\in\Sigma}\Pr(x_{i},x_{j},z_{i},z_{j}\mid z)\;T_{\Delta}(\theta_{i}(x_{i},z_{i}),\theta_{j}(x_{j},z_{j}),z)$
$=$ $\sum_{x_{i},x_{j},z_{i},z_{j}\in\Sigma}\sum_{y\in\Sigma}\Pr(y\mid z)\Pr(x_{i},x_{j},z_{i},z_{j}\mid y,z)\;T_{\Delta}(\theta_{i}(x_{i},z_{i}),\theta_{j}(x_{j},z_{j}),z)$ (By the law of total probability)
$=$ $\sum_{x_{i},x_{j},z_{i},z_{j}\in\Sigma}\sum_{y\in\Sigma}\Pr(y\mid z)\Pr(x_{i},x_{j}\mid y,z)\Pr(z_{i},z_{j}\mid y,z)\;T_{\Delta}(\theta_{i}(x_{i},z_{i}),\theta_{j}(x_{j},z_{j}),z)$ ($X$ and $Z$ are independent conditioned on $Y$.)
$\leq$ $\sum_{x_{i},x_{j},z_{i},z_{j}\in\Sigma}\sum_{y\in\Sigma}\Pr(y\mid z)\Pr(x_{i},x_{j}\mid y,z)\left(\Pr(z_{i}\mid z)\Pr(z_{j}\mid z)+\epsilon+\frac{\epsilon_{i}+\epsilon_{j}}{\Pr(y|z)}\right)\;T_{\Delta}(\theta_{i}(x_{i},z_{i}),\theta_{j}(x_{j},z_{j}),z).$ (By Lemma B.2)
We can further write this into three terms.
$=$ $\sum_{x_{i},x_{j},z_{i},z_{j}\in\Sigma}\sum_{y\in\Sigma}\Pr(y\mid z)\Pr(x_{i},x_{j}\mid y,z)\Pr(z_{i}\mid z)\Pr(z_{j}\mid z)\;T_{\Delta}(\theta_{i}(x_{i},z_{i}),\theta_{j}(x_{j},z_{j}),z)$ (3)
$+\epsilon\;\sum_{x_{i},x_{j},z_{i},z_{j}\in\Sigma}\sum_{y\in\Sigma}\Pr(y\mid z)\Pr(x_{i},x_{j}\mid y,z)\;T_{\Delta}(\theta_{i}(x_{i},z_{i}),\theta_{j}(x_{j},z_{j}),z)$ (4)
$+(\epsilon_{i}+\epsilon_{j})\sum_{x_{i},x_{j},z_{i},z_{j}\in\Sigma}\sum_{y\in\Sigma}\Pr(x_{i},x_{j}\mid y,z)\;T_{\Delta}(\theta_{i}(x_{i},z_{i}),\theta_{j}(x_{j},z_{j}),z).$ (5)
First note that
Eq. 3 $=\sum_{x_{i},x_{j},z_{i},z_{j}\in\Sigma}\Pr(z_{i}\mid z)\Pr(z_{j}\mid z)\;T_{\Delta}(\theta_{i}(x_{i},z_{i}),\theta_{j}(x_{j},z_{j}),z)\sum_{y\in\Sigma}\Pr(x_{i},x_{j},y\mid z)$
$=\sum_{x_{i},x_{j},z_{i},z_{j}\in\Sigma}\Pr(z_{i}\mid z)\Pr(z_{j}\mid z)\;T_{\Delta}(\theta_{i}(x_{i},z_{i}),\theta_{j}(x_{j},z_{j}),z)\Pr(x_{i},x_{j}\mid z)$
Next, we upper bound the second and the third term respectively. By reordering the summations,
Eq. 4 $=\epsilon\;\sum_{x_{i},x_{j},z_{i},z_{j}\in\Sigma}T_{\Delta}(\theta_{i}(x_{i},z_{i}),\theta_{j}(x_{j},z_{j}),z)\;\sum_{y\in\Sigma}\Pr(x_{i},x_{j},y\mid z)$
$=\epsilon\;\sum_{x_{i},x_{j},z_{i},z_{j}\in\Sigma}T_{\Delta}(\theta_{i}(x_{i},z_{i}),\theta_{j}(x_{j},z_{j}),z)\;\Pr(x_{i},x_{j}\mid z)$
$\leq\epsilon\;\sum_{x_{i},x_{j},z_{i},z_{j}\in\Sigma}\Pr(x_{i},x_{j}\mid z)$ ($T_{\Delta}\leq 1$)
$\leq\epsilon|\Sigma|^{2}.$ ($\sum_{x_{i},x_{j}\in\Sigma}\Pr(x_{i},x_{j}\mid z)=1$)
Eq. 5 $=(\epsilon_{i}+\epsilon_{j})\;\sum_{y\in\Sigma}\sum_{z_{i},z_{j}\in\Sigma}\sum_{x_{i},x_{j}\in\Sigma}\Pr(x_{i},x_{j}\mid y,z)T_{\Delta}(\theta_{i}(x_{i},z_{i}),\theta_{j}(x_{j},z_{j}),z)$
$\leq(\epsilon_{i}+\epsilon_{j})\;\sum_{y\in\Sigma}\sum_{z_{i},z_{j}\in\Sigma}\sum_{x_{i},x_{j}\in\Sigma}\Pr(x_{i},x_{j}\mid y,z)$ ($T_{\Delta}\leq 1$)
$\leq(\epsilon_{i}+\epsilon_{j})|\Sigma|^{3}.$ ($\sum_{x_{i},x_{j}\in\Sigma}\Pr(x_{i},x_{j}\mid y,z)=1$)
Combining everything,
$\sum_{x_{i},x_{j},z_{i},z_{j}\in\Sigma}\Pr(x_{i},x_{j},z_{i},z_{j}\mid z)\;T_{\Delta}(\theta_{i}(x_{i},z_{i}),\theta_{j}(x_{j},z_{j}),z)$
$\leq$ $\sum_{x_{i},x_{j},z_{i},z_{j}\in\Sigma}\Pr(z_{i}\mid z)\Pr(z_{j}\mid z)\;\Pr(x_{i},x_{j}\mid z)\;T_{\Delta}(\theta_{i}(x_{i},z_{i}),\theta_{j}(x_{j},z_{j}),z)+\epsilon|\Sigma|^{2}+(\epsilon_{i}+\epsilon_{j})|\Sigma|^{3}.$ (6)
Finally, combining Eq. 2 and Eq. 6,
$S_{i}(\theta_{i},\theta_{j})$
$\leq$ $\sum_{z\in\Sigma}\Pr(Z=z)\sum_{z_{i},z_{j}\in\Sigma}\Pr(z_{i}\mid z)\Pr(z_{j}\mid z)\sum_{x_{i},x_{j}\in\Sigma}\left(\Pr(x_{i},x_{j}\mid z)-\Pr(x_{i}\mid z)\Pr(x_{j}\mid z)\right)$
$\cdot T_{\Delta}(\theta_{i}(x_{i},z_{i}),\theta_{j}(x_{j},z_{j}),z)+(\epsilon+\epsilon_{i}+\epsilon_{j})|\Sigma|^{2}+(\epsilon_{i}+\epsilon_{j})|\Sigma|^{3}$ (7)
For simplicity, let $\hat{\epsilon}=(\epsilon+\epsilon_{i}+\epsilon_{j})|\Sigma|^{2}+(\epsilon_{i}+\epsilon_{j})|\Sigma|^{3}$.
$=$ $\sum_{z\in\Sigma}\Pr(Z=z)\sum_{z_{i},z_{j}\in\Sigma}\Pr(z_{i}\mid z)\Pr(z_{j}\mid z)\sum_{x_{i},x_{j}\in\Sigma}\Delta_{x_{i},x_{j},z}\;T_{\Delta}(\theta_{i}(x_{i},z_{i}),\theta_{j}(x_{j},z_{j}),z)+\hat{\epsilon}$
$\leq$ $\sum_{z\in\Sigma}\Pr(Z=z)\sum_{z_{i},z_{j}\in\Sigma}\Pr(z_{i}\mid z)\Pr(z_{j}\mid z)\sum_{x_{i},x_{j}\in\Sigma}\Delta_{x_{i},x_{j},z}\;T_{\Delta}(x_{i},x_{j},z)+\hat{\epsilon}$ (By the design of $T_{\Delta}$)
$=$ $\sum_{z\in\Sigma}\Pr(Z=z)\sum_{x_{i},x_{j}\in\Sigma}\Delta_{x_{i},x_{j},z}\;T_{\Delta}(x_{i},x_{j},z)+\hat{\epsilon}$
$=$ $S_{i}(\tau,\tau)+\hat{\epsilon}.$
Lastly, we show that if agent $i$ plays a naive strategy $\nu_{i}$, the expected score is strictly smaller than truth-telling. Note that because a naive strategy is a convex combination between truth-telling $\tau$ and an uninformative strategy $\mu_{i}$, it suffices to show that the expected score of any uninformative strategy is strictly smaller than truth-telling. Because $\mu_{i}(X_{i},Z_{i})$ is independent of $X_{i}$, we can reoder the summations in Eq. 7 in the following way. Let $\mu=\mu_{i}(X_{i},Z_{i})$.
$S_{i}(\mu_{i},\theta_{j})$
$\leq$ $\sum_{z\in\Sigma}\Pr(Z=z)\sum_{z_{i},z_{j}\in\Sigma}\Pr(z_{i}\mid z)\Pr(z_{j}\mid z)\sum_{x_{j}\in\Sigma}T_{\Delta}(\mu,\theta_{j}(x_{j},z_{j}),z)$
$\sum_{x_{i}\in\Sigma}\left(\Pr(x_{i},x_{j}\mid z)-\Pr(x_{i}\mid z)\Pr(x_{j}\mid z)\right)+\hat{\epsilon}$ (by Eq. 7)
$=$ $\sum_{z\in\Sigma}\Pr(Z=z)\sum_{z_{i},z_{j}\in\Sigma}\Pr(z_{i}\mid z)\Pr(z_{j}\mid z)\sum_{x_{j}\in\Sigma}T_{\Delta}(\mu,\theta_{j}(x_{j},z_{j}),z)\left(\Pr(x_{j}\mid z)-\Pr(x_{j}\mid z)\right)+\hat{\epsilon}$
$=$ $\hat{\epsilon}.$
Note that the expected score under truth-telling is exactly $I_{\text{TVD}}(X_{i};X_{j}\mid Z)$. Therefore, whenever $I_{\text{TVD}}(X_{i};X_{j}\mid Z)>\hat{\epsilon}$, we have $S_{i}(\tau,\tau)>S_{i}(\mu_{i},\theta_{j})$ for any uninformative strategy $\mu_{i}$ and any strategy $\theta_{j}$. This completes the proof. ∎
Proof of Lemma B.1.
$\Pr(X_{i},Z_{i}\mid Z)$ $=\sum_{y\in\Sigma}\Pr(Y=y\mid Z)\Pr(X_{i},Z_{i}\mid Y=y,Z)$ (By the law to total probability)
$=\sum_{y\in\Sigma}\Pr(Y=y\mid Z)\Pr(X_{i}\mid Y=y,Z)\Pr(Z_{i}\mid Y=y,Z)$ ($X_{i}$ and $Z_{i}$ are independent conditioned on $Y$.)
$\geq\sum_{y\in\Sigma}\Pr(Y=y\mid Z)\Pr(X_{i}\mid Y=y,Z)\frac{\Pr(Z_{i}\mid Z)\Pr(Y=y\mid Z)-\epsilon_{i}}{\Pr(Y=y\mid Z)}$ (By 4.1)
$=\Pr(Z_{i}\mid Z)\sum_{y\in\Sigma}\Pr(Y=y\mid Z)\Pr(X_{i}\mid Y=y,Z)-\epsilon_{i}\sum_{y\in\Sigma}\Pr(X_{i}\mid Y=y,Z)$
$=\Pr(Z_{i}\mid Z)\Pr(X_{i}\mid Z)-\epsilon_{i}\sum_{y\in\Sigma}\Pr(X_{i}\mid Y=y,Z).$
Proof of Lemma B.2.
$\left|\Pr(Z_{1},Z_{2}\mid Y,Z)-\Pr(Z_{1}\mid Z)\Pr(Z_{2}\mid Z)\right|$
$=$ $\left|\Pr(Z_{1},Z_{2}\mid Y,Z)-\Pr(Z_{1}\mid Y,Z)\Pr(Z_{2}\mid Y,Z)+\Pr(Z_{1}\mid Y,Z)\Pr(Z_{2}\mid Y,Z)-\Pr(Z_{1}\mid Z)\Pr(Z_{2}\mid Z)\right|$
$\leq$ $\left|\Pr(Z_{1},Z_{2}\mid Y,Z)-\Pr(Z_{1}\mid Y,Z)\Pr(Z_{2}\mid Y,Z)\right|+\left|\Pr(Z_{1}\mid Y,Z)\Pr(Z_{2}\mid Y,Z)-\Pr(Z_{1}\mid Z)\Pr(Z_{2}\mid Z)\right|$ (Triangle inequality)
$\leq$ $\epsilon+\left|\Pr(Z_{1}\mid Y,Z)\Pr(Z_{2}\mid Y,Z)-\Pr(Z_{1}\mid Z)\Pr(Z_{2}\mid Z)\right|.$ (4.2)
Therefore, we only have to bound the second quantity.
$\left|\Pr(Z_{1}\mid Y,Z)\Pr(Z_{2}\mid Y,Z)-\Pr(Z_{1}\mid Z)\Pr(Z_{2}\mid Z)\right|$
$=$ $\left|\Pr(Z_{1}\mid Y,Z)(\Pr(Z_{2}\mid Y,Z)-\Pr(Z_{2}\mid Z))+\Pr(Z_{2}\mid Z)(\Pr(Z_{1}\mid Y,Z)-\Pr(Z_{1}\mid Z))\right|$
$\leq$ $\left|\Pr(Z_{1}\mid Y,Z)(\Pr(Z_{2}\mid Y,Z)-\Pr(Z_{2}\mid Z))\right|+\left|\Pr(Z_{2}\mid Z)(\Pr(Z_{1}\mid Y,Z)-\Pr(Z_{1}\mid Z))\right|$ (Triangle inequality)
$\leq$ $\left|\Pr(Z_{2}\mid Y,Z)-\Pr(Z_{2}\mid Z)\right|+\left|\Pr(Z_{1}\mid Y,Z)-\Pr(Z_{1}\mid Z)\right|$
$=$ $\frac{1}{\Pr(Y\mid Z)}\;\left|\Pr(Z_{2},Y\mid Z)-\Pr(Y\mid Z)\;\Pr(Z_{2}\mid Z)\right|+\frac{1}{\Pr(Y\mid Z)}\left|\Pr(Z_{1},Y\mid Z)-\Pr(Y\mid Z)\;\Pr(Z_{1}\mid Z)\right|$
$\leq$ $\frac{\epsilon_{1}+\epsilon_{2}}{\Pr(Y\mid Z)}.$ (By 4.1)
Combining everything, we complete the proof. ∎
B.2 Necessity of the Assumptions
Proof of 4.4.
Intuitively, if $Z_{i}$ and $Z_{j}$ are correlated in a different way from $X_{i}$ and $X_{j}$, then it makes $Z_{i}$ and $Z_{j}$ no longer “cheap” signals. In this case, a strategy that combines $Z_{i}$ and $X_{i}$ will improve the expected score.
Suppose $\Sigma=\{0,1\}$ contains two signals. Consider the following joint distribution for $\Pr((X_{i},Z_{i}),(X_{j},Z_{j})|Y=y,Z=z)$ where row 1 to 4 corresponds to $(X_{i},Z_{i})=(0,0),(0,1),(1,0),(1,1)$ respectively and the columns stand for $(X_{j},Z_{j})$. For simplicity, let’s suppose the following joint distribution is identical for any $y$ and $z$.
$\Pr((X_{i},Z_{i}),(X_{j},Z_{j})\mid y,z)=\begin{pmatrix}2d&0&0&0\\ 0&0&0&0\\ 0&0&\frac{1}{2}-2d&\frac{1}{2}-2d\\ 0&d&0&d\end{pmatrix}\quad\text{for any $y,z\in\{0,1\}$.}$
Note that the example is valid when $0<d\leq\frac{1}{4}$. Using this joint distribution, we can further write down the joint distribution between other pairs of variables.
$\Pr(X_{i},X_{j}\mid y,z)=\begin{pmatrix}2d&0\\ d&1-3d\end{pmatrix}\qquad\Pr(Z_{i},Z_{j}\mid y,z)=\begin{pmatrix}\frac{1}{2}&\frac{1}{2}-2d\\ 0&2d\end{pmatrix}\quad\text{for any $y,z\in\{0,1\}$.}$
Furthermore, the product of the marginal distributions are
$\Pr(X_{i}\mid y,z)\Pr(X_{j}\mid y,z)$ $=\begin{pmatrix}6d^{2}&2d\left(1-3d\right)\\ 3d\left(1-2d\right)&\left(1-2d\right)\left(1-3d\right)\end{pmatrix}$
$\Pr(Z_{i}\mid y,z)\Pr(Z_{j}\mid y,z)$ $=\begin{pmatrix}\frac{1}{2}\left(1-2d\right)&\frac{1}{2}\left(1-2d\right)\\ d&d\end{pmatrix}\quad\text{for any $y,z\in\{0,1\}$.}$
We can further write down the difference between the joint distribution and the product of the marginal distributions.
$\Pr(X_{i},X_{j}\mid y,z)-\Pr(X_{i}\mid y,z)\Pr(X_{j}\mid y,z)$ $=\begin{pmatrix}2d(1-3d)&-2d(1-3d)\\ -2d(1-3d)&2d(1-3d)\end{pmatrix}$
$\Pr(Z_{i},Z_{j}\mid y,z)-\Pr(Z_{i}\mid y,z)\Pr(Z_{j}\mid y,z)$ $=\begin{pmatrix}d&-d\\ -d&d\end{pmatrix}$
Note that in this example, $|\Pr(Z_{i},Z_{j}\mid y,z)-\Pr(Z_{i}\mid y,z)\Pr(Z_{j}\mid y,z)|=d$ for any $Z_{i},Z_{j}\in\{0,1\}$, satisfying the condition in the proposition. Furthermore, for any $k\in\{0,1\}$, the scoring function $T_{\Delta}(k,h,l)=1$ if and only if $h=l$ and 0 otherwise. Now, we compute the expected score when both agents honestly report the high-effort signal.
$S_{i}(\tau,\tau)=$ $\sum_{z\in\{0,1\}}\Pr(Z=z)\sum_{x_{i},x_{j}\in\{0,1\}}(\Pr(X_{i}=x_{i},X_{j}=x_{j}\mid z)-\Pr(X_{i}=x_{i}\mid z)\;\Pr(X_{j}=x_{j}\mid z))\;T_{\Delta}(x_{i},x_{j},z)$
$=\sum_{x_{i},x_{j}\in\{0,1\}}\left|\Pr(x_{i},x_{j}\mid z)-\Pr(x_{i}\mid z)\;\Pr(x_{j}\mid z)\right|$
$=8d(1-3d).$
We show that this score is strictly smaller than the expected score when agents report 00 if and only if $X_{i}=Z_{i}=0$ (and $X_{j}=Z_{j}=0$) and report $1$ otherwise. Let $\theta$ be the strategy such that $\theta(k,l|z)=0$ if and only if $k=l=0$ for any $z$. We can write down the joint distribution and the product of marginal distributions for $\hat{X}_{i}$ and $\hat{X}_{j}$.
$\Pr(\hat{X}_{i},\hat{X}_{j}\mid y,z)=\begin{pmatrix}2d&0\\ 0&1-2d\end{pmatrix}\qquad\Pr(\hat{X}_{i},\hat{X}_{j}\mid y,z)=\begin{pmatrix}4d^{2}&2d(1-2d)\\ 2d(1-2d)&(1-2d)^{2}\end{pmatrix}\quad\text{for any $y,z\in\{0,1\}$.}$
The difference between the above two matrices is thus
$\Pr(\hat{X}_{i},\hat{X}_{j}\mid y,z)-\Pr(\hat{X}_{i}\mid y,z)\Pr(\hat{X}_{j}\mid y,z)=\begin{pmatrix}2d(1-2d)&-2d(1-2d)\\ -2d(1-2d)&2d(1-2d)\end{pmatrix}\quad\text{for any $y,z\in\{0,1\}$.}$
This means that the expected score under the strategy profile $(\theta,\theta)$ is
$S_{i}(\theta,\theta)=$ $\sum_{z\in\{0,1\}}\Pr(Z=z)\sum_{x_{i},x_{j}\in\{0,1\}}(\Pr((\hat{X}_{i}=x_{i},\hat{X}_{j}=x_{j}\mid z)-\Pr(\hat{X}_{i}=x_{i}\mid z)\;\Pr(\hat{X}_{j}=x_{j}\mid z))\;T_{\Delta}(x_{i},x_{j},z)$
$=\sum_{x_{i},x_{j}\in\{0,1\}}\left|\Pr(\hat{X}_{i}=x_{i},\hat{X}_{j}=x_{j}\mid z)-\Pr(\hat{X}_{i}=x_{i}\mid z)\;\Pr(\hat{X}_{j}=x_{j}\mid z)\right|$
$=8d(1-2d)$
$=S_{i}(\tau,\tau)+8d^{2}.$
B.3 Information Monotonicity under Naive Strategies
Proof of 4.6.
A naive strategy is a convex combination of truth-telling and an uninformative strategy. Therefore, it suffices to show that the expected score of truth-telling is strictly higher than that of any uninformative strategy.
An uninformative strategy is essentially a mixture of two types of strategy. First, $\mu_{i}(X_{i},Z_{i})$ is independent of both $X_{i}$ and $Z_{i}$. In this case, we have shown in the proof of Theorem 4.3 that the expected score is 0, strictly less than the expected score of truth-telling, which is $I_{\text{TVD}}(X_{i};X_{j}\mid Z)$.
Second, $\mu_{i}(X_{i},Z_{i})$ depends only on $Z_{i}$; let $\mu_{i}(X_{i},Z_{i})=\hat{Z}_{i}$. Conditioned on $Z$, the Markov chain $Z_{i}\rightarrow\hat{Z}_{i}$ holds. Now, fix agent $j$’s naive strategy and let her report be $\hat{X}_{j}$. By the data processing inequality [Shannon, 1948], we have:
$I_{\text{TVD}}(\hat{Z}_{i};\hat{X}_{j}\mid Z)\leq I_{\text{TVD}}(Z_{i};\hat{X}_{j}\mid Z).$
Next, suppose agent $i$ reporting $Z_{i}$, which is the score-maximizing naive strategy regardless of agent $j$’s strategy. For any naive strategy such that $\hat{X}_{j}=\nu_{j}(X_{j},Z_{j})$ that agent $j$ plays, applying the data processing inequality to agent $j$’s report yields:
$I_{\text{TVD}}(Z_{i};\hat{X}_{j}\mid Z)\leq\max\left(I_{\text{TVD}}(Z_{i};Z_{j}\mid Z),I_{\text{TVD}}(Z_{i};X_{j}\mid Z)\right).$
Therefore, 4.5 is sufficient for information monotonicity under naive strategies. ∎
B.4 Learning the Scoring Function
Proof of 4.7.
Note that when agent $i$ and $j$ play strategy $\theta_{i}$ and $\theta_{j}$ respectively and the scoring function is $T$,
$S_{i}^{T}(\theta_{i},\theta_{j})$
$=$ $\sum_{z\in\Sigma}\Pr(Z=z)\sum_{x_{i},x_{j},z_{i},z_{j}\in\Sigma}(\Pr(x_{i},x_{j},z_{i},z_{j}\mid z)-\Pr(x_{i},z_{i}\mid z)\;\Pr(x_{j},z_{j}\mid z))\;T(\theta_{i}(x_{i},z_{i}),\theta_{j}(x_{j},z_{j}),z)$
Because 4.2 and 4.1 hold with no error, by Lemma B.2 and Lemma B.1, it is easy to show that $\Pr(x_{i},x_{j},z_{i},z_{j}\mid z)=\Pr(x_{i},x_{j}\mid z)\;\Pr(z_{i}\mid z)\;\Pr(z_{j}\mid z)$ and $\Pr(x_{i},z_{i}\mid z)=\Pr(x_{i}\mid z)\;\Pr(z_{i}\mid z)$.
$=$ $\sum_{z\in\Sigma}\Pr(Z=z)\sum_{z_{i},z_{j}\in\Sigma}\Pr(z_{i}\mid z)\;\Pr(z_{j}\mid z)$
$\sum_{x_{i},x_{j}\in\Sigma}(\Pr(x_{i},x_{j}\mid z)-\Pr(x_{i}\mid z)\;\Pr(x_{j}\mid z))\;T(\theta_{i}(x_{i},z_{i}),\theta_{j}(x_{j},z_{j}),z)$
$\leq$ $\sum_{z\in\Sigma}\Pr(Z=z)\sum_{z_{i},z_{j}\in\Sigma}\Pr(z_{i}\mid z)\;\Pr(z_{j}\mid z)\sum_{x_{i},x_{j}\in\Sigma}|\Pr(x_{i},x_{j}\mid z)-\Pr(x_{i}\mid z)\;\Pr(x_{j}\mid z)|$ (Because $T(k,h,l)$ is either 0 or 1.)
$=$ $\sum_{z\in\Sigma}\Pr(Z=z)\sum_{x_{i},x_{j}\in\Sigma}|\Pr(x_{i},x_{j}\mid z)-\Pr(x_{i}\mid z)\;\Pr(x_{j}\mid z)|$ (Marginalize $z_{i},z_{j}$)
$=$ $S_{i}^{T_{\Delta}}(\tau,\tau)$
Appendix C A Heuristic Generalization to Complex Signals
We have focused on scenarios where the signal space is small, such as answering multi-choice questions like sentiment labeling, preference elicitation, or grading. However, many real-world crowdsourcing tasks involve open-ended questions that require workers to provide more complex responses, such as image or video descriptions, idea generation, or feedback and review collection. These tasks introduce unique challenges for applying our method, including 1) defining meaningful representation (textual) responses, and 2) computing the information between two responses conditioned on the principal’s sample $Z$. In this section, we outline how to address these challenges and extend our method to accommodate tasks with complex, open-ended responses. We will illustrate our idea using peer review as an example, while we note that our method can be generalized to other similar settings.
Representing responses.
Suppose two agents review the same paper and each has a review after reading the paper, which is denoted as $X_{i}$ and $X_{j}$ respectively. Instead of working hard to obtain a high-quality review, each agent can use an LLM to generate a fake review, denoted as $Z_{i}$ and $Z_{j}$. The first challenge is how to measure the level of agreement between two reviews. The idea is to represent each review with an embedding vector and compute the similarity between a pair of embedding vectors. Depending on the information of interest, different embedding methods can be applied. For example, classic NLP embedding methods such as Word2Vec [Mikolov, 2013], TF-IDF [Manning, 2009], BERT and its variants [Devlin, 2018, Sanh, 2019], and GPT-3 [Brown et al., 2020] can be used to represent the language-level information within the responses.
Existing embedding methods may struggle when the focus is on content-level information, such as the judgments in peer reviews. For example, at the language level, an LLM-generated review may be similar to a human-written review that is rewritten for polishing by an LLM, while the information within their judgments can be very different. To address this, we consider an alternative, interpretable approach for representing complex responses. The key idea is to prompt an LLM to score a response based on a list of pre-determined characteristics.
For instance, we can prompt an LLM to classify whether a review adopts a compromising, critical, or neutral stance on specific aspects of a paper, such as writing quality, completeness of related work, novelty of the research idea, or the quality of datasets. Using this method, each textual review $X$ can be represented as a vector $\phi(X)$ of length $K$, where each entry $\phi(X)_{k}$ encodes the attitude of the review toward the $K$-th aspect. The similarity between two reviews $X_{i}$ and $X_{j}$ can then be quantified using the cosine similarity between their representations $\phi(X_{i})$ and $\phi(X_{j})$.
The same idea can be generalized beyond peer review. Suppose the principal has a pre-defined set of questions about the responses, denoted as $\mathcal{Q}$ where $|\mathcal{Q}|=K$. Each question requires a (normalized) score between $-1$ and $1$, ensuring that the inner product between vectors can serve as a meaningful similarity measure. Given a response $X_{i}$, the principal can prompt an LLM to iteratively get an answer to each question $q\in\mathcal{Q}$. This process converts each response $X_{i}$ into a vector $\phi(X_{i})$ of length $K$, and each entry is a score between $-1$ and $1$. Similar ideas have been applied to identify similar data points [Zeng et al., 2024], analyze peer review [Liang et al., 2024], and select features [Jeong et al., 2024].
Computing the conditioned similarity.
The key idea of the conditioned CA mechanism is to give an agent $i$ a higher score for having responses that correlate with another agent $j$’s responses beyond what both share with $Z$. To measure how $\phi(X_{i})$ and $\phi(X_{j})$ correlate conditioned on $\phi(Z)$, a direct application — treating each vector component as a random variable — quickly becomes infeasible due to the exponential growth of the signal space in the vector dimension $K$. Therefore, we adopt a heuristic: we compute the similarity between $\phi(X_{i})$ and $\phi(X_{j})$ after removing their shared direction with $\phi(Z)$. Specifically:
  1. 1.
    Project each vector onto the hyperplane orthogonal to $\phi(Z)$.
    Let $u_{z}=\frac{\phi(Z)}{\lVert\phi(Z)\rVert}$ be the unit vector in the direction of $\phi(Z)$. Then,
    $\phi(X_{i})^{\prime}=\phi(X_{i})-(\phi(X_{i})\cdot u_{z})u_{z},\qquad\phi(X_{j})^{\prime}=\phi(X_{j})-(\phi(X_{j})\cdot u_{z})u_{z}.$
  2. 2.
    Compute the cosine similarity between $\phi(X_{i})^{\prime}$ and $\phi(X_{j})^{\prime}$ to score agent $i$.
By projecting the embedding vectors onto the orthogonal complement of $\phi(Z)$, we remove the information that already exists in $Z$. This projection approach is also used in NLP for “debiasing” word embeddings by eliminating similarity along a “biased direction” [Bolukbasi et al., 2016].
Finally, we note that we do not assume the principal’s questions in $\mathcal{Q}$ to be i.i.d. Thus, the theoretical guarantees in Section 4.1 do not directly apply to this heuristic generalization; we instead rely on empirical evaluation to assess performance.
Appendix D Experiments (Continue)
D.1 Assumption Verifications (Continue)
Figure 5 present the TVD mutual information between LLM samples and human responses conditioned on (different) LLM samples, $I_{\text{TVD}}(Z_{i};X_{j}\mid Z)$, which is the analogous results to Fig. 2 on the remaining datasets.
(a) Hatefulness
(b) Offensiveness
Figure 5: Analogous figures to Fig. 2 for other datasets.
Table 2: TVD-Conditioned Mutual Information Between Two LLMs on the Preference Alignment Data
GPT-3.5 GPT-4 Gemma Mistral Phi
GPT-3.5 0.2255 (0.0412) 0.1973 (0.0694) 0.0865 (0.1802) 0.1224 (0.1443) 0.1311 (0.1355)
GPT-4 0.0715 (0.1537) 0.2559 (-0.0307) 0.0571 (0.1682) 0.0552 (0.1701) 0.0723 (0.1530)
Gemma 0.1013 (0.1819) 0.1767 (0.1066) 0.2453 (0.0379) 0.0655 (0.2177) 0.1174 (0.1659)
Mistral 0.0920 (0.1938) 0.1132 (0.1726) 0.0620 (0.2238) 0.2350 (0.0508) 0.0964 (0.1894)
Phi 0.0874 (0.1969) 0.1300 (0.1542) 0.0855 (0.1988) 0.0929 (0.1914) 0.2479 (0.0364)
Each entry represents $I_{\text{TVD}}(Z_{i};Z_{j}\mid Z)$ (Difference: $I_{\text{TVD}}(X_{i};X_{j}\mid Z)-I_{\text{TVD}}(Z_{i};Z_{j}\mid Z)$), where $Z_{i}$ and $Z_{j}$ are generated by the models in the row and column respectively, and $Z$ uses the same model as $Z_{i}$.
Table 2, Table 3, Table 4 and Table 5 present the mutual information $I_{\text{TVD}}(Z_{i};Z_{j}\mid Z)$ on four datasets. $Z$ and $Z_{i}$ are sampled by the same LLM indicated on the rows and $Z_{j}$ is sampled by the LLM indicated on the columns. The first observation is that all values displayed in black exceed $0.05$, sometimes significantly so, which represents the mutual information of the random reporting baseline. This implies that $I_{\text{TVD}}(Z_{i};Z_{j}\mid Z)>\epsilon$ holds consistently, rejecting hypothesis 2.
Regarding hypothesis 4, we observe that the average correlations between two LLMs are generally weaker than those between two human labelers, provided $Z_{i}$ and $Z$ are independently generated by the same model. This trend is evident in the tables, where almost all bracketed values are positive, with one exception.***The exception occurs in the preference alignment dataset, where samples of $Z_{i},Z_{j}$, and $Z$ all generated by GPT-4.
Table 3: TVD-Conditioned Mutual Information Between Two LLMs on the Hatefulness Labeling Data
GPT-3.5 GPT-4 Gemma Mistral Phi
GPT-3.5 0.0317 (0.1605) 0.0362 (0.1560) 0.0385 (0.1537) 0.0426 (0.1496) 0.0353 (0.1570)
GPT-4 0.0077 (0.1575) 0.0286 (0.1366) 0.0229 (0.1424) 0.0216 (0.1437) 0.0253 (0.1400)
Gemma 0.0075 (0.1588) 0.0183 (0.1481) 0.0758 (0.0905) 0.0464 (0.1199) 0.0543 (0.1121)
Mistral 0.0040 (0.1600) 0.0097 (0.1542) 0.0341 (0.1299) 0.0408 (0.1231) 0.0350 (0.1290)
Phi 0.0088 (0.1658) 0.0115 (0.1631) 0.0618 (0.1128) 0.0526 (0.1220) 0.0803 (0.0943)
Table 4: TVD-Conditioned Mutual Information Between Two LLMs on the Offensiveness Labeling Data
GPT-3.5 GPT-4 Gemma Mistral Phi
GPT-3.5 0.1267 (0.0809) 0.0926 (0.1149) 0.0450 (0.1625) 0.0785 (0.1290) 0.0802 (0.1273)
GPT-4 0.0419 (0.1092) 0.0920 (0.0590) 0.0520 (0.0990) 0.0867 (0.0644) 0.0794 (0.0716)
Gemma 0.0302 (0.2336) 0.0602 (0.2036) 0.1152 (0.1486) 0.0597 (0.2040) 0.0676 (0.1962)
Mistral 0.0390 (0.1407) 0.0573 (0.1225) 0.0295 (0.1502) 0.0836 (0.0961) 0.0516 (0.1281)
Phi 0.0320 (0.1604) 0.0597 (0.1327) 0.0334 (0.1591) 0.0604 (0.1321) 0.0651 (0.1273)
Table 5: TVD-Conditioned Mutual Information Between Two LLMs on the Toxicity Labeling Data
GPT-3.5 GPT-4 Gemma Mistral Phi
GPT-3.5 0.1611 (0.0063) 0.1328 (0.0347) 0.1199 (0.0476) 0.1419 (0.0255) 0.1340 (0.0334)
GPT-4 0.0587 (0.1399) 0.1035 (0.0951) 0.0755 (0.1231) 0.0673 (0.1312) 0.0652 (0.1334)
Gemma 0.0721 (0.1650) 0.0869 (0.1502) 0.1329 (0.1042) 0.0880 (0.1491) 0.1043 (0.1328)
Mistral 0.0602 (0.1413) 0.0622 (0.1394) 0.0454 (0.1561) 0.0934 (0.1081) 0.0642 (0.1373)
Phi 0.0416 (0.1756) 0.0517 (0.1655) 0.0382 (0.1790) 0.0472 (0.1701) 0.0666 (0.1507)
D.2 Detecting Cheating Agents (Continue)
D.2.1 Additional Figures and Tables
In Fig. 6, we present additional figures on the toxicity dataset, analogous to Fig. 4(a), for other types of LLM cheaters. Analogous results to Table 1 for other datasets are summarized in Table 6, Table 7, Table 8, and Table 9. These results support our main claim that, although our method is occasionally slightly outperformed by certain baselines, it consistently demonstrates the most robust performance across all tested settings.
(a) GPT-4
(b) GPT-3.5-turbo
(c) Gemma-2-2b-it
(d) Mistral-7B-Instruct-v0.3
(e) Phi-3.5-mini-Instruct
Figure 6: The area under the ROC curve (AUC) for various methods while varying the fraction of LLM cheaters. These are the analogous figures for Fig. 4(a) for the remaining LLMs on the toxicity labeling dataset.
Table 6: Average AUC (and bottom 10% quantiles) across datasets and methods with GPT-3.5 cheaters.
Dataset OA CA OAZ EM CAZ (ours)
Hatefulness 0.61 (0.40) 0.90 (0.84) 0.90 (0.85) 0.86 (0.77) 0.89 (0.86)
Offensiveness 0.80 (0.69) 0.76 (0.58) 0.48 (0.27) 0.76 (0.57) 0.83 (0.77)
Toxicity 0.75 (0.62) 0.72 (0.51) 0.47 (0.24) 0.71 (0.50) 0.78 (0.71)
Preference Alignment 0.88 (0.67) 0.86 (0.69) 0.88 (0.76) 0.96 (0.91) 0.91 (0.83)
Table 7: Average AUC (and bottom 10% quantiles) across datasets and methods with Mistral cheaters.
Dataset OA CA OAZ EM CAZ (ours)
Hatefulness 0.82 (0.75) 0.72 (0.51) 0.62 (0.48) 0.76 (0.54) 0.84 (0.78)
Offensiveness 0.71 (0.54) 0.79 (0.62) 0.63 (0.47) 0.75 (0.57) 0.89 (0.84)
Toxicity 0.65 (0.50) 0.75 (0.56) 0.58 (0.39) 0.70 (0.49) 0.84 (0.79)
Preference Alignment 0.86 (0.66) 0.89 (0.75) 0.94 (0.87) 0.94 (0.87) 0.93 (0.88)
Table 8: Average AUC (and bottom 10% quantiles) across datasets and methods with Gemma cheaters.
Dataset OA CA OAZ EM CAZ (ours)
Hatefulness 0.85 (0.79) 0.71 (0.51) 0.60 (0.42) 0.77 (0.56) 0.85 (0.80)
Offensiveness 0.75 (0.58) 0.78 (0.59) 0.61 (0.43) 0.76 (0.53) 0.90 (0.84)
Toxicity 0.70 (0.58) 0.75 (0.59) 0.56 (0.36) 0.72 (0.52) 0.85 (0.79)
Preference Alignment 0.80 (0.58) 0.87 (0.73) 0.76 (0.62) 0.92 (0.85) 0.92 (0.86)
Table 9: Average AUC (and bottom 10% quantiles) across datasets and methods with Phi cheaters.
Dataset OA CA OAZ EM CAZ (ours)
Hatefulness 0.87 (0.81) 0.67 (0.44) 0.66 (0.47) 0.77 (0.56) 0.85 (0.79)
Offensiveness 0.76 (0.64) 0.76 (0.59) 0.58 (0.35) 0.75 (0.53) 0.88 (0.83)
Toxicity 0.73 (0.62) 0.74 (0.54) 0.54 (0.35) 0.71 (0.48) 0.83 (0.77)
Preference Alignment 0.85 (0.68) 0.82 (0.63) 0.82 (0.63) 0.97 (0.94) 0.93 (0.88)
D.2.2 Detection Performance vs. Number of Samples (Continue)
Figure 7 presents similar results to Fig. 4(b) for the hatefulness, offensiveness, and toxicity labeling datasets.
(a) Hatefulness
(b) Offensiveness
(c) Toxicity
Figure 7: The area under the ROC curve (AUC) for various methods while varying the fraction of questions the principal can sample.
D.2.3 Conditioning on samples of multiple LLMs
In practice, agents may employ a variety of LLMs for low-effort response generation, and the principal lacks prior knowledge of which models they are using. To address this challenge, we propose an iterative conditioning approach, systematically removing the influence of responses generated by each LLM available to the principal. The idea is to iteratively compute the CA score conditioned on each of the LLM samples the principal has, and then take the minimum as the final score. Intuitively, human agents will perform consistently well when conditioning out any of the LLMs’ information. However, LLM cheaters will perform poorly in at least one of the iterations and thus lead to a lower final score.
In Fig. 8, we replicate Fig. 4(a) for the setting where the cheaters randomly pick two LLMs to cheat with half-half probability while the principal iteratively conditions out his samples for these two LLMs’ answers. Our method exhibits the most robust performance across all settings, echoing the previous conclusion.
We further note that the problem of LLM-assisted cheating is less of a concern when each cheater uses a unique LLM. This is because LLM responses are less correlated with each other and can be easily detected as outliers under classic crowdsourcing evaluation methods, such as the EM algorithm under the DS model.
(a) Mistral & GPT-3.5
(b) Mistral & GPT-4
(c) GPT-3.5 & GPT-4
Figure 8: Area under the ROC curve (AUC) for various methods as the fraction of LLM cheaters increases. In each case, LLM cheaters randomly choose between two LLMs to generate responses, and the score is defined as the minimum of the two conditioned CA scores, each conditioned on samples from one of the two LLMs.
D.3 Complex Signals
We further test the idea illustrated in Appendix C that generalizes our method to high-dimensional settings.
In this experiment, we randomly sample 500 papers from the ICLR 2024 OpenReview data and select three human reviews per paper as human responses $X$. For each of the same batch of papers, we further use each of the five LLMs to generate three AI reviews using three similar-meaning yet different prompts (Appendix E). We randomly pick one of the generated reviews as the principal’s sample $Z$, and the other two generated reviews will be used to simulate LLM-cheaters’ responses.
We prompt GPT-4 to process each human review and LLM-generated review into a $19$-dimension vector with the prompt shown in Appendix E. We further simulate a review embedding tensor with a size of $n\times m\times 19$, where there are $n=188$ reviewers and $m=500$ papers in total. Each reviewer randomly reviews $k\in\{5,8\}$ papers. An $\alpha_{llm}$ fraction of reviewers are LLM cheaters whose review vectors are replaced with a particular type of LLM-generated reviews. An $\alpha_{p}$ fraction of reviewers are prior cheaters, where the score in each entry of the review vector is randomly sampled according to the marginal distribution in the dataset. An $\alpha_{r}$ fraction of reviewers are random cheaters, where the score in each entry of the review vector is sampled uniformly at random. In our experiments, we independently sample each of $\alpha_{llm}$, $\alpha_{p}$, and $\alpha_{r}$ uniformly from the interval $[0,0.1]$ and take average over $100$ trials. While constructing the review embedding tensor, we guarantee that each paper has approximate $3$ reviews so that the variance of the number of reviewers per paper is controlled. Again, we consider five types of LLMs and for each type of LLM cheater, the principal’s sample $Z$ is an independent generation of the same LLM under a different prompt.
We consider three methods to score the reviewers.
  • Cosine Similarity: Scores each reviewer based on the average cosine similarity between their review vector and the review vectors of other reviewers for the same paper.
  • Projected Cosine Similarity: Applies the method described in Appendix C, where each review vector is projected onto the subspace orthogonal to the principal reviewer’s sampled embedding before computing similarity.
  • Negative Similarity to $Z$: Scores each reviewer by taking the negative average cosine similarity between their review vector and the principal reviewer’s sampled vector $Z$.
We present the results in Fig. 9. It is clear that conditioning out the principal’s sampled LLM reviews is more effective than computing the cosine similarity directly. Furthermore, naively removing reviews that are similar to $Z$ has a large variance, whose performance greatly depends on the existence of random and prior cheaters.
We further note that there is a big room for improvement. First, from a methodology perspective, the 19-dimensional embedding is heuristically chosen and is not optimized for distinguishing human-AI reviews. Second, from the data perspective, human reviews are naturally very noisy. Our method can be potentially applied to other crowdsourcing tasks where human responses are more consistent. We view these as future directions.
(a) $k=5$ reviews per reviewer
(b) $k=8$ reviews per reviewer
Figure 9: A comparison of the area under the ROC curve (AUC) of various methods in the peer review dataset.
Appendix E Prompts Templates
The prompt templates for two different dataset settings are provided in Table E and Table E, respectively.
Prompt Template for Toxicity Labeling System Prompt: As a helpful system judgment assistant, your task is to answer the following questions for the given comment. You should label a comment as hateful if it expresses hatred towards a targeted group or is intended to be derogatory, to humiliate, or to insult the members of the group. You should label a comment as offensive if it contains hurtful, derogatory, or obscene comments. You should label a comment as toxic if it is a rude, disrespectful, or unreasonable comment that is likely to make readers want to leave a discussion. Question 1: Do you think this comment is hateful? — Yes or No. Question 2: Do you think this comment is offensive? — Yes or No. Question 3: Do you think this comment is toxic? — Yes or No. Comment: [Comment] User Prompt: Now, please evaluate the following comment and return the final answers in JSON format:
"question 1": {"hateful": <yes/no>},
"question 2": {"offensive": <yes/no>},
"question 3": {"toxic": <yes/no>}
Prompt Template for Preference Alignment System Prompt:
Please act as an impartial judge and evaluate the quality of the responses provided by two AI assistants to the user question displayed below. You should choose the assistant that follows the user’s instructions and answers the user’s question better. Your evaluation should consider factors such as the helpfulness, harmlessness, truthfulness, and level of detail of their responses. Begin your evaluation by comparing the two responses and provide a short explanation. Avoid any position biases and ensure that the order in which the responses were presented does not influence your decision. Do not allow the length of the responses to influence your evaluation. Do not favor certain names of the assistants. Be as objective as possible.
After providing your explanation, output your final verdict by strictly following a 5-point Likert scale: A is clearly better A is slightly better Tie B is slightly better B is clearly better Question: [Question] Response A: [Response A] Response B: [Response B] User Prompt: Now, please evaluate the responses and return the final verdict in JSON format:
    "final verdict": <A-is-clearly-better / A-is-slightly-better, Tie,
    B-is-slightly-better / B-is-clearly-better>,
    "short explanation": <your short explanation within one sentence>
Prompt Template for Paper Review 19-Dimensional Judgment Embedding Generation System Prompt:
You will be given a review of a computer science paper. Such a review is typically composed of a summary of the corresponding paper, reasons to accept, reasons to reject, and questions for the authors. Upon receiving the review, your task is to generate an embedding of the review in a 19-dimensional space according to the following instructions:
1. Identify and ignore the summary and the questions section of the review (if any). 2. Separate the review into advantages and disadvantages. 3. Further decompose the advantages and disadvantages into independent judgments. 4. You will be given 19 judgment categories. For each category: If a positive-toned judgment falls into the category, assign a score of 1. If a negative-toned judgment falls into the category, assign a score of -1. If no judgment is related to the category, assign a score of 0. 5. Return the scores for all 19 categories as a dictionary. Do not include any additional explanation. The judgment categories are: Clarity and Presentation Related Work Theory Soundness Experiment Design Significance of the Problem Novelty of Method Practicality of Method Comparison to Previous Studies Implications of the Research Reproducibility Ethical Aspects Relevance to Conference Experiment Execution Scope and Generalizability Utility and Applicability Defense Effectiveness Real-world Applicability Algorithm Efficiency Data Generation and Augmentation User Prompt: Now, please evaluate the paper review and return the scores for each category in the following JSON format:
    "Clarity and Presentation": <-1/0/1>,
    "Related Work": <-1/0/1>,
    "Theory Soundness": <-1/0/1>,
    "Experiment Design": <-1/0/1>,
    "Significance of the Problem": <-1/0/1>,
    "Novelty of Method": <-1/0/1>,
    "Practicality of Method": <-1/0/1>,
    "Comparison to Previous Studies": <-1/0/1>,
    "Implications of the Research": <-1/0/1>,
    "Reproducibility": <-1/0/1>,
    "Ethical Aspects": <-1/0/1>,
    "Relevance to Conference": <-1/0/1>,
    "Experiment Execution": <-1/0/1>,
    "Scope and Generalizability": <-1/0/1>,
    "Utility and Applicability": <-1/0/1>,
    "Defense Effectiveness": <-1/0/1>,
    "Real-world Applicability": <-1/0/1>,
    "Data Generation and Augmentation": <-1/0/1>,
    "Algorithm Efficiency": <-1/0/1>

### v2

@@
-Evaluating LLM-corrupted Crowdsourcing Data Without Ground Truth
+Evaluating LLM-Contaminated Crowdsourcing Data Without Ground Truth
 Abstract
@@
 However, the increasing use of large language models (LLMs) by crowdsourcing workers poses a significant challenge: datasets intended to reflect human input may be compromised by LLM-generated responses.
-Existing LLM detection approaches often rely on high-dimension training data such as text, making them unsuitable for annotation tasks like multiple-choice labeling.
-In this work, we investigate the potential of peer prediction — a mechanism that evaluates the information within workers’ responses without using ground truth — to mitigate LLM-assisted cheating in crowdsourcing with a focus on annotation tasks.
+Existing LLM detection approaches often rely on high-dimensional training data such as text, making them unsuitable for structured annotation tasks like multiple-choice labeling.
+In this work, we investigate the potential of peer prediction—a mechanism that evaluates the information within workers’ responses—to mitigate LLM-assisted cheating in crowdsourcing with a focus on annotation tasks.
 Our approach quantifies the correlations between worker answers while conditioning on (a subset of) LLM-generated labels available to the requester.
-Building on prior research, we propose a training-free scoring mechanism with theoretical guarantees under a crowdsourcing model that accounts for LLM collusion.
+Building on prior research, we propose a training-free scoring mechanism with theoretical guarantees under a novel model that accounts for LLM collusion.
 We establish conditions under which our method is effective and empirically demonstrate its robustness in detecting low-effort cheating on real-world crowdsourcing datasets.
 1 Introduction
-High-quality human feedback plays a crucial role in shaping modern AI systems, such as ChatGPT, to be trustworthy, safe, and aligned with human preference (Ouyang et al., 2022).
-This feedback is usually collected through crowdsourcing, where a requester (principal) posts a set of tasks and recruits workers (agents) — who may be trained experts or crowdsourcing workers — to complete them.
-For example, researchers widely use platforms like Amazon Mechanical Turk and Prolific for tasks such as data labeling and survey administration.
+High-quality human feedback plays a crucial role in shaping modern AI systems, such as ChatGPT, to be trustworthy, safe, and aligned with human preference ouyang2022training.
+This feedback is usually collected through crowdsourcing, where a requester posts tasks and recruits trained experts or general participants to complete them.
+For example, platforms like Amazon Mechanical Turk and Prolific are widely used by researchers for tasks such as data labeling and survey administration.
 However, the integrity of this data collection process is increasingly threatened by the very AI models it seeks to improve.
-Evidence suggests that many agents are using large language models (LLMs) to complete tasks such as abstract summarization (Veselovsky et al., 2023) and peer review (Liang et al., 2024), leading to datasets that no longer reflect genuine human input.
-While LLM-assisted annotations or responses may sometimes be more accurate or coherent than human’s (Gilardi et al., 2023; Törnberg, 2023; Kaikaus et al., 2023), their widespread use clashes with the goal of gathering diverse and authentic human opinions.
+Evidence suggests that many agents rely on large language models (LLMs) to complete tasks such as abstract summarization (veselovsky2023artificial) and peer review (liang2024mapping), leading to datasets that no longer reflect genuine human input, a phenomenon known as LLM contamination.
+While LLM-assisted annotations or responses may sometimes be more accurate or coherent than human’s (Gilardi_2023; törnberg2023chatgpt4; 10386425), their widespread use undermines the fundamental goal of gathering diverse and authentic human opinions.
 This issue is particularly concerning in applications that depend on human subjectivity or expertise, such as preference labeling and toxicity detection, where human judgment remains the gold standard.
-How to detect potential LLM cheaters and evaluate the quality of a human-contributed dataset?
-Existing black-box methods for detecting LLM-generated content primarily focus on distinguishing human-written text from AI-generated text by analyzing statistical differences in writing style (Gehrmann et al., 2019; Mao et al., 2024; Koike et al., 2024; Hu et al., 2023).
-While sometimes effective in open-ended text generation tasks, these training-based approaches are not well-suited for annotation tasks, where the discrete responses lack the rich textual features needed for classification.
-In this paper, we develop a theoretically robust scoring metric to evaluate the information in human contributions that is not captured by LLM responses.
-Our method builds on peer prediction, a scoring mechanism designed to incentivize truthful reporting in the absence of ground truth (Miller et al., 2005; Shnayder et al., 2016).
+How to detect potential LLM-misuse and evaluate the quality of a human-contributed dataset?
+Existing black-box methods for detecting LLM-generated content primarily focus on distinguishing human-written text from AI-generated text by analyzing statistical differences in writing style (gehrmann2019gltr; mao2024raidar; koike2024outfox; hu2023radar).
+While effective in open-ended text generation tasks, these training-based approaches are not well-suited for structured annotation tasks like multiple-choice labeling, where the discrete responses lack the rich textual features needed for classification.
+In this paper, we develop a theoretically robust scoring metric to evaluate the quality of human contributions in addition to existing LLM answers, effectively distinguishing workers with less informative responses.
+Our method builds on peer prediction, a scoring mechanism designed to incentivize truthful reporting in the absence of ground truth miller2005eliciting; shnayder2016informed.
 Traditional peer prediction methods evaluate a worker based on the correlation between her responses and those of her peers.
-Under the standard model, every worker independently observes only high-effort signals — denoted as $X_{i}$ for worker $i$ — after working on a set of i.i.d.
+In the standard model, every worker independently observes only high-effort signals, denoted as $X_{i}$ for worker $i$, after working on a set of i.i.d.
 questions.
@@
 In our settings, LLM responses serve as such cheap signals.
-To address this limitation, we extend peer prediction by measuring correlations between workers’ responses while conditioning on a cheap signal $Z$ that is observable to the requester.
+To address this limitation, we extend peer prediction by measuring correlations between workers’ responses while conditioning on a signal $Z$ that is observable to the requester.
 For example, $Z_{i}$ and $Z$ could represent responses to the questions independently generated by the same LLM or different LLMs.
-Intuitively, by conditioning on $Z$, our method scores the information within workers’ responses beyond what can be obtained from LLMs alone, ensuring that human effort is properly rewarded while LLM-assisted cheating is penalized.
-Theoretically, we identify the conditions under which our mechanism maintains information monotonicity under two regimes: (1) for any form of cheating strategy and (2) for a subset of practical cheating strategies, which we refer to as naive strategies.
-Under a naive strategy, rather than reporting honestly on all assigned tasks, the agent sometimes reports noise (random reporting) or copies the LLM output, thereby mixing high-effort signals with uninformative signals.
-To achieve information monotonicity in the general case, we require that for any two workers $i$ and $j$, $Z_{i}$ (the cheap signal of $i$) is approximately uncorrelated with both $X_{j}$ (the high-effort signal of $j$) and $Z_{j}$ (the cheap signal of $j$), once we condition on $Z$ (see Assumptions 4.1 and 4.2).
-For naive strategies, a weaker condition suffices: conditioning on $Z$, the correlation between high-effort signals $(X_{i},X_{j})$ is stronger than the correlation between either $(Z_{i},X_{j})$ or $(Z_{i},Z_{j})$ on average (see 4.5).
+Intuitively, by conditioning on $Z$, our method scores the information within workers’ responses beyond what can be obtained from LLMs alone, ensuring that human effort is properly rewarded while AI-reliant is penalized.
+Theoretically, we identify the conditions under which our mechanism maintains information monotonicity under two regimes: (1) for any form of manipulation strategy and (2) for a subset of practical cheating strategies, which we refer to as lazy-reporting strategies.
+A lazy-reporting strategy, for instance, includes blindly relying on LLM on a fraction of the tasks while answering the rest truthfully.
+To achieve information monotonicity in the general case, for an arbitrary pair of workers, $i$’s cheap signal $Z_{i}$ has to be approximately uncorrelated with $j$’s high-effort signal $X_{j}$ and their cheap signal $Z_{j}$, conditioned on $Z$ (Assumptions 4.1 and 4.2).
+For lazy-reporting strategies, a weaker condition suffices: conditioning on $Z$, the correlation between high-effort signals $(X_{i},X_{j})$ is stronger than the correlation between both $(Z_{i},X_{j})$ and $(Z_{i},Z_{j})$ (4.5).
 Empirically, we test our assumptions using two subjective labeling datasets and five types of commonly used LLMs.
@@
 However, the correlations between independent samples of LLMs responses are usually non-zero, suggesting that Assumptions 4.1 and 4.2 are too strong to hold in practice.
-This implies that it is feasible to use peer prediction to detect cheating with naive strategies, but detecting arbitrary cheating — where agents choose to cheat even after seeing the high-effort signal — may require substantial ground truth verifications.
-Finally, under scenarios where 4.5 holds, we simulate cheating agents and evaluate the effectiveness of our method in identifying them.
-Echoing our theoretical insights, we show that our approach has the most robust performance on mixed crowds compared with all the baselines.
+Finally, under scenarios where 4.5 holds, we evaluate the effectiveness of our method in detecting low-effort agents.
+Echoing our theoretical insights, we show that our approach has the most robust performance on mixed crowd compared with all the baselines.
 2 Related Work
-Our study is relevant to the usage of LLMs on crowdsourcing platforms.
-On one hand, if used properly, LLMs can assist better human-AI collaborations or even replace human labor on particular crowdsourcing tasks.
-For example, Cegin et al.
-(2023) find that ChatGPT can, to some extent, replace crowdsourcing on paraphrase generation for intent classification tasks.
-Kaikaus et al.
-(2023) focus on emotion, sentiment, and cognitive dissonance annotations, where they found that GPT labels are more consistent and reliable than human annotators.
-Ashktorab et al.
-(2021) demonstrate that “batch labeling”, a framework of using AI to assist human labeling by allowing a single labeling action to apply to multiple records, can increase the overall labeling accuracy and increase the speed.
-On the other hand, however, many studies have warned of the dangers of improperly using LLMs in various crowdsourcing tasks.
-del Rio-Chanona et al.
-(2023) show that over recent years, LLMs have decreased the activities on Stack Overflow, the leading online Q&A platform for computer programming, by 16%.
-Regarding the quality of LLM-generated content, evidences exist showing that LLMs tend to have higher bias than human annotators on qualitative analysis (Ashwin et al., 2023), and they tend to exhibit a high degree of unreliability in questions answered in the context of undergraduate computer science.
-However, LLMs evolve rapidly — a considerable amount of recent work has focused on getting LLMs to emulate human expertise via extensive crowdsourced data (Bach et al., 2022; Zhang et al., 2025; Chiang and Lee, 2023; Köpf et al., 2023).
-We anticipate that the gap between LLMs and human performance will continue to shrink in the near future.
-What’s more, the usage of LLMs in crowdsourcing has uncontrollably increased.
-In a study, it is reported that as high as 40% of MTurk workers used LLMs to cheat in text summarization tasks (Veselovsky et al., 2023).
-Furthermore, using a pre-trained LLM, Liang et al.
-(2024) reported that the use of LLMs in paper writing has significantly increased in recent years — the largest increase in CS papers was as high as 17%.
-Even worse, an empirical study suggests that humans tend to over-reliant on AI suggestions, even when they are incorrect (Campero et al., 2022).
-Taking into account both the opportunities and challenges, our work aims to address LLM cheating and foster a healthy and reliable crowdsourcing environment.
+Our study relates to the studies on LLMs usage in crowdsourcing.
+cegin2023chatgpt find that ChatGPT can partially replace human workers in paraphrase generation, and 10386425 report that GPT labels are more consistent than humans in sentiment annotations.
+ashktorab2021ai demonstrate that “batch labeling”, a framework of using AI to assist human labeling by allowing a single labeling action to apply to multiple records, can increase the overall labeling accuracy and increase the speed.
+However, several studies also highlight potential downsides: del2023large report a 16% decline in Stack Overflow activity following the rise of LLMs, and ashwin2023using find that LLMs exhibit greater bias than human annotators in qualitative analysis.
+In this context, our paper introduces a scoring mechanism that quantitatively measures the informativeness of human reports relative to the information already provided by AI.
 Our work builds on advancements in information elicitation, particularly peer prediction.
-First introduced by Miller et al.
-(2005), peer prediction mechanisms aim to incentivize truthful reporting without requiring ground truth verification.
-The core idea is to reward agents based on how well their responses predict those of others.
-The original mechanism relied on the principal’s knowledge of the correlation between agents’ signals (the information they observe).
-Later, Dasgupta and Ghosh (2013) extended this idea, removing the need for such knowledge by assuming agents answer multiple i.i.d.
-binary questions.
-Subsequent works have expanded this framework to non-binary (but finite) signals (Shnayder et al., 2016; Kong and Schoenebeck, 2019b), infinite signals (Schoenebeck and Yu, 2020), truthful guarantees with fewer questions (Kong, 2020b), heterogeneous agents (Agarwal et al., 2017), adversarial agents (Schoenebeck et al., 2021), and broader cheating strategies (Zhang and Schoenebeck, 2023b).
-As a generalization of the correlated agreement (CA) mechanism (Shnayder et al., 2016), our method can score the quality of human responses conditioned on an external information source.
-We chose the CA mechanism for its simplicity and robust empirical performance, as alternative mechanisms often have complex designs or exhibit high output variance (Burrell and Schoenebeck, 2021; Zhang and Schoenebeck, 2023a; Xu et al., 2024).
-Our approach is also inspired by Kong and Schoenebeck (2018), who proposed using conditioned mutual information to elicit hierarchical information.
-Unlike their framework, which assumes an information hierarchy where more effort yields strictly more informative signals, our method does not require hierarchical structures or sequential elicitation.
-Instead, we assume the principal has access to samples of low-effort information and focus on empirically verifying the conditions under which our method works (or not).
-Lastly, we propose a heuristic generalization of our method to handle textual and other forms of complex data (Appendix C).
-This approach is related to the discussion of applying peer prediction to elicit textual information using LLMs (Lu et al., 2024; Wu and Hartline, 2024).
-The studies on LLM-content detection are also related to this work.
-Common approaches for LLM content detection include watermarking (Gu et al., 2022; Kirchenbauer et al., 2023; Lee et al., 2023; Zhao et al., 2023; Yang et al., 2023a), which embeds identifiable markers into data, models, or generated text; statistical detection, which leverage differences in output by analyzing logits statistics (Su et al., 2023; Vasilatos et al., 2023) or vocabulary statistics (Gehrmann et al., 2019; Mao et al., 2024; Yang et al., 2023b); and deep learning-based methods, which train models to distinguish AI-generated content from human-written text (Koike et al., 2024; Hu et al., 2023; Shah et al., 2023).
-For a more comprehensive review, we refer to the survey paper (Wu et al., 2024).
+As a generalization of the correlated agreement (CA) mechanism (shnayder2016informed), our method can score the quality of human responses conditioned on an external information source.
+We chose the CA mechanism for its simplicity and robust empirical performance (burrell2021measurement; zhang2022high).
+Our approach is also inspired by kong2018eliciting, who proposed using conditioned mutual information to elicit hierarchical information.
+Their framework assumes an information hierarchy where more effort yields strictly more informative signals, so low-effort signals can be elicited first, which are used to further elicit high-effort signals.
+In contrast, our method assumes that the principal has access to samples of low-effort signals, which is more straightforward and practical in the LLM-influenced crowdsourcing settings.
+Lastly, we propose a heuristic generalization of our method to handle textual data (Appendix F), which is related to methods that elicit textual information using LLMs (textualpp2024; wu2024elicitationgpt).
+The studies on LLM-content detection are also related.
+Common approaches for LLM content detection include watermarking (gu2022watermarking; lee2023wrote; zhao2023provable; yang2023watermarking), which embeds identifiable markers into data, models, or generated text; statistical detection, which leverage differences in output by analyzing logits statistics (su2023detectllm; vasilatos2023howkgpt) or vocabulary statistics (gehrmann2019gltr; mao2024raidar); and deep learning-based methods, which train models to distinguish AI-generated content from human-written text (koike2024outfox; hu2023radar; shah2023detecting).
 However, these methods typically rely on abundant textual data for training and detection.
-In contrast, our approach is training-free and addresses cheating in scenarios where the critical distinction between human-written and AI-generated content lies in the information level rather than the vocabulary level.
-For instance, our method is designed for tasks like answering multiple-choice questions, where no direct textual response is available.
+In contrast, our approach is training-free and addresses LLM-contamination in non-textual responses scenarios, such as answering multi-choice questions.
 3 Model
@@
 questions.
-The principal distributes these questions to $n$ agents such that each agent answers multiple questions and each question is answered by more than one agent.
+The principal (requester) distributes these questions to $n$ agents (workers) such that each agent answers multiple questions and each question is answered by more than one agent.
 We first present the information structure when two agents, denoted as $i$ and $j$, answer an arbitrary question.
@@
 For example, if the answer to the question is either “yes” or “no”, then $|\Sigma|=2$.
-Suppose agent $i$ and $j$ both answer this question.
-After careful thinking, they each observe a signal $X_{i}$ and $X_{j}$ respectively.
+Suppose agents $i$ and $j$ both answer this question.
+After exerting effort, they each observe a signal $X_{i}$ and $X_{j}$ respectively.
 For simplicity, we assume the signal space is the same as the ground truth space, i.e.
 $X_{i},X_{j}\in\Sigma$.
-We can interpret $X_{i}$ and $X_{j}$ as the high-effort signals, which are agents’ best guess of $Y$ after exerting a substantial amount of effort.
-In addition to their high-effort signals, agents can obtain cheap signals $Z_{i},Z_{j}\in\Sigma$ that may or may not depend on $Y$.
-The cheap signals are usually strongly correlated, sometimes even more correlated than agents’ high-effort signals.
+We can interpret $X_{i}$ and $X_{j}$ as the full-effort signals that the agents can obtain which are their best guesses of $Y$.
+In addition to full-effort signals, agents can obtain cheap signals $Z_{i},Z_{j}\in\Sigma$ that may or may not depend on $Y$.
+Cheap signals are usually mutually correlated with each other, sometimes even more correlated than agents’ full-effort signals.
 For example, these can be the responses of two (potentially different) LLMs.
-We present the causal relationship of these variables in Fig.
-1.
-Note that $X_{i}$ and $Z_{i}$ are independent conditioned on $Y$, meaning that the high-effort signal is only correlated with the cheap signal via the ground truth.
+We present the causal relationship among these variables in Figure 1.
+Note that we assume $X_{i}$ and $Z_{i}$ are independent conditioned on $Y$, meaning that the full-effort signal is only correlated with the cheap signal via the ground truth.
 However, $Z_{i}$ and $Z_{j}$ may have additional correlations conditioned on $Y$.
-This may happen when agents cheat with LLMs trained on similar data.
-Strategies  The principal’s goal is to collect high-effort signals from agents.
-This means that we consider scenarios where human expertise remains the gold standard, or AI underperforms human agents.
+This may happen when agents rely on LLMs trained on similar data.
+Strategies  The principal’s goal is to recover the ground truth as accurately as possible.
+We consider scenarios where AI underperforms human agents, motivating the principal to seek full-effort information from human agents.
 Such situations arise when the notion of ground truth is inherently subjective, shaped by human preferences—such as when labelers are asked to rank LLM responses according to personal preference, a common practice in reinforcement learning with human feedback.
@@
 Let $\tau$ be the truth-telling strategy, i.e.
-$\tau(X_{i},Z_{i})=X_{i}$, and let $\mu_{i}$ denote an uninformative strategy such that $\mu_{i}(X_{i},Z_{i})$ is independent of $X_{i}$.
-The definitions for $\hat{X}_{j}$ and $\theta_{j}$ are analogous.
-We highlight a special type of strategy, called the naive strategy, which is more natural and common on crowdsourcing platforms.
+$\tau(X_{i},Z_{i})=X_{i}$, and let $\mu_{i}$ denote a no-effort strategy such that $\mu_{i}(X_{i},Z_{i})$ is independent of $X_{i}$ conditioned on $Z_{i}$.
+The argument for $\hat{X}_{j}$ and $\theta_{j}$ are analogous.
+We highlight a special type of strategy, called the lazy-reporting strategy, which is more practical and common on crowdsourcing platforms.
 Definition 3.1.
-A naive strategy $\nu\in\Theta_{\nu}$ is a convex combination between truth-telling and an uninformative strategy $\mu$, i.e.
+A lazy-reporting strategy $\nu\in\Theta_{\nu}$ is a convex combination between truth-telling and a no-effort strategy $\mu$, i.e.
 $\nu(X_{i},Z_{i})=\begin{cases}X_{i}&\text{with probability }p_{i}<1,\\ \mu(X_{i},Z_{i})&\text{otherwise.}\end{cases}$
-Examples of naive strategies include: exerting effort and reporting $X_{i}$ for some questions while randomly selecting answers for the rest; or fully relying on an LLM and always reporting $Z_{i}$.
-In contrast, non-naive strategies are typically complex and impractical, often resembling malicious behavior rather than mere low-effort cheating.
+In particular, $\nu$ does not make $\hat{X}_{i}$ depend jointly on both $X_{i}$ and $Z_{i}$ on the same question.
+Examples of lazy-reporting strategies include: exerting effort and reporting $X_{i}$ for some questions while randomly selecting answers for the rest; or fully relying on an LLM and always reporting $Z_{i}$.
+In contrast, non-lazy reporting strategies are typically complex and impractical, often resembling malicious rather than mere low-effort behavior.
 For example, an agent may first exert effort to obtain $X_{i}$ for each question and then mix it with $Z_{i}$ by reporting “yes” if either $X_{i}$ or $Z_{i}$ is “yes”, and “no” otherwise.
 Principal Information  The principal can observe the reports $\hat{X}_{i}$ and $\hat{X}_{j}$, but not the underlying signals $X_{i},X_{j}$ and $Z_{i},Z_{j}$.
-Furthermore, the principal cannot directly evaluate human responses using the ground truth $Y$.
+This implies that the principal cannot directly evaluate human responses using the ground truth $Y$.
 Without knowing $Y$, the principal cannot distinguish the real information $X_{i}$ and the cheap signal $Z_{i}$ without further assumptions.
 Our method assumes that the principal can observe a signal $Z$ that is (strongly) correlated with the agents’ cheap signals.
-In the context of LLM cheating, for example, the principal can generate answers to a fraction of the questions using a popular LLM.
-The effectiveness of our method necessarily depends on the quality of $Z$ — how well the principal’s information can block the correlation between agents’ cheap signals.
+In the context of LLM contamination, for example, the principal can generate answers to a fraction of the questions using a popular LLM.
+The effectiveness of our method necessarily depends on the quality of ZZ—how well the principal’s information can block the correlation between agents’ cheap signals.
 We will introduce our method and the assumptions in Section 4 and empirically verify the assumptions in Section 5.2.
 3.1 Objectives and Problem Statement
-In practice, the principal may collect responses from multiple agents, resulting in an $n\times m$ response matrix $\hat{X}$ where each entry $\hat{X}_{i,q}\in\Sigma$ if agent $i$ answers question $q$ and $\hat{X}_{i,q}=\phi$ otherwise.
+In practice, the principal may collect responses from multiple agents, resulting in an $n\times m$ response matrix $\hat{X}$ where each entry $\hat{X}_{i,k}\in\Sigma$ if agent $i$ answers question $k$ and $\hat{X}_{i,k}=\emptyset$ otherwise.
 Given $\hat{X}$ and the principal’s samples of the cheap signal $Z$, a scoring mechanism designs a score for each agent.
@@
 Definition 3.2.
-A mechanism is information monotone for naive strategies if, for any agent $i$, $S_{i}(X_{i},X_{-i}\mid Z)>S_{i}(\hat{X}_{i},\hat{X}_{-i}\mid Z)$ for any naive strategy $\nu_{i}\in\Theta_{\nu}$ with $\hat{X}_{i}=\nu_{i}(X_{i},Z_{i})$.
-A mechanism is $\epsilon$-information monotone if it is information monotone for naive strategies and, additionally, $S_{i}(X_{i},X_{-i}\mid Z)\geq S_{i}(\hat{X}_{i},\hat{X}_{-i}\mid Z)-\epsilon$ holds for any strategy $\theta_{i}\in\Theta$ where $\hat{X}_{i}=\theta_{i}(X_{i},Z_{i})$.
-At the individual level, information monotone scores can be used to identify unreliable labelers or LLM cheaters.
+A mechanism is information monotone for lazy-reporting strategies if, for any agent $i$, $S_{i}(X_{i},X_{-i}\mid Z)>S_{i}(\hat{X}_{i},\hat{X}_{-i}\mid Z)$ for any lazy-reporting strategy $\nu_{i}\in\Theta_{\nu}$ with $\hat{X}_{i}=\nu_{i}(X_{i},Z_{i})$.
+A mechanism is $\epsilon$-information monotone if it is information monotone for lazy-reporting strategies and, additionally, $S_{i}(X_{i},X_{-i}\mid Z)\geq S_{i}(\hat{X}_{i},\hat{X}_{-i}\mid Z)-\epsilon$ holds for any strategy $\theta_{i}\in\Theta$ with $\hat{X}_{i}=\theta_{i}(X_{i},Z_{i})$.
+At the individual level, information monotone scores can be used to identify low-effort agents.
 At the dataset level, the score distribution across agents can help interpret the quality of the crowdsourced dataset.
-We note that the same concept is termed informed truthfulness or informed strategy-proofness in the peer prediction literature, where the objective is to incentivize agents to report truthful information (Shnayder et al., 2016).
+We note that the same concept is termed informed truthfulness or informed strategy-proofness in the peer prediction literature, where the objective is to incentivize agents to report truthful information shnayder2016informed.
 4 Our Method
-Classic peer prediction methods often struggle with cheap signals that are highly correlated with one another but only weakly correlated with the ground truth.
-In this section, we generalize a prior work — the correlated agreement (CA) mechanism (introduced in Appendix A) — and propose a mechanism that achieves information monotonicity when the principal can obtain some noisy samples of the cheap signal.
-In the context of LLM cheating, these samples can be LLM-generated labels for the crowdsourced questions.
-.
-We name our generalization as the conditioned CA mechanism, which scores an agent based on the correlation between her responses and her peers’ conditioned on the principal’s samples of $Z$.
+Classic peer prediction methods often struggle with cheap signals, which are highly correlated with one another but only weakly correlated with the ground truth.
+In this section, we generalize a prior work—the correlated agreement (CA) mechanism (introduced in Appendix B)—and propose a mechanism that achieves information monotonicity when the principal can obtain some noisy samples of the cheap signal.
+In the context of LLM contamination, these samples can be the labels generated by a popular LLM on a fraction of the crowdsourced questions.
+We call our generalization the conditioned CA mechanism, which scores an agent based on the correlation between her responses and her peers’ conditioned on the principal’s samples of $Z$.
 Our method has two main steps.
@@
 A scoring function determines whether a pair of signals “agree” or not, which is learned based on data.
-We first estimate the empirical conditioned joint distribution $\tilde{P}(\hat{X}_{i},\hat{X}_{j}\mid Z)$, computed using the frequency of observing a pair of signals on the same question conditioned on $Z$.***We can use all agents’ reports to estimate the same distribution because we assume agents are homogeneous and questions are i.i.d..
+We first estimate the empirical conditioned joint distribution $\tilde{P}(\hat{X}_{i},\hat{X}_{j}\mid Z)$, computed using the frequency of observing a pair of signals on the same question conditioned on ZZ.
+[^1] We can use all agents’ reports to estimate the same distribution because we assume agents are homogeneous and questions are i.i.d..
 This assumption is primarily made for practical concerns, as there is usually a shortage of data to estimate the distribution accurately for each pair of agents.
@@
 $\tilde{\Delta}_{h,l,k}=\tilde{P}\left(\hat{X}_{i}=h,\hat{X}_{j}=l\mid Z=k\right)-\tilde{P}\left(\hat{X}_{i}=h\mid Z=k\right)\tilde{P}\left(\hat{X}_{j}=l\mid Z=k\right),$
-which is the difference between the joint distribution and the product of marginal distributions of the reports, conditioned on $Z$.
-This gives us the scoring function $T_{\tilde{\Delta}}=\text{Sign}(\tilde{\Delta})$, i.e.
-$T_{\tilde{\Delta}}(h,l,k)=1$ if $\tilde{\Delta}_{h,l,k}>0$ and 00 otherwise.
-This step becomes unnecessary if the true scoring function $T_{\Delta}=\text{Sign}(\Delta)$ is already known, where $\Delta$ is defined analogously to $\tilde{\Delta}$, but based on the true signals $X_{i}$ and $X_{j}$.
-For example, for binary questions with positively correlated signals, $\Delta_{\cdot,\cdot,k}$ is an identity matrix for any $k\in\Sigma$.
+[…]
 Compute the bonus/penalty score.
-We then repeatedly sample a bonus question $q$ and a penalty question $q^{\prime}$ while ensuring $Z_{q}=Z_{q^{\prime}}=k$.
+Fixing a $k\in\Sigma$, we then repeatedly sample a bonus question $q$ and a penalty question $q^{\prime}$ while ensuring $Z_{q}=Z_{q^{\prime}}=k$.
 For each pair of samples, the score for agent $i$ is:
 $T_{\tilde{\Delta}}(\hat{X}_{i,q},\hat{X}_{j,q},k)-T_{\tilde{\Delta}}(\hat{X}_{i,q},\hat{X}_{j,q^{\prime}},k),$
-where $j$ is a randomly selected agent who answers question $q^{\prime}$.
-The final score for agent $i$ is obtained by averaging over all sampled pairs of $q$ and $q^{\prime}$.
-Algorithm 1 presents the procedures of our method.
-Input: The crowdsourced dataset $\hat{X}$, the samples of cheap signal $Z$, the (estimated) scoring function $T_{\tilde{\Delta}}$
-Estimate the empirical marginal distribution of $Z$, i.e. $\tilde{P}(Z=\sigma)$ for $\sigma\in\Sigma$.
-for agent $i\in[n]$ do
-       Let $M_{i}$ be the set of questions answered by agent $i$ Initialize $s_{i}=0$;
-       for $k\in\Sigma$ do
-             Initialize $x=0$;
-             for question $q\in M_{i}\cap M_{Z}(k)$ do
-                   Randomly select a peer $j$ who answers $q$;
-                   $x\mathrel{+}=T_{\tilde{\Delta}}(\hat{X}_{i,q},\hat{X}_{j,q},k)$;
-                   Randomly select a question $q^{\prime}\in M_{j}\cap M_{Z}(k)$;
-                   $x\mathrel{-}=T_{\tilde{\Delta}}(\hat{X}_{i,q},\hat{X}_{j,q^{\prime}},k)$
-            $s_{i}\mathrel{+}=x\cdot\tilde{P}(Z=k)/|M_{i}\cap M_{Z}(k)|$
-return $s$
-Algorithm 1 The conditioned CA mechanism
+where $j$ is a randomly selected agent who answers both $q$ and $q^{\prime}$.
+The final score for agent $i$ is obtained by averaging over all $k\in\Sigma$ and the sampled pairs of $q$ and $q^{\prime}$.
 Intuitively, the conditioned CA mechanism encourages agents to agree on the same question and penalizes agents when they agree on two distinct questions, given that the principal’s sampled signal is identical on both questions.
@@
 This results in a score that is close to zero.
+We defer more details to Appendix C.
 4.1 Theoretical Insights
@@
 Before we start, we first present an intuitive way to interpret the score computed by our mechanism.
-In a prior work, Kong and Schoenebeck (2019a) show that the score of the CA mechanism can be interpreted as a type of mutual information between agents’ reports called the total variation distance (TVD) mutual information.
-Inspired by this prior work, we show that when agents report the high-effort signal and the underlying scoring function $T_{\Delta}$ is applied, the expected score computed by the conditioned CA mechanism is essentially the TVD conditioned mutual information:
+In a prior work, FrameworkKong2019 show that the score of the CA mechanism can be interpreted as a type of mutual information between agents’ reports called the total variation distance (TVD) mutual information.
+Inspired by this prior work, we show that when agents report the high-effort signal and the underlying scoring function $T_{\Delta}$ is applied, the expected score computed by the conditioned CA mechanism is essentially the conditional TVD mutual information:
 $I_{\text{TVD}}(X_{i};X_{j}\mid Z)=\sum_{z\in\Sigma}\Pr(Z=z)\sum_{\sigma,\sigma^{\prime}\in\Sigma}\left|\Pr(X_{i}=\sigma,X_{j}=\sigma^{\prime}\mid z)-\Pr(X_{i}=\sigma\mid z)\Pr(X_{j}=\sigma^{\prime}\mid z)\right|.$
+The proof of this observation can be found in Section D.1.
 4.1.1 Conditions for $\epsilon$-Information Monotonicity
@@
 Assumption 4.1.
-We assume $Z_{i}$ and $Y$ are approximately independent conditioned on $Z$, i.e. $|\Pr(Z_{i},Y\mid Z)-\Pr(Z_{i}\mid Z)\Pr(Y\mid Z)|\leq\epsilon_{i}$ for any $i\in[n]$.
+We assume $Z_{i}$ and $Y$ are approximately independent conditioned on $Z$, i.e. $\left|\Pr(Z_{i},Y\mid Z)-\Pr(Z_{i}\mid Z)\Pr(Y\mid Z)\right|\leq\epsilon_{i}$ for any $i\in[n]$.
 Assumption 4.2.
@@
 If $T_{\Delta}$ is known and $I_{\text{TVD}}(X_{i};X_{j}\mid Z)>\hat{\epsilon}\coloneqq\left((\epsilon+\epsilon_{i}+\epsilon_{j})|\Sigma|^{2}+(\epsilon_{i}+\epsilon_{j})|\Sigma|^{3}\right)$, then under 4.1 and 4.2, the conditioned CA mechanism is $\hat{\epsilon}-$information monotone.
-We defer the proof to Section B.1.
+We defer the proof to Section D.1.
 When agents report the high-effort signals truthfully, the expected score is the TVD mutual information between $X_{i}$ and $X_{j}$.
-Theorem 4.3 suggests that if $Z$ approximately blocks the correlation between $Z_{i}$ with $X_{j}$ and $Z_{j}$, then any cheating strategy can only decrease the TVD mutual information between the reports $\hat{X}_{i}$ and $\hat{X}_{j}$ up to an error.
-In the context of LLM cheating, suppose the principal suspects that some agents are using LLMs to generate their labels.
-Our mechanism suggests that the principal can generate labels using a (potentially different) LLM and compute the CA score conditioned on these generated labels.
-According to Theorem 4.3, if the principal’s generated labels satisfy two conditions — 1) they block correlations among the labels generated by potential cheating agents (4.2), and 2) they block correlations between those labels and the ground truth (4.1) — then the conditioned CA mechanism’s score is information monotone.
+Theorem 4.3 suggests that if $Z$ approximately blocks the correlation between $Z_{i}$ with $X_{j}$ and $Z_{j}$, then any manipulation can only decrease the TVD mutual information between $\hat{X}_{i}$ and $\hat{X}_{j}$ up to an error.
 We further show the necessity of the assumptions for information monotonicity.
-We construct an example to show that if $Z_{i}$ and $Z_{j}$ have some correlation that is not captured by $Z$, then there exists a signal structure such that the score computed by the conditioned CA mechanism is not maximized by reporting $X_{i}$ and $X_{j}$.
+The following proposition suggests that if $Z_{i}$ and $Z_{j}$ have some correlation that is not captured by $Z$, then there exists a signal structure following the relationship in Figure 1, such that the score computed by the conditioned CA mechanism is not maximized by reporting $X_{i}$ and $X_{j}$.
 Proposition 4.4.
 Suppose $T_{\Delta}$ is known.
-If there exist $(z_{i},z_{j},y,z)$ such that $|\Pr(Z_{i}=z_{i},Z_{j}=z_{j}\mid Y=y,Z=z)-\Pr(Z_{i}=z_{i}\mid Y=y,Z=z)\Pr(Z_{j}=z_{j}\mid Y=y,Z=z)|=d$ for some $0<d\leq\frac{1}{4}$, then there exists a joint distribution $\Pr(X_{i},X_{j},Z_{i},Z_{j},Y,Z)$ that satisfies the causal relationship described in Fig.
-1 and a strategy profile $(\theta_{i},\theta_{j})\neq(\tau,\tau)$ such that $S_{i}(\hat{X}_{i},\hat{X}_{j})\geq S_{i}(X_{i},X_{j})+8d^{2}$ where $\hat{X}_{i}=\theta_{i}(X_{i},Z_{i})$ and $\hat{X}_{j}=\theta_{j}(X_{j},Z_{j})$.
-Before we provide the full proof (deferred to Section B.2), we briefly introduce the high-level idea.
+If there exist $(z_{i},z_{j},y,z)$ such that $|\Pr(Z_{i}=z_{i},Z_{j}=z_{j}\mid Y=y,Z=z)-\Pr(Z_{i}=z_{i}\mid Y=y,Z=z)\Pr(Z_{j}=z_{j}\mid Y=y,Z=z)|=d$ for some $0<d\leq\frac{1}{4}$, then there exists a joint distribution $\Pr(X_{i},X_{j},Z_{i},Z_{j},Y,Z)$ that satisfies the causal relationship described in Figure 1 and a strategy profile $(\theta_{i},\theta_{j})\neq(\tau,\tau)$ such that $S_{i}(\hat{X}_{i},\hat{X}_{j}\mid Z)\geq S_{i}(X_{i},X_{j}\mid Z)+8d^{2}$ where $\hat{X}_{i}=\theta_{i}(X_{i},Z_{i})$ and $\hat{X}_{j}=\theta_{j}(X_{j},Z_{j})$.
+Before we provide the full proof (deferred to Section D.2), we briefly introduce the high-level idea.
 Consider a signal structure with a binary signal space $\Sigma=\{0,1\}$.
@@
 Note that although this example only indicates that 4.2 is necessary, the same recipe can be straightforwardly applied to show the necessity of 4.1.
-4.1.2 Conditions for Information Monotonicity Under Naive Strategies
+4.1.2 Conditions for Information Monotonicity Under Lazy-reporting Strategies
 We know 4.1 and 4.2 are necessary for the conditioned CA mechanism to be $\epsilon$-information monotone.
-However, these assumptions are designed to guard against complex, unrealistic strategies, such as those where cheaters first obtain the high-effort signal and then blend it with the cheap signal.
-If we instead restrict our focus to naive strategies, these assumptions on $Z$ can be significantly relaxed.
+However, these assumptions are designed to guard against complex, unrealistic strategies, such as those where agents first obtain the full-effort signal and then blend it with the cheap signal.
+If we instead restrict our focus to lazy-reporting strategies, these assumptions on $Z$ can be significantly relaxed.
 Assumption 4.5.
-For any pair of agents $i$ and $j$, we assume the TVD conditioned mutual information between $X_{i}$, $X_{j}$, $Z_{i}$, $Z_{j}$ satisfies that
+For any pair of agents $i$ and $j$, we assume the conditional TVD mutual information between $X_{i}$, $X_{j}$, $Z_{i}$, $Z_{j}$ satisfies that
 $I_{\text{TVD}}(X_{i};X_{j}\mid Z)>\max\left(I_{\text{TVD}}(Z_{i};Z_{j}\mid Z),I_{\text{TVD}}(Z_{i};X_{j}\mid Z)\right).$
 Proposition 4.6.
-If $T_{\Delta}$ is known and 4.5 holds, the conditioned CA mechanism is information monotone for naive strategies.
-We defer the proof to Section B.3.
-Intuitively, a naive strategy is essentially a mixture of truth-telling and uninformative strategies.
-We show that when both agents follow naive strategies, the score-maximizing equilibrium must fall into one of the following three cases: 1) both agents exert full effort, achieving an expected score of $I_{\text{TVD}}(X_{i};X_{j}\mid Z)$; 2) both agents report cheap signals, with expected score $I_{\text{TVD}}(Z_{i};Z_{j}\mid Z)$; 3) one agent exerts full effort while the other reports a cheap signal, yielding expected score $I_{\text{TVD}}(Z_{i};X_{j}\mid Z)$.
-Therefore, the inequality in 4.5 ensures that full effort strictly dominates naive strategies, establishing the proposition.
+If $T_{\Delta}$ is known and 4.5 holds, the conditioned CA mechanism is information monotone for lazy-reporting strategies.
+We defer the proof to Section D.3.
+Intuitively, a lazy-reporting strategy is essentially a mixture of truth-telling and no-effort strategies.
+We show that when both agents follow lazy-reporting strategies, the score-maximizing equilibrium must fall into one of the following three cases: 1) both agents exert full effort, achieving an expected score of $I_{\text{TVD}}(X_{i};X_{j}\mid Z)$; 2) both agents report cheap signals, with expected score $I_{\text{TVD}}(Z_{i};Z_{j}\mid Z)$; 3) one agent exerts full effort while the other reports a cheap signal, yielding expected score $I_{\text{TVD}}(Z_{i};X_{j}\mid Z)$.
+Therefore, the inequality in 4.5 ensures that full effort strictly dominates lazy-reporting strategies, establishing the proposition.
 Note that 4.5 is significantly weaker than 4.1 and 4.2, which require both $I_{\text{TVD}}(Z_{i};Z_{j}\mid Z)$ and $I_{\text{TVD}}(Z_{i};X_{j}\mid Z)$ to be close to zero.
-4.1.3 Learning the scoring function $T_{\Delta}$.
+Remark 4.7.
 Thus far, we have assumed that the true scoring function $T_{\Delta}$ is known.
-For complex signal structures, especially when the signal space is large, we have to learn the scoring function from the data.
-This introduces a concern: if a significant fraction of agents are cheating, the learned scoring function $T_{\tilde{\Delta}}$ might be manipulated in a way that benefits the cheating reports.
-We mitigate this concern using the following proposition.
-Let $S_{i}^{T}(\theta_{i},\theta_{j})$ be the expected score computed under the scoring function $T$ when agents’ strategies are $\theta_{i}$ and $\theta_{j}$ respectively.
-Proposition 4.7.
-Suppose 4.1 and 4.2 hold with no error such that $\epsilon=\epsilon_{i}=0$ for any $i\in[n]$.
-Then, for any scoring function $T\in\{0,1\}^{|\Sigma|\times|\Sigma|}$ and any strategy profile $(\theta_{i},\theta_{j})$, $S_{i}^{T}(\theta_{i},\theta_{j})\leq S_{i}^{T_{\Delta}}(\tau,\tau)$, where $\tau$ denote the truth-telling strategy.
-We defer the proof to Section B.4.
-Intuitively, the proposition holds because when $T_{\Delta}$ is known and agents are truth-telling, the expected score is $I_{\text{TVD}}(X_{i};X_{j}\mid Z)$.
-However, if $T_{\Delta}$ is mislearned, the expected score is upper bounded by $I_{\text{TVD}}(\hat{X}_{i};\hat{X}_{j}\mid Z)$.
-Since our assumptions ensure that $Z_{i}$ and $Z_{j}$ are independent given $Z$, the data processing inequality guarantees that $I_{\text{TVD}}(\hat{X}_{i};\hat{X}_{j}\mid Z)<I_{\text{TVD}}(X_{i};X_{j}\mid Z)$.
-4.7 implies that the expected score of the conditioned CA mechanism is maximized if every agent exerts full effort and reports $X_{i}$, while any strategy that manipulates the true scoring function will only (weakly) decrease the expected score.***Here we assume that the principal can learn the correct $T_{\Delta}$ if everyone reports their high-effort signal.
+In Appendix E, we show that the theoretical guarantees still hold even when we have to learn the scoring function from the noisy and potentially corrupted data.
 5 Experiments
@@
 Next, we simulate a mixed, noisy crowd of labelers and assess whether our method can detect unreliable labelers better than several well-established baselines.
+The datasets and code of our experiments are available at https://github.com/yichiz97/LLM_contamination.
 5.1 Datasets
-Hatefulness/Offensiveness/Toxicity Labeling  We first consider a toxicity labeling dataset which includes $3459$ social media user comments posted in response to political news posts and videos on Twitter, YouTube, and Reddit in August 2021 (Schöpke-Gonzalez et al., 2025).
+Hatefulness/Offensiveness/Toxicity Labeling  We first consider a toxicity labeling dataset which includes $3459$ social media user comments posted in response to political news posts and videos on Twitter, YouTube, and Reddit in August 2021 (schopke2025using).
 MTurk workers are assigned these comments and are requested to annotate them as hateful, offensive, and/or toxic.
 The definitions of hatefulness, offensiveness, and toxicity are selected from highly cited previous literature.
-In this dataset, every question is answered by $5$ workers, and each worker answers $73$ questions on average.
-Preference Alignment
-We further consider an LLM alignment dataset where each question is a comparison between two answers to the same prompt generated by two different LLMs.
+Every question is answered by $5$ workers and each worker answers $73$ questions on average.
+Preference Alignment  We further use an alignment dataset where each question compares two LLM-generated answers to the same prompt.
 Labelers rate their preference on a 0–4 scale (0: A clearly better, 4: B clearly better, 2: tie).
-The comparison is scored from several more detailed angles, such as the helpfulness and the truthfulness of the answer.
-Nonetheless, we only focus on the overall preference.
-In total, there are 10 4611046110\,46110 461 questions where each question is answered by two normal labelers and two experts.
-On average, each labeler answers around $180$ questions.
-All labelers are recruited from Prolific.
-For more details, refer to Miranda et al.
-(2024).
-5.2 Large Language Model Usage
-In this work, we select five well-known LLMs to generate labels or annotations for all used datasets, including GPT-3.5-turbo, GPT-4, Gemma-2-2b-it***https://huggingface.co/google/gemma-2-2b-it, Phi-3.5-mini-Instruct ***https://huggingface.co/microsoft/Phi-3.5-mini-instruct, and Mistral-7B-Instruct-v0.3***https://huggingface.co/mistralai/Mistral-7B-Instruct-v0.3.
-The corresponding prompt templates are provided in Appendix E.
-Note that we choose only five models because other commonly used LLMs, such as LLaMA and Gemini, refuse to answer toxicity-related questions.
-In Appendix C, we use GPT-4 to process each human review and LLM-generated review into a $19$-dimension vector to implement our method.
+The comparison is scored from several more detailed angles, while we only focus on the overall preference.
+The dataset contains 10 46110\,461 questions, each labeled by two crowd workers and two experts.
+On average, each labeler answers about 180 questions.
+For more details, refer to miranda2024hybrid.
 5.2 Assumption Verification
@@
 The first two hypotheses correspond to 4.1 and 4.2 respectively and the other two hypotheses correspond to 4.5.
-More broadly, our results in this section provide valuable insights into the correlations between different LLMs and their correlations compared with human labelers in the context of subjective labeling questions.
-For each considered LLM, we prompt it to independently generate responses to each question in the dataset three times using the default temperature.
+More broadly, our results in this section provide valuable insights into the correlations between different LLMs and their correlations compared with human agents in the context of subjective labeling questions.
+For each considered LLM (as detailed in Section G.1), we prompt it to independently generate responses to each question in the dataset three times using the default temperature.
 These responses are treated as samples of the cheap signals $Z_{i}$, $Z_{j}$, and $Z$.
 Furthermore, we randomly select two human responses in the original dataset as samples of $X_{i}$ and $X_{j}$, respectively.
-Using these samples across all tasks, we estimate the conditioned joint distributions and then compute the conditioned mutual information.
-We report the results for $I_{\text{TVD}}(Z_{i};X_{j}\mid Z)$ on two datasets in Fig.
-2, while results for other datasets and for $I_{\text{TVD}}(Z_{i};Z_{j}\mid Z)$ are deferred to Section D.1.
-(a) Hatefulness
-(b) Preference Alignment
-Figure 2: The TVD mutual information between $Z_{i}$ and $X_{j}$ conditioned on $Z$, i.e.
-$I_{\text{TVD}}(Z_{i};X_{j}\mid Z)$.
+Using these samples across all tasks, we estimate the conditioned joint distributions and then compute the conditional mutual information.
+We report the results for $I_{\text{TVD}}(Z_{i};X_{j}\mid Z)$ in Figure 2, while results for $I_{\text{TVD}}(Z_{i};Z_{j}\mid Z)$ are deferred to Section G.3.
+(a) Preference Alignment
+(b) Hatefulness
+(c) Toxicity
+(d) Offensiveness
+Figure 2: The TVD mutual information between $X_{i}$ and $X_{j}$ (for “Human”, “Normal worker”, and “expert”) or $Z_{i}$ and $X_{j}$ (for remaining bars) conditioned on $Z$, i.e.
+$I_{\text{TVD}}(X_{i};X_{j}\mid Z)$ or $I_{\text{TVD}}(Z_{i};X_{j}\mid Z)$.
 LLM names on the x-axis denote the models used to sample $Z$ while each bar in the same group represents the model used to sample $Z_{i}$.
-We use “Random Cheaters” as a reference, who randomly selects an answer to each question according to the prior.
+We use “Random Agents” as a reference, who randomly selects an answer to each question according to the prior.
 Hypotheses 1 and 2 generally do not hold.
-In Fig.
-2, we observe that when $Z_{i}$ and $Z$ are sampled from different models, the mutual information between $Z_{i}$ and $X_{j}$ is often significantly higher than that expected from the random reporting (red bar).
+In Figure 2, we observe that when $Z_{i}$ and $Z$ are sampled from different models, the mutual information between $Z_{i}$ and $X_{j}$ is often significantly higher than that expected from the random reporting (red bar).
 Even when $Z_{i}$ and $Z$ are independently sampled by the same model, conditioning on $Z$ reduces but does not eliminate this correlation.
-This indicates that independently generated responses from the same LLM can remain dependent after conditioning on another independent sample, contradicting Hypothesis 1.
-As presented in Section D.1, this pattern continues to hold for $I_{\text{TVD}}(Z_{i};Z_{j}\mid Z)$, rejecting hypothesis 2.
+This indicates that even independently generated responses from the same LLM can remain dependent after conditioning on another independent sample, contradicting Hypothesis 1.
+As presented in Section G.3, this pattern continues to hold for $I_{\text{TVD}}(Z_{i};Z_{j}\mid Z)$, rejecting hypothesis 2.
 Hypothesis 3 and 4 generally hold when conditioning on the “right” model.
-Encouragingly, the conditioned correlation between an LLM and a human labeler is consistently weaker than that between two human labelers when $Z_{i}$ and $Z$ are sampled from the same model.
-For example, in Fig.
-2(b), the unconditioned mutual information between GPT-4-generated labels and human labels initially exceeds that of two expert labels.
-However, after conditioning on independent GPT-4 responses, $I_{\text{TVD}}(X_{i};X_{j}\mid Z)$ becomes significantly higher than $I_{\text{TVD}}(Z_{i};X_{j}\mid Z)$.
+Encouragingly, the conditioned correlation between an LLM and a human agent is consistently weaker than that between two human agents when $Z_{i}$ and $Z$ come from the same model.
+For example, in Figure 2(a), the unconditional mutual information between GPT-4-generated labels and human labels exceeds that of two expert labels.
+However, after conditioning on GPT-4 responses, $I_{\text{TVD}}(X_{i};X_{j}\mid Z)$ becomes significantly higher than $I_{\text{TVD}}(Z_{i};X_{j}\mid Z)$.
 This pattern is consistent across all tested datasets, providing strong support for hypothesis 3.
 Results for Hypothesis 4 are deferred to the appendix.
-Our empirical results suggest that as of the time of this study and based on the subjective labeling tasks we tested, human labelers exhibit distinct correlation patterns that no LLM replicates.
-5.4 Detecting Cheating Agents
-In this subsection, we focus on identifying cheating agents with naive strategies and comparing various methods based on their area under the ROC curve (AUC).
+Our empirical results suggest that as of the time of this study and based on the subjective labeling tasks we tested, human agents exhibit distinct correlation patterns that no LLM replicates.
+5.3 Detecting Low-effort Agents
+In this subsection, we focus on identifying low-effort agents with lazy-reporting strategies and comparing various methods based on their area under the ROC curve (AUC).
 5.3.1 Simulating Noisy Crowds
-Real-world crowdsourcing datasets do not include labels identifying whether an agent is cheating or working diligently.
-To enable evaluation, we simulate cheating agents and treat the original dataset labels as representing “high-effort” responses.
-Specifically, we replace a randomly selected fraction of labelers in the original dataset with cheaters adopting naive strategies.
-To reflect the noisy nature of real-world crowdsourcing data, we simulate three types of cheaters, each representing a common naive strategy observed in practice (Yang et al., 2019; Becker et al., 2021):
+Real-world crowdsourcing datasets do not include labels identifying whether an agent is exerting low effort or working diligently.
+To enable evaluation, we simulate low-effort agents and treat the original dataset labels as representing “full-effort” responses.
+Specifically, we replace a randomly selected fraction of agents in the original dataset with simulated low-effort agents adopting lazy-reporting strategies.
+To reflect the noisy nature of real-world crowdsourcing data, we simulate three types of low-effort agents, each corresponding to a common lazy-reporting strategy observed in practice yang2019modeling; becker2021crowd:
   1. 1.
-    An $\alpha_{llm}$ fraction of labelers are LLM cheaters, who report the LLM-generated labels on all assigned questions.
+    An $\alpha_{llm}$ fraction are LLM-reliant agents, who report the LLM-generated labels on all assigned questions.
   2. 2.
-    An $\alpha_{r}$ fraction of labelers are random cheaters, who generate labels by independently sampling from the marginal distribution over labels for each assigned question.
+    An $\alpha_{r}$ fraction are random agents, who generate labels by independently sampling from the marginal distribution over labels for each assigned question.
   3. 3.
-    An $\alpha_{b}$ fraction of labelers are biased cheaters, who report the dataset’s majority label on 90% of questions and choose uniformly at random on the rest.
-Figure 3: Distribution of the conditioned CA scores with $(\alpha_{llm},\alpha_{r},\alpha_{b})=(0.1,0.05,0.05)$, where GPT-4 generates samples for $Z_{i}$, $Z_{j}$, and $Z$ using the preference alignment dataset.
-Figure 3 shows the distribution of conditioned CA scores, where both the LLM cheaters’ responses and the principal’s signals are independently sampled from separate GPT-4 outputs.
-As shown, human workers achieve higher scores than any type of cheater on average.
-However, some workers fall below the average score of cheaters.
-Several factors contribute to this issue.
-For example, a subset of workers may themselves be cheaters in the dataset, or they may have been assigned many tasks that are inherently difficult to classify.
-Consequently, no detection method can achieve perfect accuracy.
-Next, we will examine how well our method distinguishes workers from cheaters compared to other baselines.
+    An $\alpha_{b}$ fraction are biased agents, who report the dataset’s majority label on 90% of questions and choose uniformly at random on the rest.
+Figure 3(a) shows the distribution of conditioned CA scores, where both the LLM-reliant agents’ responses and the principal’s signals are independently sampled from separate GPT-4 outputs.
+As shown, human workers achieve higher scores than any type of low-effort agent on average.
+However, some workers score below the average low-effort agent.
+One possible explanation is that a subset of workers in the dataset may also have exerted low effort.
+Due to these factors, no detection method can achieve perfect accuracy.
+(a) Score distribution
+(b) AUC
+Figure 3: (a) Distribution of the conditioned CA scores with $(\alpha_{llm},\alpha_{r},\alpha_{b})=(0.1,0.05,0.05)$.
+(b) AUC scores of various methods as the fraction of LLM-reliant agents increases.
+In both panels, GPT-4 generates samples for $Z_{i}$, $Z_{j}$, and $Z$ using the preference alignment dataset.
 5.3.2 Baseline Measurements
-We introduce the key ideas of the following four baselines.
+We introduce the key ideas of the following four baselines and defer the details to Section G.2.
   1. 1.
-    We consider the Output Agreement (OA) mechanism that directly counts the frequency of agreements between two agent’s responses.
-Specifically, while scoring labeler $i$, we iteratively pair her with a different labeler $j$ and let $M_{i,j}$ denote the set of questions answered by both labelers.
-Then, the score of labeler $i$ is the average of the empirical probability of agreement, i.e.
-$\frac{1}{n}\sum_{j\in[n]\backslash\{i\}}\sum_{q\in M_{i,j}}\frac{\mathbbm{1}\left[\hat{X}_{i,q}=\hat{X}_{j,q}\right]}{|M_{i,j}|}$, and let $\frac{0}{0}=0$.
+    We consider the Output Agreement (OA) mechanism that directly counts the frequency of agreements between two agents’ responses.
   2. 2.
-    We use the Correlated Agreement (CA) mechanism as a baseline, which is introduced in Appendix A.
-The expected CA score is the unconditioned mutual information $I_{\text{TVD}}(\hat{X}_{i};\hat{X}_{j})$.
+    We use the Correlated Agreement (CA) mechanism as a baseline, which is introduced in Appendix B.
+The expected CA score is the unconditional mutual information $I_{\text{TVD}}(\hat{X}_{i};\hat{X}_{j})$.
   3. 3.
-    Another common method to estimate the reliability of the noisy agents is the Expectation-Maximization (EM) algorithm under the DS model (Dawid and Skene, 1979).
-Under the DS model, each agent $i$ makes mistakes according to a confusion matrix $\Gamma^{i}$ where $\Gamma^{i}_{h,l}$ denotes the probability that labeler $i$ reports $h$ when the ground truth answer is $l$.
-The EM algorithm first initializes a prior of the ground truth for each question and a confusion matrix for each agent.
-Then, it iteratively goes through an E-step and an M-step.
-In the E-step, the algorithm computes the posterior of the ground truth of each question given the prior and the confusion matrix.
-In the M-step, the confusion matrix of each agent is updated using the posterior computed in the E-step.
-The above two steps are iterated until the likelihood of the data matrix $\hat{X}$ converges.
-Finally, we score the reliability of agent $i$ as $\sum_{h\in\Sigma}\Gamma^{i}_{h,h}\tilde{P}(h)$ where $\tilde{P}(h)$ is the empirical marginal distribution of observing $h$ in the dataset.
+    Another common method to estimate the reliability of the noisy agents is the Expectation-Maximization (EM) algorithm under the DS model (dawid1979maximum).
+The algorithm iteratively updates a confusion matrix that captures the noise level of each agent.
+After convergence, we use the trace of each agent’s confusion matrix as a reliability score.
   4. 4.
     Lastly, we consider the Conditioned Agreement ($\text{OA}_{Z}$) which counts the frequency of agreements only when the agreement is different from the principal’s sample on that question.
-Similar to the agreement metric, the score of labeler $i$ is the average of the empirical probability of agreement conditioned on $Z$, i.e.
-$\frac{1}{n}\sum_{j\in[n]\backslash\{i\}}\sum_{q\in M_{i,j}}\frac{\mathbbm{1}\left[\hat{X}_{i,q}=\hat{X}_{j,q}\&\hat{X}_{i,q}\neq Z_{q}\right]}{|M_{i,j}|}$.
-Intuitively, if labeler $i$ uses an LLM whose generated labels are similar to $Z$, then $\hat{X}_{i,q}\neq Z_{q}$ will happen with a small probability, meaning that the conditioned agreement score will be small.
-We note that there are many more heuristic methods to estimate the reliability of crowdsourcing agents (Yin et al., 2008; Embretson and Reise, 2013; Piech et al., 2013).
+We note that there are many more heuristic methods to estimate the reliability of crowdsourcing agents 4415269; embretson2013item; Piech2013.
 Many of these methods are variants of the EM algorithm under a different context-dependent model.
@@
 5.3.3 Results
-(a) AUC v.s. fraction of LLM cheaters.
-(b) AUC v.s. sampled questions.
-Figure 4: The area under the ROC curve (AUC) for various methods on the preference alignment dataset as we vary (a) the fraction of LLM cheaters and (b) the proportion of questions the principal can sample.
-In (a), GPT-4 generates samples for $Z_{i}$, $Z_{j}$, and $Z$; in (b), AUC curves for all five LLM-cheater types are shown (omitting error bars for clarity).
-Given a dataset $\hat{X}$ and samples of the cheap signal $Z$, we use our method as well as the baselines to score each labeler.
-We use the AUC score to measure the ability of each discussed metric to distinguish the simulated cheaters from the original human labelers.
-Figure 4(a) presents a comparison of different measures while varying the fraction of LLM cheaters in the crowds.
-For each value of $\alpha_{llm}$ ranging from 00 to $0.2$, we randomly sample $\alpha_{r}$ and $\alpha_{b}$ uniformly from $[0,0.2]$ and report the average and error bars over $50$ such trials.
-Our result suggests that classic methods, such as the EM algorithm and the CA mechanism, perform well when the number of LLM cheaters is low.
-However, their AUCs degrade significantly as the fraction of LLM cheaters increases.
-On the other hand, the performance of $\text{OA}_{Z}$, which scores agents based on the similarity between their reports and $Z$, tends to have large variance and performs poorly when there are few LLM cheaters.
+Given a dataset $D$ and samples of the cheap signal $Z$, we use our method as well as the baselines to score each agent.
+We use the AUC score to measure the ability of each discussed metric to distinguish the simulated low-effort agents from the original human agents.
+In particular, let $I^{+}$ be the set of human agents and $I^{-}$ be the set of simulated low-effort agents.
+We compute the AUC score for each method as follows:
+$\mathrm{AUC}=\frac{1}{|I^{+}|\cdot|I^{-}|}\sum_{i\in I^{+}}\sum_{j\in I^{-}}\left(\mathds{1}[S_{i}>S_{j}]+0.5\cdot\mathds{1}[S_{i}=S_{j}]\right).$
+Figure 3(b) presents a comparison of different methods.
+For each value of $\alpha_{llm}$ ranging from 0 to $0.2$, we randomly sample $\alpha_{r}$ and $\alpha_{b}$ uniformly from $[0,0.2]$ and report the average and error bars over $50$ such trials.
+Our result suggests that classic methods, such as the EM algorithm and the CA mechanism, perform well when the number of LLM-reliant agents is low.
+However, their AUCs degrade significantly as the fraction of LLM-reliant agents increases.
+On the other hand, the performance of $\text{OA}_{Z}$, which scores agents based on the similarity between their reports and $Z$, tends to have large variance and performs poorly when there are few LLM-reliant agents.
 In contrast, our proposed conditioned CA mechanism has the most robust performance for the mixed crowds and has the lowest variance.
-Table 1 summarizes the results from Figure 4(a), reporting the average AUC score and the bottom 10% quantile for each method on each dataset w.r.t.
-the randomness of cheater fractions.
-A high average AUC reflects strong overall performance across random cheater compositions, while a high 10% quantile indicates robustness to the worst-case scenarios.
+Table 1 summarizes the results from Figure 3(b), reporting the average AUC score and the bottom 10% quantile for each method on each dataset, where the average is w.r.t.
+every tested combination of $(\alpha_{llm},\alpha_{r},\alpha_{b})$.
+A high average AUC reflects strong overall performance, while a high 10% quantile indicates robustness to the worst-case scenarios.
 The results support our earlier claim that our method consistently demonstrates the highest robustness across all tested settings.
-Table 1: Average AUC (and bottom 10% quantiles) across datasets and methods with GPT-4 cheaters.
+Table 1: Average AUC (and bottom 10% quantiles) across datasets and methods with GPT-4-reliant agents.
 Dataset OA CA OAZ EM CAZ (ours)
@@
 Preference Alignment 0.65 (0.40) 0.70 (0.43) 0.70 (0.45) 0.68 (0.41) 0.85 (0.77)
-Our experiments do not claim that our method outperforms all baselines in every scenario.
-Traditional methods like EM often perform better when all agents are random or biased cheaters.
+We note that our experiments do not claim that our method outperforms all baselines in every scenario.
+Traditional methods like EM often perform better when the crowd consists entirely of random or biased agents.
 However, our method demonstrates the most robust performance across all settings due to its theoretical guarantees.
 This is particularly useful as, in practice, the designer usually has little clue about the composition of the crowds.
-In contrast, the baselines exhibit cases where the AUC drops below 0.5, meaning the algorithm completely misclassifies good workers as cheating agents.
-Detection Performance vs. Number of Samples
-We further examine how many questions the principal needs to sample to effectively detect cheaters.
-In Fig.
-4(b), the AUC of the conditioned CA mechanism is shown with the LLM-cheater fraction fixed at 0.15, while the fraction of random and biased cheaters is each independently sampled from $[0,0.2]$.
-Similar plots for other datasets are provided in Section D.2.2.
-As the sampling rate increases from 20% to 100%, the average AUC also increases.
-While the results vary across the datasets and the LLMs that cheaters use, we observe that, except for the hatefulness labeling dataset with GPT-3.5 cheaters, sampling 50% of the questions is sufficient to outperform the baseline CA mechanism.
-Furthermore, in cases where human and AI labels are well separated, sampling just 20% can suffice.
+In contrast, the baselines exhibit cases where the AUC drops below 0.5, meaning the algorithm completely misclassifies good workers as low-effort agents.
 Additional results in the appendix.
-First, Section D.2.1 presents additional figures and tables for other datasets and LLMs, further supporting the claims made in the main body.
-Furthermore, when the principal is uncertain about which LLM the cheaters are using, we propose a method that computes the CA score conditioned on each candidate LLM’s outputs and uses the minimum score across them as the final score (Section D.2.3).
-Finally, in Appendix C and Section D.3, we present a heuristic extension of our method to open-ended textual responses.
+First, Section G.4.1 presents analogous figures and tables for additional datasets and LLMs, further supporting the claims made in the main body.
+Second, we investigate how many questions the principal needs to sample for effective detection of low-effort agents (Section G.4.2).
+The results suggest that sampling LLM labels on 50% of the questions is typically sufficient to outperform the CA baseline.
+In cases where human and AI labels are well separated, sampling just 20% can suffice.
+Third, when the principal is uncertain about which LLMs the agents use, we propose a method that computes the CA score conditioned on each candidate LLM’s outputs and uses the minimum score across them as the final score (Section G.4.3).
+Finally, in Appendix F and Section G.5, we present a heuristic extension of our method to open-ended textual responses.
 This serves as a preliminary step toward generalizing our approach beyond labeling tasks, with further refinement and large-scale validation left for future work.
-6 Limitations
+6 Conclusion
+This work addresses the challenge of LLM contamination in crowdsourcing by developing a theoretically grounded quality measure for annotation tasks.
+Leveraging an extension of peer prediction, our method evaluates worker responses beyond what LLMs can generate, ensuring reliable human contributions.
+We establish theoretical guarantees under different low-effort strategies and validate our approach empirically on real-world datasets.
+Our results demonstrate that the proposed method effectively detects low-effort agents while maintaining robustness across various conditions.
+Future work can focus on refining our method to better differentiate between harmful LLM usage and productive human-AI collaboration.
+This includes optimizing scoring mechanisms to recognize when LLMs are used as assistive tools rather than substitutes for human effort and developing adaptive detection techniques that adjust based on task characteristics and worker behavior.
+Broader Impact
+On the positive side, our work can improve the robustness and fairness of data collection pipelines used in training AI systems, thereby enhancing the reliability of downstream decision-making or other usage of AI.
+However, there is a potential risk that insights into reporting strategies or detection weaknesses could be misused to evade quality control.
+We believe this risk is mitigated by focusing on defensive mechanisms and open discussion of vulnerabilities to promote transparency and better system design.
+Furthermore, like many crowdsourcing algorithms for label-quality control, our method operates without access to ground truth.
+Consequently, a low score does not always indicate a low-effort worker, while it may also reflect a diligent agent holding a minority opinion.
+Therefore, while such verification-free algorithms are highly scalable, they should be used with a human-in-the-loop, allowing the principal to carefully review low-scoring agents and prevent potential ethical concerns.
+Acknowledgment
+This work was partly carried out at the DIMACS Center at Rutgers University.
+Appendix A Limitations
 We acknowledge several limitations of our method.
 First, our method is primarily designed for crowdsourcing datasets comprising subjective questions or tasks where human experts are considered the gold standard.
-If LLMs significantly surpass human performance, the correlations between human responses conditioned on $Z$ may no longer reflect the ground truth $Y$.
+If LLMs significantly surpass human performance, the correlations between human responses may no longer reflect the ground truth $Y$.
+This violates the assumption that, conditioned on $Y$, agents’ high-effort signals are independent.
 In such cases, applying our method could mistakenly reward incorrect correlations.
@@
 In particular, when LLM responses are diverse, e.g.
-every cheating agent applies a different model or there are only a few cheating agents, heuristic methods such as the EM algorithm work well.
-This is because diverse or sparse cheaters are less correlated with the majority of the crowd and can be easily flagged as outliers using the classic methods.
+when each agent relies on a different model or only a few agents use LLM assistance, heuristic methods such as the EM algorithm work well.
+This is because diverse or sparse low-effort agents are less correlated with the majority of the crowd and can be easily flagged as outliers using the classic methods.
 Despite this, our method demonstrates the most robust performance across a wide range of settings.
-Future work could explore integrating the principal’s samples of potential cheating signals into heuristic approaches like EM to enhance their performance and optimize detection.
+Future work could incorporate the principal’s sampled LLM signals into heuristic methods such as EM to improve their performance and enhance detection under LLM contamination.
 Third, our experiments evaluate only five language models, as some models, such as LLaMA and Gemini, refuse to respond to toxicity-related prompts.
 While this is sufficient to demonstrate the method’s effectiveness, additional evaluations across more diverse models and task types would provide a more comprehensive assessment.
-7 Conclusion
-This work addresses the challenge of LLM-assisted cheating in crowdsourcing by developing a theoretically grounded quality measure for annotation tasks.
-Leveraging an extension of peer prediction, our method evaluates human information beyond what LLMs can generate, ensuring reliable human contributions.
-We establish theoretical guarantees under different cheating strategies and validate our approach empirically on real-world datasets.
-Our results demonstrate that our method effectively detects low-effort cheating while maintaining robustness across various conditions.
-Future work can focus on refining our method to better differentiate between harmful cheating and productive human-AI collaboration.
-This includes optimizing scoring mechanisms to recognize when LLMs are used as assistive tools rather than substitutes for human effort and developing adaptive detection techniques that adjust based on task characteristics and worker behavior.
-Acknowledgment
-This work was partly carried out at the DIMACS Center at Rutgers University. Jinlong Pang and Yang Liu are partially supported by the National Science Foundation (NSF) under grant IIS-2143895.
-Appendix A Preliminary
+Appendix B Preliminary
 What do we know about the problem?
-Under the assumption that there are no cheap signals, the peer prediction literature suggests that estimating the mutual information between $X_{i}$ and $X_{j}$ is an information monotone metric [Kong and Schoenebeck, 2019a].
-There are various ways to estimate the mutual information using agents’ responses to multiple questions [Kong, 2020a, Schoenebeck and Yu, 2020] and many information monotone metrics that can be explained using mutual information [Dasgupta and Ghosh, 2013, Shnayder et al., 2016].
-In this section, we discuss a simple yet intuitive mechanism called the correlated agreement (CA) mechanism [Shnayder et al., 2016].
+Our work builds on advancements in information elicitation, particularly peer prediction.
+First introduced by miller2005eliciting, peer prediction mechanisms aim to incentivize truthful reporting without requiring ground truth verification.
+The core idea is to reward agents based on how well their responses predict those of others.
+The original mechanism relied on the principal’s knowledge of the correlation between agents’ signals (the information they observe).
+Later, dasgupta2013crowdsourced extended this idea, removing the need for such knowledge by assuming agents answer multiple i.i.d.
+binary questions.
+Subsequent works have expanded this framework to non-binary (but finite) signals [shnayder2016informed, kong2019information], infinite signals [schoenebeck2020learning], truthful guarantees with fewer questions [kong2020dominantly], heterogeneous agents [Agarwal2017], adversarial agents [rowdycrowds], and more general reporting strategies [zhangomni2023].
+Under the assumption that there are no cheap signals, the peer prediction literature suggests that estimating the mutual information between $X_{i}$ and $X_{j}$ is an information monotone metric FrameworkKong2019.
+There are various ways to estimate the mutual information using agents’ responses to multiple questions dmiKong2020, schoenebeck2020learning and many information monotone metrics that can be explained using mutual information dasgupta2013crowdsourced, shnayder2016informed.
+In this section, we discuss a simple yet intuitive mechanism called the correlated agreement (CA) mechanism [shnayder2016informed].
 The CA mechanism rewards agents when they “agree” on the same question and penalizes them when they “agree” on distinct questions.
 Instead of directly checking agreement, the mechanism redefines agreement based on the correlation between signals.
-A pair of signals $\sigma,\sigma^{\prime}\in\Sigma$ is correlated agreed if $\Pr(X_{1}=\sigma,X_{2}=\sigma^{\prime})-\Pr(X_{1}=\sigma)\Pr(X_{2}=\sigma^{\prime})>0$, meaning that $\sigma$ and $\sigma^{\prime}$ are more likely to appear on the same question rather than on distinct questions.
-In particular, the mechanism utilizes the delta matrix $\Delta_{\sigma,\sigma^{\prime}}=\Pr(X_{1}=\sigma,X_{2}=\sigma^{\prime})-\Pr(X_{1}=\sigma)\Pr(X_{2}=\sigma^{\prime})$ to design the scoring function $T_{\Delta}=\text{Sign}(\Delta)$ where $\text{Sign}(x)=1$ if $x>0$ and 0 otherwise.
+A pair of signals $\sigma,\sigma^{\prime}\in\Sigma$ is correlated agreed if $\Pr(X_{i}=\sigma,X_{j}=\sigma^{\prime})-\Pr(X_{i}=\sigma)\Pr(X_{j}=\sigma^{\prime})>0$, meaning that $\sigma$ and $\sigma^{\prime}$ are more likely to appear on the same question rather than on distinct questions.
+In particular, the mechanism utilizes the delta matrix $\Delta_{\sigma,\sigma^{\prime}}=\Pr(X_{i}=\sigma,X_{j}=\sigma^{\prime})-\Pr(X_{i}=\sigma)\Pr(X_{j}=\sigma^{\prime})$ to design the scoring function $T_{\Delta}=\operatorname{sign}(\Delta)$ where $\operatorname{sign}(x)=1$ if $x>0$ and 0 otherwise.
 In practice, the mechanism can be implemented in two ways depending on the information available to the principal.
@@
 when questions have binary answers, $T_{\Delta}$ is likely to be a $2\times 2$ identical matrix.
-Second, if $T_{\Delta}$ is unknown, the mechanism can first estimate the joint distribution, denoted as $\widetilde{\Pr}(X_{1},X_{2})$ using the report matrix $\hat{X}$.
-Then, an estimate of $T_{\Delta}$ can be computed using the estimated Delta matrix $\tilde{\Delta}_{\sigma,\sigma^{\prime}}=\widetilde{\Pr}(X_{1}=\sigma,X_{2}=\sigma^{\prime})-\widetilde{\Pr}(X_{1}=\sigma)\Pr(X_{2}=\sigma^{\prime})$.
+Second, if $T_{\Delta}$ is unknown, the mechanism can first estimate the joint distribution, denoted as $\widetilde{\Pr}(\hat{X}_{i},\hat{X}_{j})$ using the report matrix $\hat{X}$.
+Then, an estimate of $T_{\Delta}$ can be computed using the estimated Delta matrix $\tilde{\Delta}_{\sigma,\sigma^{\prime}}=\widetilde{\Pr}(\hat{X}_{i}=\sigma,\hat{X}_{j}=\sigma^{\prime})-\widetilde{\Pr}(\hat{X}_{i}=\sigma)\widetilde{\Pr}(\hat{X}_{j}=\sigma^{\prime})$.
 Next, Algorithm 1 can be applied.
@@
 return $s$
-Algorithm 2 The Correlated Agreement Mechanism [Shnayder et al., 2016]
-Appendix B Proofs
-B.1 Sufficiency of the Assumptions
+Algorithm 1 The Correlated Agreement Mechanism [shnayder2016informed]
+Appendix C Details of the Conditioned Correlated Agreement Mechanism
+After learning the scoring function, the conditioned CA mechanism scores agents according to Algorithm 2.
+At a high level, the mechanism will iteratively fix a signal $k\in\Sigma$ and look at the questions on which $Z_{q}=k$.
+Let $M_{Z}(k)$ be the set of questions on which $Z=k$.
+Then, an agent $i$ will receive a higher score if she agrees with another agent $j$ on the same question more often than on two distinct questions conditioning that the questions are drawn from $M_{Z}(k)$.
+This measures the additional correlation between two independent agents conditioned on $Z$.
+Input: The crowdsourced dataset $\hat{X}$, the samples of cheap signal $Z$, the (estimated) scoring function $T_{\tilde{\Delta}}$
+Estimate the empirical marginal distribution of $Z$, i.e. $\tilde{P}(Z=\sigma)$ for $\sigma\in\Sigma$.
+for agent $i\in[n]$ do
+   Let $M_{i}$ be the set of questions answered by agent $i$;
+   Initialize $s_{i}=0$;
+ for $k\in\Sigma$ do
+      If $M_{i}\cap M_{Z}(k)=\emptyset$, continue;
+      Initialize $x=0$;
+    for question $q\in M_{i}\cap M_{Z}(k)$ do
+         Randomly select a peer $j$ who answers $q$;
+       $x\mathrel{+}=T_{\tilde{\Delta}}(\hat{X}_{i,q},\hat{X}_{j,q},k)$;
+         Randomly select a question $q^{\prime}\in M_{j}\cap M_{Z}(k)$;
+       $x\mathrel{-}=T_{\tilde{\Delta}}(\hat{X}_{i,q},\hat{X}_{j,q^{\prime}},k)$
+    $s_{i}\mathrel{+}=x\cdot\tilde{P}(Z=k)/|M_{i}\cap M_{Z}(k)|$
+return $s=(s_{1},\ldots,s_{n})$
+Algorithm 2 The conditioned CA mechanism
+Appendix D Proofs
+D.1 Sufficiency of the Assumptions
 Proof of Theorem 4.3.
@@
 $=$ $\sum_{z\in\Sigma}\Pr(Z=z)\sum_{x_{i},x_{j}\in\Sigma}(\Pr(x_{i},x_{j}\mid z)-\Pr(x_{i}\mid z)\;\Pr(x_{j}\mid z))\;T_{\Delta}(x_{i},x_{j},z)$
-$=$ $\sum_{z\in\Sigma}\Pr(Z=z)\sum_{x_{i},x_{j}\in\Sigma}|\Pr(x_{i},x_{j}\mid z)-\Pr(x_{i}\mid z)|$
+$=$ $\sum_{z\in\Sigma}\Pr(Z=z)\sum_{x_{i},x_{j}\in\Sigma}|\Pr(x_{i},x_{j}\mid z)-\Pr(x_{i}\mid z)\;\Pr(x_{j}\mid z)|$
 $=$ $I_{\text{TVD}}(X_{i};X_{j}\mid Z)$
@@
 Lower bound the penalty score.  We first provide a lower bound for $\Pr(X_{i},Z_{i}\mid Z)$, which is summarized in the following lemma.
-Lemma B.1.
+Lemma D.1.
 Under 4.1, $\Pr(X_{i},Z_{i}\mid Z)\geq\Pr(Z_{i}\mid Z)\Pr(X_{i}\mid Z)-\epsilon_{i}\sum_{y\in\Sigma}\Pr(X_{i}\mid Y=y,Z)$.
@@
 $\sum_{x_{i},x_{j},z_{i},z_{j}\in\Sigma}\Pr(x_{i},z_{i}\mid z)\;\Pr(x_{j},z_{j}\mid z)\;T_{\Delta}(\theta_{i}(x_{i},z_{i}),\theta_{j}(x_{j},z_{j}),z)$
-$\geq$ $\sum_{x_{i},x_{j},z_{i},z_{j}\in\Sigma}\left(\Pr(z_{i}\mid z)\Pr(x_{i}\mid z)-\epsilon_{i}\sum_{y\in\Sigma}\Pr(x_{i}\mid y,z)\right)\;\Pr(x_{j},z_{j}\mid z)\;T_{\Delta}(\theta_{i}(x_{i},z_{i}),\theta_{j}(x_{j},z_{j}),z).$ (By Lemma B.1)
+$\geq$ $\sum_{x_{i},x_{j},z_{i},z_{j}\in\Sigma}\left(\Pr(z_{i}\mid z)\Pr(x_{i}\mid z)-\epsilon_{i}\sum_{y\in\Sigma}\Pr(x_{i}\mid y,z)\right)\;\Pr(x_{j},z_{j}\mid z)\;T_{\Delta}(\theta_{i}(x_{i},z_{i}),\theta_{j}(x_{j},z_{j}),z).$ (By Lemma D.1)
 By reordering the summations,
@@
 First note that $T_{\Delta}\leq 1$.
-Therefore, the second term is lower bounded by $-\epsilon_{i}\sum_{y,z_{i}\in\Sigma}\sum_{x_{j},z_{j}\in\Sigma}\Pr(x_{j},z_{j}\mid z)\sum_{x_{i}\in\Sigma}\Pr(x_{i}\mid y,z)$.
+Therefore, the second term is lower bounded by $\epsilon_{i}\sum_{y,z_{i}\in\Sigma}\sum_{x_{j},z_{j}\in\Sigma}\Pr(x_{j},z_{j}\mid z)\sum_{x_{i}\in\Sigma}\Pr(x_{i}\mid y,z)$.
 Marginalizing over $x_{j},z_{j}$ and $x_{i}$ and observing that $y,z_{i}\in\Sigma$, the quality is lower-bounded by $\epsilon_{i}|\Sigma|^{2}$.
@@
 $=$ $\sum_{x_{i},x_{j},z_{i},z_{j}\in\Sigma}\Pr(z_{i}\mid z)\Pr(x_{i}\mid z)\;\left(\Pr(z_{j}\mid z)\Pr(x_{j}\mid z)-\epsilon_{j}\sum_{y\in\Sigma}\Pr(x_{j}\mid y,z)\right)$ (1)
-$\qquad\qquad\cdot T_{\Delta}(\theta_{i}(x_{i},z_{i}),\theta_{j}(x_{j},z_{j}),z)-\epsilon_{i}|\Sigma|^{2}$ (By Lemma B.1)
+$\qquad\qquad\cdot T_{\Delta}(\theta_{i}(x_{i},z_{i}),\theta_{j}(x_{j},z_{j}),z)-\epsilon_{i}|\Sigma|^{2}$ (By Lemma D.1)
 $\geq$ $\sum_{x_{i},x_{j},z_{i},z_{j}\in\Sigma}\Pr(z_{i}\mid z)\Pr(x_{i}\mid z)\Pr(z_{j}\mid z)\Pr(x_{j}\mid z)\;T_{\Delta}(\theta_{i}(x_{i},z_{i}),\theta_{j}(x_{j},z_{j}),z)-(\epsilon_{i}+\epsilon_{j})|\Sigma|^{2}.$ (2)
 Upper bound the bonus score. We first present a lemma that bounds the difference between $\Pr(Z_{i},Z_{j}\mid Y,Z)$ and $\Pr(Z_{i}\mid Z)\Pr(Z_{j}\mid Z)$.
-Lemma B.2.
+Lemma D.2.
 Under 4.1 and 4.2, $\left|\Pr(Z_{i},Z_{j}\mid Y,Z)-\Pr(Z_{i}\mid Z)\Pr(Z_{j}\mid Z)\right|\leq\epsilon+\frac{\epsilon_{i}+\epsilon_{j}}{\Pr(Y|Z)}$.
@@
 $=$ $\sum_{x_{i},x_{j},z_{i},z_{j}\in\Sigma}\sum_{y\in\Sigma}\Pr(y\mid z)\Pr(x_{i},x_{j}\mid y,z)\Pr(z_{i},z_{j}\mid y,z)\;T_{\Delta}(\theta_{i}(x_{i},z_{i}),\theta_{j}(x_{j},z_{j}),z)$ ($X$ and $Z$ are independent conditioned on $Y$.)
-$\leq$ $\sum_{x_{i},x_{j},z_{i},z_{j}\in\Sigma}\sum_{y\in\Sigma}\Pr(y\mid z)\Pr(x_{i},x_{j}\mid y,z)\left(\Pr(z_{i}\mid z)\Pr(z_{j}\mid z)+\epsilon+\frac{\epsilon_{i}+\epsilon_{j}}{\Pr(y|z)}\right)\;T_{\Delta}(\theta_{i}(x_{i},z_{i}),\theta_{j}(x_{j},z_{j}),z).$ (By Lemma B.2)
+$\leq$ $\sum_{x_{i},x_{j},z_{i},z_{j}\in\Sigma}\sum_{y\in\Sigma}\Pr(y\mid z)\Pr(x_{i},x_{j}\mid y,z)\left(\Pr(z_{i}\mid z)\Pr(z_{j}\mid z)+\epsilon+\frac{\epsilon_{i}+\epsilon_{j}}{\Pr(y|z)}\right)\;T_{\Delta}(\theta_{i}(x_{i},z_{i}),\theta_{j}(x_{j},z_{j}),z).$ (By Lemma D.2)
 We can further write this into three terms.
@@
 First note that
-Eq. 3 $=\sum_{x_{i},x_{j},z_{i},z_{j}\in\Sigma}\Pr(z_{i}\mid z)\Pr(z_{j}\mid z)\;T_{\Delta}(\theta_{i}(x_{i},z_{i}),\theta_{j}(x_{j},z_{j}),z)\sum_{y\in\Sigma}\Pr(x_{i},x_{j},y\mid z)$
+Equation 3 $=\sum_{x_{i},x_{j},z_{i},z_{j}\in\Sigma}\Pr(z_{i}\mid z)\Pr(z_{j}\mid z)\;T_{\Delta}(\theta_{i}(x_{i},z_{i}),\theta_{j}(x_{j},z_{j}),z)\sum_{y\in\Sigma}\Pr(x_{i},x_{j},y\mid z)$
 $=\sum_{x_{i},x_{j},z_{i},z_{j}\in\Sigma}\Pr(z_{i}\mid z)\Pr(z_{j}\mid z)\;T_{\Delta}(\theta_{i}(x_{i},z_{i}),\theta_{j}(x_{j},z_{j}),z)\Pr(x_{i},x_{j}\mid z)$
 Next, we upper bound the second and the third term respectively. By reordering the summations,
-Eq. 4 $=\epsilon\;\sum_{x_{i},x_{j},z_{i},z_{j}\in\Sigma}T_{\Delta}(\theta_{i}(x_{i},z_{i}),\theta_{j}(x_{j},z_{j}),z)\;\sum_{y\in\Sigma}\Pr(x_{i},x_{j},y\mid z)$
+Equation 4 $=\epsilon\;\sum_{x_{i},x_{j},z_{i},z_{j}\in\Sigma}T_{\Delta}(\theta_{i}(x_{i},z_{i}),\theta_{j}(x_{j},z_{j}),z)\;\sum_{y\in\Sigma}\Pr(x_{i},x_{j},y\mid z)$
 $=\epsilon\;\sum_{x_{i},x_{j},z_{i},z_{j}\in\Sigma}T_{\Delta}(\theta_{i}(x_{i},z_{i}),\theta_{j}(x_{j},z_{j}),z)\;\Pr(x_{i},x_{j}\mid z)$
@@
 $\leq\epsilon|\Sigma|^{2}.$ ($\sum_{x_{i},x_{j}\in\Sigma}\Pr(x_{i},x_{j}\mid z)=1$)
-Eq. 5 $=(\epsilon_{i}+\epsilon_{j})\;\sum_{y\in\Sigma}\sum_{z_{i},z_{j}\in\Sigma}\sum_{x_{i},x_{j}\in\Sigma}\Pr(x_{i},x_{j}\mid y,z)T_{\Delta}(\theta_{i}(x_{i},z_{i}),\theta_{j}(x_{j},z_{j}),z)$
+Equation 5 $=(\epsilon_{i}+\epsilon_{j})\;\sum_{y\in\Sigma}\sum_{z_{i},z_{j}\in\Sigma}\sum_{x_{i},x_{j}\in\Sigma}\Pr(x_{i},x_{j}\mid y,z)T_{\Delta}(\theta_{i}(x_{i},z_{i}),\theta_{j}(x_{j},z_{j}),z)$
 $\leq(\epsilon_{i}+\epsilon_{j})\;\sum_{y\in\Sigma}\sum_{z_{i},z_{j}\in\Sigma}\sum_{x_{i},x_{j}\in\Sigma}\Pr(x_{i},x_{j}\mid y,z)$ ($T_{\Delta}\leq 1$)
@@
 $\leq$ $\sum_{x_{i},x_{j},z_{i},z_{j}\in\Sigma}\Pr(z_{i}\mid z)\Pr(z_{j}\mid z)\;\Pr(x_{i},x_{j}\mid z)\;T_{\Delta}(\theta_{i}(x_{i},z_{i}),\theta_{j}(x_{j},z_{j}),z)+\epsilon|\Sigma|^{2}+(\epsilon_{i}+\epsilon_{j})|\Sigma|^{3}.$ (6)
-Finally, combining Eq. 2 and Eq. 6,
+Finally, combining Equation 2 and Equation 6,
 $S_{i}(\theta_{i},\theta_{j})$
@@
 $=$ $S_{i}(\tau,\tau)+\hat{\epsilon}$.
-Lastly, we show that if agent $i$ plays a naive strategy $\nu_{i}$, the expected score is strictly smaller than truth-telling.
-Note that because a naive strategy is a convex combination between truth-telling $\tau$ and an uninformative strategy $\mu_{i}$, it suffices to show that the expected score of any uninformative strategy is strictly smaller than truth-telling.
-Because $\mu_{i}(X_{i},Z_{i})$ is independent of $X_{i}$, we can reoder the summations in Eq.
-7 in the following way.
+Lastly, we show that if agent $i$ plays a lazy-reporting strategy $\nu_{i}$, the expected score is strictly smaller than truth-telling.
+Note that because a lazy-reporting strategy is a convex combination between truth-telling $\tau$ and a no-effort strategy $\mu_{i}$, it suffices to show that the expected score of any no-effort strategy is strictly smaller than truth-telling.
+Because $\mu_{i}(X_{i},Z_{i})$ is independent of $X_{i}$, we can reoder the summations in Equation 7 in the following way.
 Let $\mu=\mu_{i}(X_{i},Z_{i})$.
@@
 $\leq$ $\sum_{z\in\Sigma}\Pr(Z=z)\sum_{z_{i},z_{j}\in\Sigma}\Pr(z_{i}\mid z)\Pr(z_{j}\mid z)\sum_{x_{j}\in\Sigma}T_{\Delta}(\mu,\theta_{j}(x_{j},z_{j}),z)$
-$\sum_{x_{i}\in\Sigma}\left(\Pr(x_{i},x_{j}\mid z)-\Pr(x_{i}\mid z)\Pr(x_{j}\mid z)\right)+\hat{\epsilon}$ (by Eq. 7)
+$\sum_{x_{i}\in\Sigma}\left(\Pr(x_{i},x_{j}\mid z)-\Pr(x_{i}\mid z)\Pr(x_{j}\mid z)\right)+\hat{\epsilon}$ (by Equation 7)
 $=$ $\sum_{z\in\Sigma}\Pr(Z=z)\sum_{z_{i},z_{j}\in\Sigma}\Pr(z_{i}\mid z)\Pr(z_{j}\mid z)\sum_{x_{j}\in\Sigma}T_{\Delta}(\mu,\theta_{j}(x_{j},z_{j}),z)\left(\Pr(x_{j}\mid z)-\Pr(x_{j}\mid z)\right)+\hat{\epsilon}$
@@
 Note that the expected score under truth-telling is exactly $I_{\text{TVD}}(X_{i};X_{j}\mid Z)$.
-Therefore, whenever $I_{\text{TVD}}(X_{i};X_{j}\mid Z)>\hat{\epsilon}$, we have $S_{i}(\tau,\tau)>S_{i}(\mu_{i},\theta_{j})$ for any uninformative strategy $\mu_{i}$ and any strategy $\theta_{j}$.
+Therefore, whenever $I_{\text{TVD}}(X_{i};X_{j}\mid Z)>\hat{\epsilon}$, we have $S_{i}(\tau,\tau)>S_{i}(\mu_{i},\theta_{j})$ for any no-effort strategy $\mu_{i}$ and any strategy $\theta_{j}$.
 This completes the proof.
 ∎
-Proof of Lemma B.1.
-$\Pr(X_{i},Z_{i}\mid Z)$ $=\sum_{y\in\Sigma}\Pr(Y=y\mid Z)\Pr(X_{i},Z_{i}\mid Y=y,Z)$ (By the law to total probability)
+Proof of Lemma D.1.
+$\Pr(X_{i},Z_{i}\mid Z)$ $=\sum_{y\in\Sigma}\Pr(Y=y\mid Z)\Pr(X_{i},Z_{i}\mid Y=y,Z)$ (By the law of total probability)
 $=\sum_{y\in\Sigma}\Pr(Y=y\mid Z)\Pr(X_{i}\mid Y=y,Z)\Pr(Z_{i}\mid Y=y,Z)$ ($X_{i}$ and $Z_{i}$ are independent conditioned on $Y$.)
@@
 $=\Pr(Z_{i}\mid Z)\Pr(X_{i}\mid Z)-\epsilon_{i}\sum_{y\in\Sigma}\Pr(X_{i}\mid Y=y,Z).$
-Proof of Lemma B.2.
-$\left|\Pr(Z_{1},Z_{2}\mid Y,Z)-\Pr(Z_{1}\mid Z)\Pr(Z_{2}\mid Z)\right|$
-$=$ $\left|\Pr(Z_{1},Z_{2}\mid Y,Z)-\Pr(Z_{1}\mid Y,Z)\Pr(Z_{2}\mid Y,Z)+\Pr(Z_{1}\mid Y,Z)\Pr(Z_{2}\mid Y,Z)-\Pr(Z_{1}\mid Z)\Pr(Z_{2}\mid Z)\right|$
-$\leq$ $\left|\Pr(Z_{1},Z_{2}\mid Y,Z)-\Pr(Z_{1}\mid Y,Z)\Pr(Z_{2}\mid Y,Z)\right|+\left|\Pr(Z_{1}\mid Y,Z)\Pr(Z_{2}\mid Y,Z)-\Pr(Z_{1}\mid Z)\Pr(Z_{2}\mid Z)\right|$ (Triangle inequality)
-$\leq$ $\epsilon+\left|\Pr(Z_{1}\mid Y,Z)\Pr(Z_{2}\mid Y,Z)-\Pr(Z_{1}\mid Z)\Pr(Z_{2}\mid Z)\right|.$ (4.2)
+Proof of Lemma D.2.
+$\left|\Pr(Z_{i},Z_{j}\mid Y,Z)-\Pr(Z_{i}\mid Z)\Pr(Z_{j}\mid Z)\right|$
+$=$ $\left|\Pr(Z_{i},Z_{j}\mid Y,Z)-\Pr(Z_{i}\mid Y,Z)\Pr(Z_{j}\mid Y,Z)+\Pr(Z_{i}\mid Y,Z)\Pr(Z_{j}\mid Y,Z)-\Pr(Z_{i}\mid Z)\Pr(Z_{j}\mid Z)\right|$
+$\leq$ $\left|\Pr(Z_{i},Z_{j}\mid Y,Z)-\Pr(Z_{i}\mid Y,Z)\Pr(Z_{j}\mid Y,Z)\right|+\left|\Pr(Z_{i}\mid Y,Z)\Pr(Z_{j}\mid Y,Z)-\Pr(Z_{i}\mid Z)\Pr(Z_{j}\mid Z)\right|$ (Triangle inequality)
+$\leq$ $\epsilon+\left|\Pr(Z_{i}\mid Y,Z)\Pr(Z_{j}\mid Y,Z)-\Pr(Z_{i}\mid Z)\Pr(Z_{j}\mid Z)\right|.$ (4.2)
 Therefore, we only have to bound the second quantity.
-$\left|\Pr(Z_{1}\mid Y,Z)\Pr(Z_{2}\mid Y,Z)-\Pr(Z_{1}\mid Z)\Pr(Z_{2}\mid Z)\right|$
-$=$ $\left|\Pr(Z_{1}\mid Y,Z)(\Pr(Z_{2}\mid Y,Z)-\Pr(Z_{2}\mid Z))+\Pr(Z_{2}\mid Z)(\Pr(Z_{1}\mid Y,Z)-\Pr(Z_{1}\mid Z))\right|$
-$\leq$ $\left|\Pr(Z_{1}\mid Y,Z)(\Pr(Z_{2}\mid Y,Z)-\Pr(Z_{2}\mid Z))\right|+\left|\Pr(Z_{2}\mid Z)(\Pr(Z_{1}\mid Y,Z)-\Pr(Z_{1}\mid Z))\right|$ (Triangle inequality)
-$\leq$ $\left|\Pr(Z_{2}\mid Y,Z)-\Pr(Z_{2}\mid Z)\right|+\left|\Pr(Z_{1}\mid Y,Z)-\Pr(Z_{1}\mid Z)\right|$
-$=$ $\frac{1}{\Pr(Y\mid Z)}\;\left|\Pr(Z_{2},Y\mid Z)-\Pr(Y\mid Z)\;\Pr(Z_{2}\mid Z)\right|+\frac{1}{\Pr(Y\mid Z)}\left|\Pr(Z_{1},Y\mid Z)-\Pr(Y\mid Z)\;\Pr(Z_{1}\mid Z)\right|$
+$\left|\Pr(Z_{i}\mid Y,Z)\Pr(Z_{j}\mid Y,Z)-\Pr(Z_{i}\mid Z)\Pr(Z_{j}\mid Z)\right|$
+$=$ $\left|\Pr(Z_{i}\mid Y,Z)(\Pr(Z_{j}\mid Y,Z)-\Pr(Z_{j}\mid Z))+\Pr(Z_{j}\mid Z)(\Pr(Z_{i}\mid Y,Z)-\Pr(Z_{i}\mid Z))\right|$
+$\leq$ $\left|\Pr(Z_{i}\mid Y,Z)(\Pr(Z_{j}\mid Y,Z)-\Pr(Z_{j}\mid Z))\right|+\left|\Pr(Z_{j}\mid Z)(\Pr(Z_{i}\mid Y,Z)-\Pr(Z_{i}\mid Z))\right|$ (Triangle inequality)
+$\leq$ $\left|\Pr(Z_{j}\mid Y,Z)-\Pr(Z_{j}\mid Z)\right|+\left|\Pr(Z_{i}\mid Y,Z)-\Pr(Z_{i}\mid Z)\right|$
+$=$ $\frac{1}{\Pr(Y\mid Z)}\;\left|\Pr(Z_{j},Y\mid Z)-\Pr(Y\mid Z)\;\Pr(Z_{j}\mid Z)\right|+\frac{1}{\Pr(Y\mid Z)}\left|\Pr(Z_{i},Y\mid Z)-\Pr(Y\mid Z)\;\Pr(Z_{i}\mid Z)\right|$
 $\leq$ $\frac{\epsilon_{1}+\epsilon_{2}}{\Pr(Y\mid Z)}.$ (By 4.1)
 Combining everything, we complete the proof. ∎
-B.2 Necessity of the Assumptions
+D.2 Necessity of the Assumptions
 Proof of 4.4.
@@
 Now, we compute the expected score when both agents honestly report the high-effort signal.
-$S_{i}(\tau,\tau)=$ $\sum_{z\in\{0,1\}}\Pr(Z=z)\sum_{x_{i},x_{j}\in\{0,1\}}(\Pr(X_{i}=x_{i},X_{j}=x_{j}\mid z)-\Pr(X_{i}=x_{i}\mid z)\;\Pr(X_{j}=x_{j}\mid z))\;T_{\Delta}(x_{i},x_{j},z)$
+$S_{i}(X_{i},X_{j}\mid Z)=$ $\sum_{z\in\{0,1\}}\Pr(Z=z)$
+$\quad\sum_{x_{i},x_{j}\in\{0,1\}}(\Pr(X_{i}=x_{i},X_{j}=x_{j}\mid z)-\Pr(X_{i}=x_{i}\mid z)\;\Pr(X_{j}=x_{j}\mid z))\;T_{\Delta}(x_{i},x_{j},z)$
 $=\sum_{x_{i},x_{j}\in\{0,1\}}\left|\Pr(x_{i},x_{j}\mid z)-\Pr(x_{i}\mid z)\;\Pr(x_{j}\mid z)\right|$
 $=8d(1-3d).$
-We show that this score is strictly smaller than the expected score when agents report 00 if and only if $X_{i}=Z_{i}=0$ (and $X_{j}=Z_{j}=0$) and report $1$ otherwise.
+We show that this score is strictly smaller than the expected score when agents report 0 if and only if $X_{i}=Z_{i}=0$ (and $X_{j}=Z_{j}=0$) and report $1$ otherwise.
 Let $\theta$ be the strategy such that $\theta(k,l|z)=0$ if and only if $k=l=0$ for any $z$.
 We can write down the joint distribution and the product of marginal distributions for $\hat{X}_{i}$ and $\hat{X}_{j}$.
-$\Pr(\hat{X}_{i},\hat{X}_{j}\mid y,z)=\begin{pmatrix}2d&0\\ 0&1-2d\end{pmatrix}\qquad\Pr(\hat{X}_{i},\hat{X}_{j}\mid y,z)=\begin{pmatrix}4d^{2}&2d(1-2d)\\ 2d(1-2d)&(1-2d)^{2}\end{pmatrix}\quad\text{for any $y,z\in\{0,1\}$.}$
+$\Pr(\hat{X}_{i},\hat{X}_{j}\mid y,z)=\begin{pmatrix}2d&0\\ 0&1-2d\end{pmatrix}$
+$\Pr(\hat{X}_{i}\mid y,z)\Pr(\hat{X}_{j}\mid y,z)=\begin{pmatrix}4d^{2}&2d(1-2d)\\ 2d(1-2d)&(1-2d)^{2}\end{pmatrix}\quad\text{for any $y,z\in\{0,1\}$.}$
 The difference between the above two matrices is thus
@@
 This means that the expected score under the strategy profile $(\theta,\theta)$ is
-$S_{i}(\theta,\theta)=$ $\sum_{z\in\{0,1\}}\Pr(Z=z)\sum_{x_{i},x_{j}\in\{0,1\}}(\Pr((\hat{X}_{i}=x_{i},\hat{X}_{j}=x_{j}\mid z)-\Pr(\hat{X}_{i}=x_{i}\mid z)\;\Pr(\hat{X}_{j}=x_{j}\mid z))\;T_{\Delta}(x_{i},x_{j},z)$
+$S_{i}(\hat{X}_{i},\hat{X}_{j}\mid Z)=$ $\sum_{z\in\{0,1\}}\Pr(Z=z)$
+$\quad\sum_{x_{i},x_{j}\in\{0,1\}}(\Pr((\hat{X}_{i}=x_{i},\hat{X}_{j}=x_{j}\mid z)-\Pr(\hat{X}_{i}=x_{i}\mid z)\;\Pr(\hat{X}_{j}=x_{j}\mid z))\;T_{\Delta}(x_{i},x_{j},z)$
 $=\sum_{x_{i},x_{j}\in\{0,1\}}\left|\Pr(\hat{X}_{i}=x_{i},\hat{X}_{j}=x_{j}\mid z)-\Pr(\hat{X}_{i}=x_{i}\mid z)\;\Pr(\hat{X}_{j}=x_{j}\mid z)\right|$
 $=8d(1-2d)$
-$=S_{i}(\tau,\tau)+8d^{2}.$
-B.3 Information Monotonicity under Naive Strategies
+$=S_{i}(X_{i},X_{j}\mid Z)+8d^{2}.$
+D.3 Information Monotonicity under Lazy-reporting Strategies
 Proof of 4.6.
-A naive strategy is a convex combination of truth-telling and an uninformative strategy.
-Therefore, it suffices to show that the expected score of truth-telling is strictly higher than that of any uninformative strategy.
-An uninformative strategy is essentially a mixture of two types of strategy.
+A lazy-reporting strategy is a convex combination of truth-telling and a no-effort strategy.
+Therefore, it suffices to show that the expected score of truth-telling is strictly higher than that of any no-effort strategy.
+A no-effort strategy is essentially a mixture of two types of strategy.
 First, $\mu_{i}(X_{i},Z_{i})$ is independent of both $X_{i}$ and $Z_{i}$.
@@
 Conditioned on $Z$, the Markov chain $Z_{i}\rightarrow\hat{Z}_{i}$ holds.
-Now, fix agent $j$’s naive strategy and let her report be $\hat{X}_{j}$.
-By the data processing inequality [Shannon, 1948], we have:
+Now, fix agent $j$’s lazy-reporting strategy and let her report be $\hat{X}_{j}$.
+By the data processing inequality shannon1948mathematical, we have:
 $I_{\text{TVD}}(\hat{Z}_{i};\hat{X}_{j}\mid Z)\leq I_{\text{TVD}}(Z_{i};\hat{X}_{j}\mid Z).$
-Next, suppose agent $i$ reporting $Z_{i}$, which is the score-maximizing naive strategy regardless of agent $j$’s strategy.
-For any naive strategy such that $\hat{X}_{j}=\nu_{j}(X_{j},Z_{j})$ that agent $j$ plays, applying the data processing inequality to agent $j$’s report yields:
+Next, suppose agent $i$ reporting $Z_{i}$, which is the score-maximizing lazy-reporting strategy regardless of agent $j$’s strategy.
+For any lazy-reporting strategy such that $\hat{X}_{j}=\nu_{j}(X_{j},Z_{j})$ that agent $j$ plays, applying the data processing inequality to agent $j$’s report yields:
 $I_{\text{TVD}}(Z_{i};\hat{X}_{j}\mid Z)\leq\max\left(I_{\text{TVD}}(Z_{i};Z_{j}\mid Z),I_{\text{TVD}}(Z_{i};X_{j}\mid Z)\right).$
-Therefore, 4.5 is sufficient for information monotonicity under naive strategies. ∎
-B.4 Learning the Scoring Function
-Proof of 4.7.
+Therefore, 4.5 is sufficient for information monotonicity under lazy-reporting strategies. ∎
+Appendix E Learning the Scoring Function
+In Section 4.1, we assumed that the underlying scoring function $T_{\Delta}$ is known.
+For complex signal structures, especially when the signal space is large, we have to learn the scoring function from the data.
+This introduces a concern: if a significant fraction of agents rely on LLMs, the learned scoring function $T_{\tilde{\Delta}}$ might be manipulated in a way that benefits the LLM-generated reports.
+We mitigate this concern using the following proposition.
+Let $S_{i}^{T}(\theta_{i},\theta_{j})$ be the expected score computed under the scoring function $T$ when agents’ strategies are $\theta_{i}$ and $\theta_{j}$ respectively.
+Proposition E.1.
+Suppose 4.1 and 4.2 hold with no error such that $\epsilon=\epsilon_{i}=0$ for any $i\in[n]$.
+Then, for any scoring function $T\in\{0,1\}^{|\Sigma|\times|\Sigma|}$ and any strategy profile $(\theta_{i},\theta_{j})$, $S_{i}^{T}(\theta_{i},\theta_{j})\leq S_{i}^{T_{\Delta}}(\tau,\tau)$, where $\tau$ denote the truth-telling strategy.
+Intuitively, the proposition holds because when $T_{\Delta}$ is known and agents are truth-telling, the expected score is $I_{\text{TVD}}(X_{i};X_{j}\mid Z)$.
+However, if $T_{\Delta}$ is mislearned, the expected score is upper bounded by $I_{\text{TVD}}(\hat{X}_{i};\hat{X}_{j}\mid Z)$.
+Since our assumptions ensure that $Z_{i}$ and $Z_{j}$ are independent given $Z$, the data processing inequality guarantees that $I_{\text{TVD}}(\hat{X}_{i};\hat{X}_{j}\mid Z)<I_{\text{TVD}}(X_{i};X_{j}\mid Z)$.
+E.1 implies that the expected score of the conditioned CA mechanism is maximized if every agent exerts full effort and reports $X_{i}$, while any strategy that manipulates the true scoring function will only (weakly) decrease the expected score.
+[^2] Here we assume that the principal can learn the correct $T_{\Delta}$ if everyone reports their full-effort signal.
+Proof of E.1.
 Note that when agent $i$ and $j$ play strategy $\theta_{i}$ and $\theta_{j}$ respectively and the scoring function is $T$,
@@
 $=$ $\sum_{z\in\Sigma}\Pr(Z=z)\sum_{x_{i},x_{j},z_{i},z_{j}\in\Sigma}(\Pr(x_{i},x_{j},z_{i},z_{j}\mid z)-\Pr(x_{i},z_{i}\mid z)\;\Pr(x_{j},z_{j}\mid z))\;T(\theta_{i}(x_{i},z_{i}),\theta_{j}(x_{j},z_{j}),z)$
-Because 4.2 and 4.1 hold with no error, by Lemma B.2 and Lemma B.1, it is easy to show that $\Pr(x_{i},x_{j},z_{i},z_{j}\mid z)=\Pr(x_{i},x_{j}\mid z)\;\Pr(z_{i}\mid z)\;\Pr(z_{j}\mid z)$ and $\Pr(x_{i},z_{i}\mid z)=\Pr(x_{i}\mid z)\;\Pr(z_{i}\mid z)$.
+Because 4.2 and 4.1 hold with no error, by Lemma D.2 and Lemma D.1, it is easy to show that $\Pr(x_{i},x_{j},z_{i},z_{j}\mid z)=\Pr(x_{i},x_{j}\mid z)\;\Pr(z_{i}\mid z)\;\Pr(z_{j}\mid z)$ and $\Pr(x_{i},z_{i}\mid z)=\Pr(x_{i}\mid z)\;\Pr(z_{i}\mid z)$.
 $=$ $\sum_{z\in\Sigma}\Pr(Z=z)\sum_{z_{i},z_{j}\in\Sigma}\Pr(z_{i}\mid z)\;\Pr(z_{j}\mid z)$
@@
 $=$ $S_{i}^{T_{\Delta}}(\tau,\tau)$
-Appendix C A Heuristic Generalization to Complex Signals
+Appendix F A Heuristic Generalization to Complex Signals
 We have focused on scenarios where the signal space is small, such as answering multi-choice questions like sentiment labeling, preference elicitation, or grading.
@@
 Depending on the information of interest, different embedding methods can be applied.
-For example, classic NLP embedding methods such as Word2Vec [Mikolov, 2013], TF-IDF [Manning, 2009], BERT and its variants [Devlin, 2018, Sanh, 2019], and GPT-3 [Brown et al., 2020] can be used to represent the language-level information within the responses.
+For example, classic NLP embedding methods such as Word2Vec mikolov2013efficient, TF-IDF manning2009introduction, BERT and its variants devlin2018bert, sanh2019distilbert, and GPT-3 brown2020language can be used to represent the language-level information within the responses.
 Existing embedding methods may struggle when the focus is on content-level information, such as the judgments in peer reviews.
@@
 Suppose the principal has a pre-defined set of questions about the responses, denoted as $\mathcal{Q}$ where $|\mathcal{Q}|=K$.
-Each question requires a (normalized) score between $-1$ and $1$, ensuring that the inner product between vectors can serve as a meaningful similarity measure.
+Each question requires a (normalized) score between 1-1 and $1$, ensuring that the inner product between vectors can serve as a meaningful similarity measure.
 Given a response $X_{i}$, the principal can prompt an LLM to iteratively get an answer to each question $q\in\mathcal{Q}$.
-This process converts each response $X_{i}$ into a vector $\phi(X_{i})$ of length $K$, and each entry is a score between $-1$ and $1$.
-Similar ideas have been applied to identify similar data points [Zeng et al., 2024], analyze peer review [Liang et al., 2024], and select features [Jeong et al., 2024].
+This process converts each response $X_{i}$ into a vector $\phi(X_{i})$ of length $K$, and each entry is a score between 1-1 and $1$.
+Similar ideas have been applied to identify similar data points zeng2024similar, analyze peer review liang2024mapping, and select features jeong2024llmselect.
 Computing the conditioned similarity.
@@
 By projecting the embedding vectors onto the orthogonal complement of $\phi(Z)$, we remove the information that already exists in $Z$.
-This projection approach is also used in NLP for “debiasing” word embeddings by eliminating similarity along a “biased direction” [Bolukbasi et al., 2016].
+This projection approach is also used in NLP for “debiasing” word embeddings by eliminating similarity along a “biased direction” bolukbasi2016man.
 Finally, we note that we do not assume the principal’s questions in $\mathcal{Q}$ to be i.i.d.
 Thus, the theoretical guarantees in Section 4.1 do not directly apply to this heuristic generalization; we instead rely on empirical evaluation to assess performance.
-Appendix D Experiments (Continue)
-D.1 Assumption Verifications (Continue)
-Figure 5 present the TVD mutual information between LLM samples and human responses conditioned on (different) LLM samples, $I_{\text{TVD}}(Z_{i};X_{j}\mid Z)$, which is the analogous results to Fig.
-2 on the remaining datasets.
-(a) Hatefulness
-(b) Offensiveness
-Figure 5: Analogous figures to Fig. 2 for other datasets.
-Table 2: TVD-Conditioned Mutual Information Between Two LLMs on the Preference Alignment Data
+Appendix G Experiments (Continue)
+G.1 Large Language Models
+In this work, we select five well-known LLMs to generate labels or annotations for all used datasets, including GPT-3.5-turbo, GPT-4, Gemma-2-2b-it [^3] https://huggingface.co/google/gemma-2-2b-it, Phi-3.5-mini-Instruct  [^4] https://huggingface.co/microsoft/Phi-3.5-mini-instruct, and Mistral-7B-Instruct-v0.3555https://huggingface.co/mistralai/Mistral-7B-Instruct-v0.3.
+The corresponding prompt templates are provided in Appendix H.
+Note that we choose only five models because other commonly used LLMs such as LLaMA and Gemini refuse to answer toxicity-related questions.
+G.2 Details of the Baseline Methods
+Output Agreement (OA)  Perhaps the most straightforward idea to score an agent is to count the frequency of agreements between her responses and another agents’ responses.
+Specifically, while scoring agent $i$, we iteratively pair her with a different agent $j$ and let $M_{i,j}$ denote the set of questions answered by both agents.
+Then, the score of agent $i$ is the average of the empirical probability of agreement, i.e.
+$\frac{1}{n}\sum_{j\in[n]\backslash\{i\}}\sum_{q\in M_{i,j}}\frac{\mathbbm{1}[\hat{X}_{i,q}=\hat{X}_{j,q}]}{|M_{i,j}|}$ where $\hat{X}$ is the crowdsourced dataset with rows representing agents and columns representing questions.
+[^6] Let $\frac{0}{0}=0$.
+A score of 1 thus indicates perfect agreement while a score of 0 indicates no agreement at all.
+Correlated Agreement (CA)  We use the CA mechanism shnayder2016informed as a baseline which does not incorporate the principal’s information $Z$.
+Details of the mechanism have been presented in Algorithm 1.
+Expectation-Maximization (EM)  The EM algorithm is a well-known approach to estimating the reliabilities of noisy agents and better recovering the ground truth considering the reliabilities of agents [dawid1979maximum].
+Each agent $i$ is assumed to make mistakes according to a confusion matrix $\Gamma^{i}$ where $\Gamma^{i}_{h,l}$ denotes the probability that agent $i$ reports $h$ when the ground truth answer is $l$.
+The EM algorithm first initializes a prior of the ground truth for each question and a confusion matrix for each agent.
+Then, it iteratively goes through an E-step and an M-step.
+In the E-step, the algorithm computes the posterior of the ground truth of each question given the prior and the confusion matrix.
+In the M-step, the confusion matrix of each agent is updated using the posterior computed in the E-step.
+The above two steps are iterated until the likelihood of $D$ converges.
+Finally, we score the reliability of agent $i$ as $\sum_{h\in\Sigma}\Gamma^{i}_{h,h}\tilde{P}(h)$ where $\tilde{P}(h)$ is the empirical marginal distribution of observing $h$ in the dataset.
+Conditioned Agreement ($\text{OA}_{Z}$)  The first three measures do not utilize the principal’s information about the cheap signal, denoted as $Z$.
+Note that in our experiments, $Z$ is the vector of labels generated by a particular LLM.
+The conditioned agreement metric scores a agent if she can provide agreements to another agent in addition to $Z$.
+Similar to the agreement metric, the score of agent $i$ is the average of the empirical probability of agreement conditioned on $Z$, i.e.
+$\frac{1}{n}\sum_{j\in[n]\backslash\{i\}}\sum_{q\in M_{i,j}}\frac{\mathds{1}[\hat{X}_{i,q}=\hat{X}_{j,q}\&\hat{X}_{i,q}\neq Z_{q}]}{|M_{i,j}|}$.
+Intuitively, if agent $i$ uses an LLM whose generated labels are similar to $Z$, then $\hat{X}_{i,q}\neq Z_{q}$ will happen with a small probability, meaning that the conditioned agreement score will be small.
+G.3 Assumption Verifications (Continue)
+Table 2, Table 3, Table 4 and Table 5 present the mutual information $I_{\text{TVD}}(Z_{i};Z_{j}\mid Z)$ on four datasets.
+$Z$ and $Z_{i}$ are sampled by the same LLM indicated on the rows and $Z_{j}$ is sampled by the LLM indicated on the columns.
+The first observation is that all values displayed in black exceed $0.05$, which represents the mutual information of the random reporting baseline.
+This implies that $I_{\text{TVD}}(Z_{i};Z_{j}\mid Z)>\epsilon$ holds consistently, rejecting hypothesis 2.
+Regarding hypothesis 4, we observe that the average correlations between two LLMs are generally weaker than those between two human agents, provided $Z_{i}$ and $Z$ are independently generated by the same model.
+This trend is evident in the tables, where almost all bracketed values are positive, with one exception.
+[^7] The exception occurs in the preference alignment dataset, where samples of $Z_{i},Z_{j}$, and $Z$ all generated by GPT-4.
+Table 2: Conditional TVD Mutual Information Between Two LLMs on the Preference Alignment Data
 GPT-3.5 GPT-4 Gemma Mistral Phi
@@
 Each entry represents $I_{\text{TVD}}(Z_{i};Z_{j}\mid Z)$ (Difference: $I_{\text{TVD}}(X_{i};X_{j}\mid Z)-I_{\text{TVD}}(Z_{i};Z_{j}\mid Z)$), where $Z_{i}$ and $Z_{j}$ are generated by the models in the row and column respectively, and $Z$ uses the same model as $Z_{i}$.
-Table 2, Table 3, Table 4 and Table 5 present the mutual information $I_{\text{TVD}}(Z_{i};Z_{j}\mid Z)$ on four datasets.
-$Z$ and $Z_{i}$ are sampled by the same LLM indicated on the rows and $Z_{j}$ is sampled by the LLM indicated on the columns.
-The first observation is that all values displayed in black exceed $0.05$, sometimes significantly so, which represents the mutual information of the random reporting baseline.
-This implies that $I_{\text{TVD}}(Z_{i};Z_{j}\mid Z)>\epsilon$ holds consistently, rejecting hypothesis 2.
-Regarding hypothesis 4, we observe that the average correlations between two LLMs are generally weaker than those between two human labelers, provided $Z_{i}$ and $Z$ are independently generated by the same model.
-This trend is evident in the tables, where almost all bracketed values are positive, with one exception.***The exception occurs in the preference alignment dataset, where samples of $Z_{i},Z_{j}$, and $Z$ all generated by GPT-4.
-Table 3: TVD-Conditioned Mutual Information Between Two LLMs on the Hatefulness Labeling Data
+Table 3: Conditional TVD Mutual Information Between Two LLMs on the Hatefulness Labeling Data
 GPT-3.5 GPT-4 Gemma Mistral Phi
@@
 Phi 0.0088 (0.1658) 0.0115 (0.1631) 0.0618 (0.1128) 0.0526 (0.1220) 0.0803 (0.0943)
-Table 4: TVD-Conditioned Mutual Information Between Two LLMs on the Offensiveness Labeling Data
+Table 4: Conditional TVD Mutual Information Between Two LLMs on the Offensiveness Labeling Data
 GPT-3.5 GPT-4 Gemma Mistral Phi
@@
 Phi 0.0320 (0.1604) 0.0597 (0.1327) 0.0334 (0.1591) 0.0604 (0.1321) 0.0651 (0.1273)
-Table 5: TVD-Conditioned Mutual Information Between Two LLMs on the Toxicity Labeling Data
+Table 5: Conditional TVD Mutual Information Between Two LLMs on the Toxicity Labeling Data
 GPT-3.5 GPT-4 Gemma Mistral Phi
@@
 Phi 0.0416 (0.1756) 0.0517 (0.1655) 0.0382 (0.1790) 0.0472 (0.1701) 0.0666 (0.1507)
-D.2 Detecting Cheating Agents (Continue)
-D.2.1 Additional Figures and Tables
-In Fig.
-6, we present additional figures on the toxicity dataset, analogous to Fig.
-4(a), for other types of LLM cheaters.
-Analogous results to Table 1 for other datasets are summarized in Table 6, Table 7, Table 8, and Table 9.
-These results support our main claim that, although our method is occasionally slightly outperformed by certain baselines, it consistently demonstrates the most robust performance across all tested settings.
+G.4 Detecting Low-effort Agents (Continue)
+G.4.1 Additional Figures and Tables
 (a) GPT-4
@@
 (e) Phi-3.5-mini-Instruct
-Figure 6: The area under the ROC curve (AUC) for various methods while varying the fraction of LLM cheaters.
-These are the analogous figures for Fig.
-4(a) for the remaining LLMs on the toxicity labeling dataset.
-Table 6: Average AUC (and bottom 10% quantiles) across datasets and methods with GPT-3.5 cheaters.
+Figure 4: The area under the ROC curve (AUC) for various methods while varying the fraction of LLM-reliant agents.
+These are the analogous figures for Figure 3(b) for the remaining LLMs on the toxicity labeling dataset.
+Table 6: Average AUC (and bottom 10% quantiles) across datasets and methods with GPT-3.5-reliant agents.
 Dataset OA CA OAZ EM CAZ (ours)
@@
 Preference Alignment 0.88 (0.67) 0.86 (0.69) 0.88 (0.76) 0.96 (0.91) 0.91 (0.83)
-Table 7: Average AUC (and bottom 10% quantiles) across datasets and methods with Mistral cheaters.
+Table 7: Average AUC (and bottom 10% quantiles) across datasets and methods with Mistral-reliant agents.
 Dataset OA CA OAZ EM CAZ (ours)
@@
 Preference Alignment 0.86 (0.66) 0.89 (0.75) 0.94 (0.87) 0.94 (0.87) 0.93 (0.88)
-Table 8: Average AUC (and bottom 10% quantiles) across datasets and methods with Gemma cheaters.
+Table 8: Average AUC (and bottom 10% quantiles) across datasets and methods with Gemma-reliant agents.
 Dataset OA CA OAZ EM CAZ (ours)
@@
 Preference Alignment 0.80 (0.58) 0.87 (0.73) 0.76 (0.62) 0.92 (0.85) 0.92 (0.86)
-Table 9: Average AUC (and bottom 10% quantiles) across datasets and methods with Phi cheaters.
+Table 9: Average AUC (and bottom 10% quantiles) across datasets and methods with Phi-reliant agents.
 Dataset OA CA OAZ EM CAZ (ours)
@@
 Preference Alignment 0.85 (0.68) 0.82 (0.63) 0.82 (0.63) 0.97 (0.94) 0.93 (0.88)
-D.2.2 Detection Performance vs. Number of Samples (Continue)
-Figure 7 presents similar results to Fig. 4(b) for the hatefulness, offensiveness, and toxicity labeling datasets.
-(a) Hatefulness
-(b) Offensiveness
-(c) Toxicity
+In Figure 4, we present additional figures on the toxicity dataset, analogous to Figure 3(b), for other LLMs.
+Analogous results to Table 1 for other datasets are summarized in Table 6, Table 7, Table 8, and Table 9.
+These results support our main claim that, although our method is occasionally slightly outperformed by certain baselines, it consistently demonstrates the most robust performance across all tested settings.
+G.4.2 Detection Performance vs. Number of Samples
+We further examine how many questions the principal needs to sample to effectively detect low-effort agents.
+In Figure 5, we report the AUC of the conditioned CA mechanism with the fraction of LLM-reliant agents fixed at $0.15$, while the fractions of random and biased agents are independently sampled from $[0,0.2]$.
+We vary the sampling rate from 20% to 100%.
+The results vary across the datasets and the LLMs that agents use.
+Except for the hatefulness labeling dataset with GPT-3.5-reliant agents, sampling 50% of the questions is sufficient to outperform the baseline CA mechanism.
+(a) Preference
+(b) Hatefulness
+(c) Offensiveness
+(d) Toxicity
 Figure 5: The area under the ROC curve (AUC) for various methods while varying the fraction of questions the principal can sample.
-D.2.3 Conditioning on samples of multiple LLMs
+G.4.3 Conditioning on samples of multiple LLMs
 In practice, agents may employ a variety of LLMs for low-effort response generation, and the principal lacks prior knowledge of which models they are using.
@@
 Intuitively, human agents will perform consistently well when conditioning out any of the LLMs’ information.
-However, LLM cheaters will perform poorly in at least one of the iterations and thus lead to a lower final score.
-In Fig.
-8, we replicate Fig.
-4(a) for the setting where the cheaters randomly pick two LLMs to cheat with half-half probability while the principal iteratively conditions out his samples for these two LLMs’ answers.
-Our method exhibits the most robust performance across all settings, echoing the previous conclusion.
-We further note that the problem of LLM-assisted cheating is less of a concern when each cheater uses a unique LLM.
-This is because LLM responses are less correlated with each other and can be easily detected as outliers under classic crowdsourcing evaluation methods, such as the EM algorithm under the DS model.
-(a) Mistral & GPT-3.5
-(b) Mistral & GPT-4
+However, LLM-reliant agents will perform poorly in at least one of the iterations and thus leading to a lower final score.
+In Figure 6, we replicate Figure 3(b) for the setting where the LLM-reliant agents randomly pick two LLMs with half-half probability while the principal iteratively conditions out his samples for these two LLMs’ answers.
+Our method exhibits the most robust performance across all datasets, echoing the previous conclusion.
+(a) Gemma & GPT-3.5
+(b) Gemma & GPT-4
 (c) GPT-3.5 & GPT-4
-Figure 8: Area under the ROC curve (AUC) for various methods as the fraction of LLM cheaters increases.
-In each case, LLM cheaters randomly choose between two LLMs to generate responses, and the score is defined as the minimum of the two conditioned CA scores, each conditioned on samples from one of the two LLMs.
-D.3 Complex Signals
-We further test the idea illustrated in Appendix C that generalizes our method to high-dimensional settings.
+Figure 6: Area under the ROC curve (AUC) for various methods as the fraction of LLM-reliant agents increases.
+In each case, LLM-reliant agents randomly choose between two LLMs to generate responses, and the score is defined as the minimum of the two conditioned CA scores, each conditioned on samples from one LLM.
+G.5 Complex Signals
+We further test the idea illustrated in Appendix F that generalizes our method to high-dimensional settings.
 In this experiment, we randomly sample 500 papers from the ICLR 2024 OpenReview data and select three human reviews per paper as human responses $X$.
-For each of the same batch of papers, we further use each of the five LLMs to generate three AI reviews using three similar-meaning yet different prompts (Appendix E).
-We randomly pick one of the generated reviews as the principal’s sample $Z$, and the other two generated reviews will be used to simulate LLM-cheaters’ responses.
-We prompt GPT-4 to process each human review and LLM-generated review into a $19$-dimension vector with the prompt shown in Appendix E.
+For each of the same batch of papers, we further use each of the five LLMs to generate three AI reviews using three similar-meaning yet different prompts (Appendix H).
+We randomly pick one of the generated reviews as the principal’s sample $Z$, and the other two generated reviews will be used to simulate the responses of LLM-reliant agents.
+We prompt GPT-4 to process each human review and LLM-generated review into a $19$-dimension vectors with the prompt shown in Appendix H.
 We further simulate a review embedding tensor with a size of $n\times m\times 19$, where there are $n=188$ reviewers and $m=500$ papers in total.
 Each reviewer randomly reviews $k\in\{5,8\}$ papers.
-An $\alpha_{llm}$ fraction of reviewers are LLM cheaters whose review vectors are replaced with a particular type of LLM-generated reviews.
-An $\alpha_{p}$ fraction of reviewers are prior cheaters, where the score in each entry of the review vector is randomly sampled according to the marginal distribution in the dataset.
-An $\alpha_{r}$ fraction of reviewers are random cheaters, where the score in each entry of the review vector is sampled uniformly at random.
+An $\alpha_{llm}$ fraction of reviewers are LLM-reliant agents whose review vectors are replaced with a particular type of LLM-generated reviews.
+An $\alpha_{p}$ fraction of reviewers are prior-reporting agents, where the score in each entry of the review vector is randomly sampled according to the marginal distribution in the dataset.
+An $\alpha_{r}$ fraction of reviewers are random agents, where the score in each entry of the review vector is sampled uniformly at random.
 In our experiments, we independently sample each of $\alpha_{llm}$, $\alpha_{p}$, and $\alpha_{r}$ uniformly from the interval $[0,0.1]$ and take average over $100$ trials.
 While constructing the review embedding tensor, we guarantee that each paper has approximate $3$ reviews so that the variance of the number of reviewers per paper is controlled.
-Again, we consider five types of LLMs and for each type of LLM cheater, the principal’s sample $Z$ is an independent generation of the same LLM under a different prompt.
+Again, we consider five types of LLMs and for each type of LLM-reliant agent, the principal’s sample $Z$ is an independent generation of the same LLM under a different prompt.
 We consider three methods to score the reviewers.
   • Cosine Similarity: Scores each reviewer based on the average cosine similarity between their review vector and the review vectors of other reviewers for the same paper.
-  • Projected Cosine Similarity: Applies the method described in Appendix C, where each review vector is projected onto the subspace orthogonal to the principal reviewer’s sampled embedding before computing similarity.
+  • Projected Cosine Similarity: Applies the method described in Appendix F, where each review vector is projected onto the subspace orthogonal to the principal reviewer’s sampled embedding before computing similarity.
   • Negative Similarity to $Z$: Scores each reviewer by taking the negative average cosine similarity between their review vector and the principal reviewer’s sampled vector $Z$.
-We present the results in Fig.
-9.
+We present the results in Figure 7.
 It is clear that conditioning out the principal’s sampled LLM reviews is more effective than computing the cosine similarity directly.
-Furthermore, naively removing reviews that are similar to $Z$ has a large variance, whose performance greatly depends on the existence of random and prior cheaters.
+Furthermore, naively removing reviews that are similar to $Z$ has a large variance, whose performance greatly depends on the existence of random agents and prior-reporting agents.
 We further note that there is a big room for improvement.
@@
 Figure 7: A comparison of the area under the ROC curve (AUC) of various methods in the peer review dataset.
-Appendix E Prompts Templates
-The prompt templates for two different dataset settings are provided in Table E and Table E, respectively.
+Appendix H Prompts Templates
+We adopt our prompt templates based on recent studies [pang2024improving, liu2024automatic, pang2025token].
+The specific templates used for the two dataset settings are presented in Table H and Table H, respectively.
 Prompt Template for Toxicity Labeling System Prompt: As a helpful system judgment assistant, your task is to answer the following questions for the given comment.